A Bibliography of Publications in Journal of the Institute of Mathematics and Its Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.10

Title word cross-reference

(1 + x^{2})^{−1/2} [494, 557]. (p(t)x′)′ + q(t)x^{2n−1} = 0 [796]. 1/2 [368]. 2 [419]. 2p [547]. 3 [419]. [0, ∞) [790]. _{2}F_{0} [312]. A + B^{∗} log(1 + CX) [251]. A + B log(1 + CX) [249]. cos(iπ/n) [590]. exp(z) [742]. f(x) [319]. J_{MN} [804]. k [544]. L^{2} [52]. L_{1} [802]. M [241]. M^{2} [754]. N [585, 652]. Nth [316]. O(h^{6}) [745, 744]. θ [475]. X [161].

-dimensional [419]. -method [475]. -monomer [652]. -pole [547]. -spline [161].

12 [711]. 15 [561].

A. [418]. A.D.I. [345, 526]. Abel [25, 695]. abscissas [36]. absorbing [186]. absorption [659]. Accelerating [433]. acceleration [343]. Accuracy [651, 15, 140, 285, 420, 508, 522, 526, 616, 657]. Accurate [223, 277, 224, 610, 721]. Acoustic [83, 84, 88, 556, 17, 497]. Acoustics [444]. action [617, 618]. actual [465]. adaptation [733]. adaptive [387, 542]. addenda [665]. Addendum [561]. Aerodynamic [441]. aerofoil [414]. Age [360, 622]. Age-dependent [360]. age-replacement [622]. aid [469]. algebra [118, 544, 48]. Algebraic [651, 151, 277, 804]. Algorithm [180, 712, 24, 98, 124, 148, 218, 293, 344, 397, 404, 408, 442, 469, 501, 538, 549, 569, 587, 600, 610, 623]. algorithms [73, 74, 75, 97, 98, 119, 231, 233, 624, 625, 653]. aligned [685]. allocation [469]. almost [427, 482, 736]. along [664, 665]. Alternating [523, 524, 285, 343, 345, 370, 371]. alternative [96, 387]. always [349]. Amplitude [7]. analogue [283, 745, 744]. Analysis [195, 591, 56, 321, 393, 405, 435, 527, 543, 600, 656]. analytic [61]. Analytical [336]. angular [380]. anisotropic [156, 413, 654, 664, 665, 684, 786]. Anodic [290, 291]. any [791]. aperiodic [7]. aperture [17, 748]. apertures [598]. apex [93]. appendix [336, 427]. Appl [561, 711]. Application [262, 264, 641, 804, 1, 159, 161, 380, 400, 403, 408, 410, 449, 536, 566]. Applications [55, 272, 71, 173, 324, 543, 558, 674, 775, 41]. applied [48, 88, 255, 393, 579, 595, 654]. approach [109, 266, 552, 645, 722]. approaches [512]. approximant [58]. approximants [57, 172, 241, 353, 351, 364, 365, 366, 412, 483, 658, 742, 804]. Approximate [699, 58, 64, 163, 203, 284, 335, 391]. approximating [646, 790]. Approximation [250, 367, 3, 249, 251, 252, 265, 373, 384, 409, 489, 528, 537, 539, 635, 720, 719, 732, 756, 757, 772, 769, 802]. Approximations [4, 61, 62, 140, 263, 280, 401, 484, 494, 557, 636, 647, 741, 770]. arbitrary [60, 408, 497, 788]. arcs [531]. areas [581]. Arising [611, 186, 217, 448, 527, 592, 593, 707]. Arithmetic [30]. Arrays [403]. arterial [519]. aspect [497]. aspects [120, 298, 636]. Assessment [622]. associated [10, 43, 313, 561, 562, 564]. asymmetrically [802]. Asymptotic [86, 131, 253, 347, 596, 1, 168, 206, 208, 322, 381, 379, 425, 452, 517, 518, 607, 725, 752, 796]. asymptotically [565]. augmented [47]. Automatic [202, 116]. autonomous [649]. average [23].

Axially [554, 183].

B [111, 179, 181, 310, 318]. B-Spline

[181, 111]. B-Splines [179, 318]. Bäcklund [159]. backward [122, 540]. Bairstow [29]. Banach [675]. Band [85, 458]. barrier [281, 391]. barriers [426, 700]. bars [154]. based [75, 539, 710, 711]. basic [283]. basis [111, 127, 530, 756, 757]. be [150]. beam [595, 720, 719]. beams [707]. bearing [688]. behaviour [425]. Belousov [633]. Bending [573, 595]. Bergman [485]. Bernoulli [112]. Bessel [346, 681, 761]. Best [802, 11, 62, 544, 770]. best-possible [11]. between [135, 145, 146, 359, 495]. Bickley [612]. Bifurcated [577]. biharmonic [45, 149]. bilinear [581]. bimaterial [43]. biomechanics [378]. Bivariational [59]. Bloch [125]. Block [320, 72, 242, 276]. bodies [358, 392, 396, 685]. body [396, 752, 758, 773]. bottoms [687]. Bound [30, 32, 458, 631]. boundaries [171, 410, 531]. Boundary [155, 303, 521, 696, 701, 15, 42, 56, 66, 69, 95, 135, 138, 139, 140, 141, 158, 176, 182, 186, 185, 183, 184, 189, 188, 190, 196, 205, 206, 229, 265, 269, 270, 275, 287, 304, 311, 315, 336, 334, 361, 372, 378, 383, 417, 419, 465, 500, 502, 508, 514, 536, 542, 559, 560, 563, 564, 594, 604, 620, 632, 640, 659, 676, 694, 707, 708, 713, 726, 731, 745]. boundary-layer [206]. boundary-value [95, 176, 205, 229, 563]. bounded [449]. Boundedness [454]. Bounds [85, 99, 235, 437, 7, 11, 16, 21, 33, 58, 59, 147, 295, 335, 419, 477, 478, 574, 717, 767]. Branch [241]. Bromwich [655]. bulk [577, 576, 766]. bulk- [766]. bulk-reacting [577, 576].

C [310]. cable [13]. calculating [64, 770]. Calculation [388, 14, 318, 323, 353, 449, 532, 641]. calculations [220, 309, 492, 511, 640]. calculus [680]. can [150]. capacitance [498]. capacity [287]. Carlo [626]. case [28]. cases [676]. Cauchy [467, 578]. cautious [387]. centres [408]. Certain [466, 494, 557, 61, 62, 195, 425, 451, 543]. changes [236]. channel [83, 84, 209, 462]. characteristics [741]. characterizing [638]. charge [340, 748]. charge-current [340]. Chebychev [343]. Chebyshev [3, 226, 227, 228, 249, 251, 252, 261, 375, 591, 770]. chemical [448, 593]. Chisholm [172, 412]. choice [203, 384, 475]. Cholesky [712]. circular [17, 114, 436, 556, 788, 748]. class [2, 20, 32, 34, 51, 61, 68, 97, 98, 119, 121, 120, 122, 166, 208, 247, 324, 330, 331, 560, 564, 695, 703, 726, 755, 772, 780]. classes [543]. Clenshaw [589, 591]. Closed [647, 464]. Closed-form [647]. closely [498]. co [297, 348, 405]. co- [297]. co-current [348]. co-ordinates [405]. coaxial [359]. coefficient [527]. coefficients [322, 325, 347, 388, 479, 480, 495, 526, 532, 646]. collection [638]. Collisionless [597]. collocation [104, 205, 334, 461, 484]. column [279]. combinations [76]. Comments [494, 668, 612, 666]. common [654]. compact [765]. comparative [315]. Comparison [243, 133, 245, 791]. Complementary [20, 169, 280, 22, 299]. complete [258]. Complex [257, 24, 95, 367, 582, 767]. components [377]. composition [652]. Compressible [640, 234, 641]. compression [354]. Computation [799, 800, 26, 111, 542, 706, 765]. Computational [53, 599, 120]. Computations [326]. Computer [358, 568]. computing [633, 764]. concentrated [287]. concentration [479, 480]. concentration-dependent [479, 480]. concept [110]. concerned [282]. Condensed [244]. Condition [99, 763]. conditional [675]. conditioned [248]. conditioning [274]. Conditions [260, 706, 67, 138, 139, 327, 334, 372, 640, 694, 696]. conducting [664, 665]. conduction [287, 322, 420, 421, 606]. cone [723]. configuration [60]. confluent [725]. conformal [410, 485, 604]. Conjugate [230, 231, 429, 46]. connected [357, 485]. connecting [316]. Considerations [799, 800, 97, 391]. consistent [605]. consistently [549]. constant [325, 347]. constants [496, 655]. constrained [75, 294, 404, 705, 738, 746]. constraints [339, 398]. constricted [697]. construction [530, 644, 684]. contact [449, 450, 601]. containing [56, 183, 430]. continuation [386, 398]. Continued [151, 516, 558]. Continuous [144, 705, 27, 777, 790]. Continuum [41]. contour [797]. control [40, 51, 116, 211, 337, 382, 424, 488, 512, 617, 618, 631, 780]. controllability [155, 789]. controllable [40]. Controlling [729]. controls [706]. convection [87, 207]. convective [46, 783]. Convergence [92, 215, 373, 429, 548, 658, 742, 24, 50, 49, 61, 75, 97, 98, 101, 136, 310, 308, 433, 438, 490, 533, 584, 585, 586, 623, 639, 644, 656, 661, 706, 738, 772]. convergent [349]. converging [195]. Converse [474]. convex [130]. convexity [190]. convolution [366]. cooled [416]. cooling [128]. coordinate [563, 564]. coordinates [616]. Copolymer [652]. Corrected [484]. Corrections [89, 91]. corrector [5]. correspond [516]. Corrigendum [82]. Cosserat [740]. cost [617, 631]. Cotes [36]. countermeasures [400]. couple [472]. coupled [145, 146, 570, 667]. coupling [389]. crack [238, 237]. cracks [43]. criterion [337, 602]. critical [196, 306, 414]. crossing [362]. crude [419]. cubic [15, 67, 68, 161, 185, 205, 385, 457, 606, 635, 647, 792]. cubic-spline [606]. current [340, 341, 348, 638]. curse [503]. Curtis [589]. Curve [160, 178, 317, 779]. curved [127, 449, 530, 531, 528, 529, 677, 757]. curves [339, 447]. Curvilinear [563, 564]. cut [157]. cycle [535]. cycles [393]. Cyclic [92]. cylinder [156, 357, 556, 692]. cylinders [114, 554]. cylindrical [137, 154, 355, 392, 420, 577, 576, 616, 681].

D [418]. dam [190]. damped [324, 468, 538, 774]. Damping [774, 473]. Data [298, 385, 541, 581, 793, 803]. dates [602]. decision [171]. deep [281]. Deficiencies [621]. defined [62, 483]. defining [581]. definite [91, 274, 442]. deflation [102, 296]. deflection [707, 744]. deformation [11, 507]. deformations [270, 554]. degeneracy [636]. degree [14, 316, 471, 499, 646]. Delay [18, 86, 134, 473, 552, 754]. delay-differential [552]. Delay-minimizing [18]. delayed [658]. delta [93, 722, 797]. Demagnetization [392]. density [313, 314, 790]. dependence [316, 705]. dependent [69, 184, 282, 297, 360, 479, 480, 675, 760]. depending [465]. Derivation [368, 509, 96, 660, 670]. derivative [523]. derivatives [26, 90, 111, 152, 153, 272, 316, 524, 742]. derived [406]. descent [79]. describing [469, 534]. design [124, 298]. determinants [312]. determination [277, 605, 646, 703]. determined [307]. determining [671]. deterministic [194]. deviator [778]. device [279]. devices [622]. diagonal [565]. Diagonalization [257]. diagonalizes [24]. Diagonally [123, 124]. diameter [175]. dielectric [642]. difference [44, 138, 139, 140, 141, 223, 224, 254, 273, 345, 369, 395, 420, 421, 425, 459, 490, 496, 513, 514, 635, 660, 661, 675, 745, 744, 774, 791]. differences [203, 492]. different [354]. differential [33, 34, 65, 86, 119, 144, 147, 164, 208, 223, 224, 228, 225, 235, 247, 261, 273, 275, 277, 303, 302, 320, 327, 335, 334, 338, 344, 380, 382, 422, 423, 448, 460, 461, 469, 473, 489, 490, 527, 548, 553, 552, 588, 614, 637, 638, 653, 650, 660, 678, 694, 705, 734, 741, 744, 754, 755, 776, 804]. differential-algebraic [804]. differentiation [122, 435, 637]. ’diffraction [167, 8, 288, 642, 664, 665, 17, 131, 380, 693, 786]. Diffusion [402, 403, 20, 32, 168, 167, 186, 282, 284, 479, 480, 527, 616, 659, 717]. dimension [234]. dimensional [12, 50, 60, 114, 282, 299, 328, 345, 348, 370, 413, 419, 459, 553, 601, 606, 635, 405]. dimensionality [503]. dimensions [462, 521, 523, 585, 627, 673, 687, 728]. Dipole [456, 783]. Dirac [658]. Direct [246, 219]. direction [90, 285, 343, 345, 370, 371, 523, 524]. directions [230]. Dirichlet [37, 56, 484, 691, 732]. disc [772]. discontinuity [696]. discontinuous [493, 776, 777]. discrete [3, 527, 581, 584, 720, 753]. Discretization [70]. discretized [383]. discs [183]. discussion [624]. disjoint [402, 403]. disk [664, 665]. disks [37, 38, 39, 359]. Dislocations [472]. dispersion [447]. dispersive [476, 684, 727]. displacement [255, 449]. dissimilar [788]. dissipation [232]. dissipative [476, 727]. distributed [51, 617, 618]. distribution [28, 340, 341, 783]. distributional [272]. Distributions [35, 76]. disturbances [329]. divergent [163]. divided [492]. dock [212, 481]. domain [53, 732, 766]. domains [402, 403, 485]. domination [574]. Doppler [519]. double [81, 82, 97, 98, 131, 483, 537]. double-rank [97, 98]. Downdating [712]. drag [773]. drop [615]. droplet [80]. drug [403]. Dual [57, 299, 399, 247, 579, 698, 699]. dualities [134]. duct [576]. ducts [577, 576]. due [213, 441, 602]. Duffing [211]. Dyke [427]. dynamic [232, 266, 324, 469, 507, 572]. dynamical [649]. dynamically [40]. Dynamics [117, 638, 766].

Edge [167, 238, 239]. effect [207, 239, 281, 334, 766]. effectiveness [622]. Effects [712, 209, 305, 502]. Efficiency [683]. efficient [108, 397]. eigen [23]. Eigenfunction [797, 764]. eigenfunctions [765]. eigenproblem [599]. eigensolution [432]. Eigenvalue [162, 429, 81, 82, 106, 105, 165, 380, 386, 419, 434, 430, 592, 603, 776, 782, 805]. Eigenvalues [610, 26, 235, 258, 667, 683, 694, 765, 767]. eigenvectors [26]. Einstein [340, 341, 342]. Elastic [154, 60, 80, 130, 129, 156, 197, 229, 255, 396, 409, 447, 449, 519, 571, 739, 778, 788]. elastic-perfectly [778]. Elastic-plastic [154]. elasticity [159, 551, 704]. elastodynamics [78]. elastostatics [238, 237, 450]. electric [166, 427, 638]. electrochemical [265, 290, 291, 297]. electrochemistry [448]. electrohydrostatic [502]. Electromagnetic [786, 288, 445, 453]. electrostatic [114, 299]. electrostatics [113]. element [127, 191, 192, 220, 221, 265, 372, 492, 508, 530, 531, 542, 546, 545, 550, 644, 738]. elements [127, 471, 530, 528, 529, 759, 768, 775]. Elimination [85, 99, 202, 243, 643, 63, 431, 510]. ellipsoid [691]. ellipsoidal [381, 379, 379]. elliptic [65, 95, 106, 105, 158, 157, 254, 303, 361, 369, 477, 513, 514, 575, 585, 594, 604, 709, 723, 729]. embedded [130]. encountered [511]. End [67]. Energy [570, 279, 478]. envelope [684]. epidemic [714, 715]. equal [316]. equally [645, 647]. Equation [55, 19, 25, 45, 86, 112, 114, 132, 148, 168, 167, 176, 226, 227, 277, 283, 285, 302, 322, 370, 382, 413, 417, 420, 421, 425, 459, 527, 555, 595, 606, 609, 612, 635, 644, 663, 720, 719, 723, 728, 743, 785, 787, 789]. Equations [201, 246, 611, 651, 1, 20, 31, 33, 34, 46, 47, 52, 58, 64, 65, 70, 103, 104, 119, 125, 144, 150, 149, 164, 163, 194, 208, 216, 217, 219, 218, 223, 224, 228, 225, 241, 245, 247, 248, 254, 256, 261, 262, 263, 264, 273, 274, 275, 303, 320, 327, 335, 334, 338, 344, 349, 351, 365, 366, 380, 381, 394, 395, 397, 422, 423, 448, 460, 461, 463, 464, 469, 473, 475, 477, 479, 480, 486, 489, 490, 496, 500, 501, 506, 509, 513, 514, 523, 524, 526, 548, 549, 550, 553, 552, 575, 587, 588, 614]. equations [616, 619, 628, 629, 637, 638, 653, 650, 657, 660, 667, 675, 678, 694, 695, 729, 734, 741, 744, 754, 755, 769, 771, 776, 777, 784, 806]. equi [14]. equi-spaced [14]. equiangular [664, 665]. equipment [621]. equivalence [112]. equivalent [661]. Errata [331, 665]. Error [33, 32, 56, 100, 147, 451, 578, 591, 600, 651, 712, 11, 16, 123, 225, 335, 435, 477, 517, 539, 596, 629, 648, 656, 729]. Errors [85, 309, 313, 647]. escape [327]. establishing [649]. Estimates [651, 123, 225, 578, 648]. estimation [65, 134, 200, 313, 314, 538, 541, 795]. estimators [702]. evaluate [655]. Evaluating [591, 90, 108, 578, 590]. Evaluation [43, 179, 181, 511, 346, 491]. eversion [129]. evolution [70, 135, 206, 358, 754]. Exact [125, 382, 518, 311, 391, 508, 629]. exactly [372]. example [40]. examples [336, 798]. exchangers [300, 301]. existence [454, 678, 772]. Expansion [681, 1, 253, 532, 608, 784]. Expansions [679, 50, 131, 196, 271, 368, 375, 381, 379, 516, 517, 518, 533, 725, 752, 797]. Explicit [16, 517, 747, 198]. explosion [306]. explosions [117]. Exponential [799, 800, 250, 252]. exponential-polynomial [250]. expressed [91, 89, 581]. extended [196]. extensible [13]. Extension [29, 572, 129, 354, 515, 537]. exterior [276]. externally [688]. Extrapolated [205, 215, 370, 371, 418]. Extrapolation [438, 690, 689]. extremum [247, 399, 579, 698, 699, 733, 762].

F. [668]. factor [21]. Factorization [712]. factors [29, 43]. failing [621]. Falsi [304]. families [768]. family [459, 702]. fast [126, 216, 219, 218, 221, 233, 369, 610, 612, 394]. Fehlberg [422, 423]. fermion [511]. fibre [270, 554, 731]. fibre-reinforced [270, 554, 731]. fibres [736]. field [415, 427, 453]. fields [340, 341, 342, 685]. film [239]. filtered [634]. filters [166]. Final [71, 474]. Final-stability [71, 474]. finding [48, 258, 397, 406, 767]. Finite [78, 550, 44, 127, 138, 139, 140, 141, 191, 192, 197, 223, 224, 250, 265, 273, 327, 337, 372, 471, 481, 492, 508, 530, 531, 528, 529, 542, 546, 545, 635, 644, 704, 738, 745, 744, 759, 768, 774, 775, 791]. finite-difference [223, 224]. finite-element [644]. first [47, 52, 103, 264, 422, 486, 489, 588]. Fitting [634, 160, 178, 317, 385, 779]. fixed [18, 256, 281]. fixed-time [18]. flat [222, 415]. flexibility [393]. flexure [573]. Flow [93, 415, 416, 697, 10, 21, 169, 185, 187, 190, 209, 286, 359, 373, 405, 414, 519, 577, 576, 586, 593, 597, 602, 620, 632, 639, 641, 640, 677, 685, 696, 698, 716, 758, 773, 785]. flows [193, 234, 289, 323, 462, 583, 772]. Fluid [718, 21, 87, 207, 222, 368, 447, 502, 527]. fluid-loaded [447]. fluids [358]. Forbidden [546]. force [255]. forced [13, 143]. forces [396]. form [91, 89, 242, 312, 571, 647]. formal [515]. Forms [244, 142, 581]. formula [612, 655, 656, 710, 711]. formulae [123, 152, 153, 198, 199, 421, 455, 464, 467, 596, 646, 710, 711]. formulas [436, 451, 637, 668, 670]. formulation [54, 265, 407, 521, 785]. formulations [724]. Forsythe [317]. Fort [525]. Fort-Frankel [525]. forward [540]. four [370]. four-dimensional [370]. Fourier [394, 49, 126, 233, 350, 440, 491, 533, 566, 591, 612, 646, 764]. fourth [138, 139, 275, 763]. fourth-order [138, 139]. fractional [518, 566]. Fractionalization [567]. fractions [151, 241, 516, 558]. Frankel [525]. Fredholm [31, 47, 52, 216, 395, 486, 506, 614]. free [10, 66, 184, 190, 287, 304, 536, 659]. free-boundary [304]. frequencies [363]. frequency [37, 38, 39, 381, 445, 691, 751]. friction [21]. frictional [739]. front [569, 750]. fronts [737]. fully [135, 455, 635]. Function [793, 73, 74, 263, 295, 294, 405, 494, 534, 539, 557, 568, 658, 683, 756, 757, 794, 797]. functional [302, 490, 631]. functions [31, 61, 62, 126, 127, 253, 262, 264, 313, 346, 367, 368, 377, 381, 379, 389, 442, 530, 541, 582, 608, 674, 681, 692, 725, 761, 765, 790]. fundamental [806]. furnaces [348]. Further [224, 391, 505].

Galerkin

[47, 56, 217, 218, 284, 500, 594, 760].

Galerkin-Kantorovich [500]. games [400]. gap [212]. gaps [35]. Garside [248]. Gauss [436, 451]. Gaussian [63, 85, 99, 202, 214, 243, 466, 582, 643, 645]. General [344, 465, 483, 138, 139, 141, 145, 146, 166, 293, 385, 559, 691, 97]. generalization [79, 333, 525, 689]. Generalized [193, 259, 434, 749, 49, 253, 305, 388, 389, 430, 504, 505, 533, 603, 608, 662, 805]. generated [241]. Generating [798, 126]. Generation [672, 412]. generator [754]. generators [753]. geophysics [428]. given [277, 773]. Givens [326, 374]. Gladwell [336]. global [220, 221, 648, 796]. Goursat [214]. governed [158, 157, 390, 789]. governing [389]. Gradient [429, 46, 231]. graph [574]. gravity [239, 289, 495]. Green [197]. Group [487, 684]. growth [360, 701].

Hadamard [197, 806]. half

[238, 255, 441, 449, 601, 739, 788]. half-plane [238, 441]. half-space [255, 449, 739]. half-spaces [788]. Hall [209]. hanging [615]. Hankel [567, 669]. harmonic [60, 150, 436, 484, 676]. HEAPS [110]. heat [176, 185, 209, 226, 227, 287, 300, 301, 322, 348, 373, 420, 421, 459, 606, 789]. heaving [213]. Heaviside [680]. heavy [156]. hemispherical [213]. Hermite [283]. heterogeneous [154]. High [140, 198, 420, 526, 529, 637, 285, 363, 471, 528, 751, 755, 381]. High-accuracy [140]. High-order [198, 528]. Higher [421, 438, 459]. highest [646]. highly [108]. Hilbert [439, 631]. Hill [555, 743]. hole [731]. homogeneous [128, 377]. Homotopy [619]. Hopf [394]. hopscotch [320, 344, 595]. horizontal [87]. hosepipes [356]. Householder [468]. Hurwitz [54]. hydromagnetic [209]. hyperbolic [86, 155, 273, 347, 521, 522, 524, 526, 741, 806]. hypercircle [247]. hypergeometric [389, 725].

I. [336, 418]. ideal [270, 294, 731]. ignition [306]. II [36, 38, 84, 98, 146, 251, 268, 291, 310, 379, 403, 480, 577, 625, 757, 800]. III [39, 94, 308]. ill [248]. ill-conditioned [248]. image [171]. imbedding [782]. immersed [357]. impact [80]. implementation [199, 221, 411]. implicit [121, 120, 123, 124, 199, 220, 273, 345, 370, 371, 382, 635]. implosions [137]. Improved [406, 226, 508]. impulsively [222, 415]. inaccurate [541]. incidence [357]. incident [281]. include [515]. inclusion [130, 580]. inclusions [705]. incompressible [129, 500]. Inconsistencies [584]. inconsistent [256, 397]. Increased [15, 616]. incrementation [210]. Indefinite [246, 318]. indentation [580]. independent [184]. induced [94, 788]. inequality [65, 265]. inextensible [736]. infinite [6, 60, 147, 212, 222, 390, 415, 576, 617, 690]. Inflation [129]. Inflation-extension [129]. Influence [429]. Inherent [588]. inhomogeneous [9, 783]. initial [128, 192, 199, 259, 470, 559, 560, 758, 804]. initial- [560]. initial-boundary [559]. initial-value [199, 259, 804]. inspection [622]. Inst [561, 711]. instability [502, 588, 677]. Integral [150, 785, 2, 1, 31, 33, 52, 53, 58, 103, 104, 112, 114, 132, 149, 216, 218, 262, 263, 264, 351, 365, 366, 394, 463, 464, 486, 506, 509, 655, 695, 765, 777, 787, 784, 797]. Integrals [581, 1, 4, 62, 108, 131, 318, 467, 491, 511, 518, 578, 669, 690]. Integration [710, 711, 109, 119, 152, 153, 338, 452, 582, 650, 761]. integro [489, 527, 548]. integro-differential [489, 527, 548]. intense [117]. intensity [43]. Interaction [269, 402, 758]. interchanging [184]. interface [502]. Interfacial [788]. interference [298]. interpolating [674]. Interpolation [180, 471, 673, 67, 161, 307, 352, 406, 436, 458, 793]. interpolatory [68, 457]. intersection [277, 447]. intersymbol [298]. Interval [30, 316, 569]. Invariant [782]. invariants [151]. inverse [4, 406, 458]. inverses [747]. Inversion [240, 25, 126, 364, 493, 613, 657, 721, 730, 735, 784]. invert [656]. investigation [17, 149, 206]. investigations [267, 268]. involving [27, 114, 152, 153, 236]. irreducible [142, 377, 547]. irregularly [307]. isentropic [716]. isoparametric [759]. isotropic [129, 159]. Iteration [162, 210, 432]. Iterative [26, 603, 119, 163, 274, 343, 433, 506, 605, 619].

J [561, 711]. J. [336, 494]. Jacobi [36, 92, 411, 784]. Jacobi-type [411]. Jacobian [203, 200]. Jarratt [248]. joined [788]. Joseph [350]. jumps [398]. junction [18].

Kantorovich [500]. Kaplun [537]. kernel [485, 658]. kernels [395, 614]. kind [47, 52, 103, 104, 263, 486, 489, 506]. kinds [264]. Kirkwood [58]. knots [14, 647]. Knudsen [169]. Kontorovich [446]. Kutta [422, 423, 121, 120, 123, 124, 198, 199, 259, 382, 629, 648].

L [336]. L.O.D. [345]. Lagrangian [547]. Laguerre [608]. lake [213]. Lamé [667, 723]. laminar [21]. Lanczos [170, 258, 460, 599, 600]. Laplace [4, 253, 364, 370, 417, 493, 613, 644, 656, 657, 721, 728]. Large [201, 86, 108, 258, 369, 397, 469, 473]. Lax [521]. layer [196, 206, 229, 405, 500, 696, 701]. layers [87, 269, 620, 632, 713]. Least [326, 801, 219, 296, 385, 404, 468, 538, 795]. least-squares [296, 385, 404]. Lebedev [446]. length [339]. lengths [203]. level [459]. Liapunov [649]. lifting [414]. like [136]. Limit [535]. limitations [198]. Limits [669, 537]. Linear [201, 246, 300, 301, 316, 651, 3, 12, 20, 33, 32, 42, 59, 64, 86, 105, 125, 134, 138, 139, 141, 147, 155, 159, 163, 192, 194, 217, 219, 225, 230, 231, 241, 245, 256, 274, 303, 325, 333, 335, 351, 365, 383, 386, 398, 404, 425, 430, 437, 448, 454, 461, 465, 473, 476, 477, 482, 501, 513, 514, 521, 522, 527, 538, 547, 548, 559, 560, 568, 593, 587, 588, 594, 607, 627, 628, 649, 659, 663, 671, 694, 699, 705, 726, 727, 735, 739, 750, 751, 754, 762, 766, 770, 771, 776, 779, 803, 804]. linearization [734]. linearly [570]. lines [536]. lining [577, 576]. Liouville [683]. Lipschitz [669]. Lipschitz-Hankel [669]. Liquid [286]. Lloyd [336]. loaded [447, 449]. loading [571, 788]. Local [677, 101]. Locally [307, 345, 553]. location [543]. locomotive [638]. Long [497]. Longitudinal [197]. loss [752]. Low [37, 38, 39, 305, 445, 691]. low-frequency [691]. lower [21, 295, 478, 631]. Lyapunov [55].

M. [427]. M.H.D. [654]. machine [260, 266]. machining [265, 290, 291, 297]. Mack [248]. made [150]. magnetic [415, 685, 713]. magnetogasdynamic [187]. magnetohydrodynamic [269, 583, 688, 698]. magnetohydrodynamics [671, 781]. maintenance [622]. make [74]. making [73]. many [295, 511, 522]. many-fermion [511]. mapping [465, 485]. Markov [558]. mass [448, 752]. Matched [752]. matching [220, 531]. material [129, 270]. Materials [736, 197, 472]. Math [561, 711]. Mathematical [290, 291, 350, 362, 766]. Mathematics [488, 579]. Matrices [202, 244, 257, 91, 89, 107, 142, 173, 203, 200, 240, 258, 274, 330, 331, 371, 392, 432, 434, 430, 610, 747]. Matrix [55, 246, 799, 800, 24, 53, 64, 433, 458, 534, 600, 662]. matter [342]. Maximum [34, 190]. Maxwell [125, 340, 341, 342]. mean [577, 576]. measure [802]. mechanics [368, 566]. media [286, 727]. medium [9, 329, 786]. Mellin [730]. melting [44, 737]. memory [794]. mesh [16, 542]. metabolism [282]. metallic [288]. meteorological [191]. Method [162, 346, 429, 591, 5, 12, 29, 46, 47, 64, 66, 69, 72, 79, 81, 82, 90, 96, 102, 113, 138, 139, 141, 158, 157, 168, 170, 171, 182, 185, 191, 192, 194, 214, 215, 216, 219, 221, 226, 240, 248, 256, 258, 265, 276, 285, 304, 317, 382, 386, 408, 411, 440, 460, 461, 475, 485, 489, 496, 506, 508, 521, 527, 530, 531, 536, 540, 546, 545, 551, 580, 589, 599, 605, 613, 638, 639, 641, 640, 649, 671, 714, 733, 767, 770, 794, 797, 806]. Methods [92, 33, 56, 101, 127, 133, 135, 163, 217, 224, 228, 245, 273, 274, 284, 310, 308, 314, 315, 332, 336, 343, 345, 370, 383, 406, 407, 410, 418, 422, 423, 459, 470, 523, 524, 526, 529, 548, 553, 578, 585, 586, 590, 619, 648, 660, 683, 690, 703, 755, 760, 764, 790, 795]. metric [407, 442, 623]. mildly [448]. mimic [368]. mine [400]. minimal [232, 452]. minimax [501, 771]. Minimization [73, 295, 10, 74, 75, 90, 97, 98, 349, 624, 625, 703, 793, 794]. Minimizing [568, 18]. minimum [34, 773]. Mixed [514, 138, 182, 194, 229, 523, 524, 676]. mixtures [41]. model [236, 287, 362, 519, 633]. models [348, 715, 718, 766]. Moderate [499]. modes [570, 576]. modification [64, 468, 806]. Modifications [248, 374]. Modified [538, 553, 591, 46, 122, 540]. modular [151]. moduli [354]. moments [12, 456, 790]. monomer [652]. monotone [135, 189]. Monte [626]. mooring [13]. motion [13, 156, 177, 222, 232, 783]. motions [40]. move [569]. move-to-the-front [569]. moving [23, 42, 171, 186, 185, 188, 265, 315, 378, 542, 748]. multi [50, 572, 627]. multi-dimensional [50]. multi-dimensions [627]. multi-plant [572]. multiderivative [550]. multidimensional [536]. multilevel [522]. multiphase [148]. Multiple [779, 29, 131, 401]. multisoliton [125]. multistep [470, 521, 527, 550, 791]. Murphy [494]. muscle [403].

natural [372, 640]. near [93, 441, 696]. nearby [175]. Nearest [503]. necessary [763]. negative [399]. neighbour [503]. network [118, 544]. Neumann [38, 513, 584]. Newton [66, 101, 332, 383, 670, 703, 794]. Node [528]. Non [473, 476, 750, 751, 20, 33, 32, 46, 59, 73, 74, 86, 89, 105, 125, 128, 138, 139, 141, 189, 192, 235, 300, 313, 314, 340, 341, 365, 383, 386, 399, 404, 425, 442, 448, 461, 465, 501, 513, 514, 521, 522, 538, 548, 568, 570, 593, 587, 607, 627, 628, 649, 659, 663, 727, 766, 770, 771]. non- [138]. non-autonomous [649]. non-definite [442]. non-homogeneous [128]. Non-linear [473, 476, 750, 751, 20, 33, 32, 59, 86, 105, 125, 139, 141, 192, 300, 365, 383, 386, 404, 425, 448, 461, 465, 501, 513, 514, 521, 522, 538, 548, 568, 593, 587, 607, 627, 628, 659, 663, 727, 766, 770, 771]. non-linearly [570]. non-monotone [189]. non-negative [399]. non-parametric [313, 314]. non-quadratic [73, 74]. non-selfadjoint [235]. non-symmetric [46, 89]. non-zero [340, 341]. Nonconservative [325, 146]. nonequilibrium [236]. nonlinear [106, 135, 140, 301, 336, 496, 661, 724]. nonpotential [724]. Note [174, 374, 643, 787, 74, 76, 120, 122, 132, 164, 225, 234, 278, 355, 447, 452, 467, 498, 544, 559, 560, 575, 595, 598, 632, 684, 686, 698, 703]. Notes [45]. null [341]. Number [99, 21, 390, 574, 791]. numbers [36, 151]. Numerical [66, 95, 152, 179, 181, 184, 267, 268, 276, 359, 417, 418, 463, 491, 609, 650, 695, 707, 708, 709, 730, 776, 777, 790, 1, 5, 10, 12, 14, 17, 45, 81, 82, 100, 108, 109, 116, 153, 158, 157, 164, 206, 214, 217, 228, 225, 275, 277, 315, 323, 336, 338, 344, 348, 361, 369, 380, 410, 419, 421, 435, 460, 461, 485, 486, 493, 543, 563, 596, 604, 613, 637, 649, 676, 721, 729, 745, 755, 764, 765, 804].

objective [73, 74]. Oblate [39]. oblique [357]. obliquely [281]. obstacles [60]. obtain [478]. Obtaining [243, 806]. occurring [131, 744]. occurs [458]. odd [14, 16, 458]. odd-degree [14]. off [309, 637]. off-step [637]. One [769, 3, 12, 91, 234, 328, 345, 348, 553, 606, 635]. one-dimensional [12, 328, 348, 635]. One-sided [769, 3]. only [295]. operational [680]. operator [769]. operators [235, 399, 699, 724, 765]. optics [125]. Optimal [40, 475, 541, 51, 166, 211, 260, 337, 424, 638, 706, 763]. optimality [746]. optimization [110, 294, 332, 437, 674, 746]. optimized [521]. Optimum [371, 422, 423, 688]. order [16, 65, 86, 107, 124, 138, 139, 141, 158, 164, 198, 208, 320, 335, 422, 423, 421, 458, 459, 483, 528, 529, 566, 588, 630, 637, 649, 719, 741, 755, 767, 7]. ordered [602]. Orderings [243, 605]. orders [438, 522]. ordinary [119, 335, 334, 338, 380, 637, 650, 694, 755, 776]. ordinate [297]. ordinates [405]. origin [253]. original [493]. orthogonal [50, 173]. orthogonality [509]. orthogonalization [292]. oscillating [279]. Oscillation [719]. oscillations [7, 130, 143, 278]. oscillator [211, 424]. oscillatory [108, 145, 146, 607, 690]. Oseen [785]. other [375]. over- [276]. over-relaxation [72]. overdetermined [219, 245, 501, 771]. overstable [207]. oxygen [186, 282].

P [248]. Padé [57, 58, 351, 483, 515, 658, 742]. Padé-approximant [58]. Padé-type [57, 483]. pair [149]. Pairs [22]. paper [494]. parabolic [12, 66, 228, 225, 261, 273, 275, 320, 382, 448, 460, 461, 475, 523, 531, 550, 553, 678, 692, 806]. parallel [175, 700]. parameter [210, 306, 380, 408, 459, 465, 475]. parameters [703]. parametric [313, 314]. Partial [431, 34, 65, 223, 224, 228, 225, 247, 261, 273, 275, 344, 432, 460, 461, 553, 678, 741]. partially [278, 286]. particle [358]. particles [527]. past [187, 222, 414, 415, 416, 685]. path [48]. path-finding [48]. Pelican [362]. penalty [294]. pencils [330, 331]. pendulum [143]. perfectly [507, 664, 665, 778]. perfectly-plastic [507]. periodic [7, 16, 112, 454, 458, 463, 470, 552, 687]. peripheral [72, 276]. Permutation [142]. permutations [242]. Perturbation [743, 113, 368, 593, 714]. perturbations [271, 679]. perturbative [552]. phase [174, 236, 440, 683, 737]. phenomena [654]. phenomenon [267, 268]. physics [350]. Piecewise [104, 68, 103, 178, 384]. piecewise-cubic [68]. pipe [698]. Pivotal [243]. Pivoting [735]. planar [737]. Plane [374, 414, 60, 93, 150, 157, 175, 238, 237, 270, 288, 380, 402, 403, 426, 441, 507, 551, 586, 639, 641, 726]. plane-wave [288]. planetary [358]. planetary-type [358]. plant [572]. plasma [509]. plastic [154, 507, 778]. plate [6, 222, 278, 415, 416, 444, 447, 731, 740, 744]. point [15, 138, 139, 140, 141, 205, 250, 256, 277, 364, 365, 366, 515, 637, 708, 745, 748]. points [241, 307, 390, 389, 516, 539, 590]. pointwise [65]. Poisson [612, 634]. Pol [424]. pole [547]. pollution [5]. Polygonal [732, 410]. Polynomial [611, 14, 102, 104, 248, 250, 296, 317, 334, 349, 412, 549, 608, 619]. Polynomials [30, 312, 68, 103, 166, 178, 185, 504, 505, 543, 784]. population [360]. porous [190, 286]. positive [91, 274]. Positivity [36]. possibility [540]. Possible [319, 11]. posteriori [335, 477]. potential [427, 477, 682]. power [182, 401, 412, 483, 516]. Practical [352, 611, 575, 589, 800]. pre [274]. pre-conditioning [274]. precision [646]. prediction [532, 569]. predictor [5]. predictor-corrector [5]. prescribed [398]. presence [212, 342, 371, 415]. present [666]. presentation [747]. pressure [342, 396, 449]. pressurized [688]. principal [467, 578, 629]. principle [174, 190, 220, 561, 562, 775]. principles [22, 19, 20, 34, 57, 169, 234, 247, 299, 399, 398, 450, 545, 560, 579, 698, 699, 762, 798]. Probability [428, 126, 790]. probe [519]. Problem [162, 10, 27, 37, 38, 66, 81, 82, 95, 106, 105, 116, 149, 171, 176, 186, 184, 189, 192, 196, 214, 229, 260, 265, 266, 270, 275, 280, 282, 284, 287, 357, 361, 373, 378, 380, 391, 393, 417, 418, 419, 437, 484, 513, 520, 559, 584, 592, 593, 602, 603, 617, 618, 659, 666, 678, 707, 708, 722, 731, 732, 739, 740, 745, 752, 776, 781, 789, 805]. Problems [611, 2, 12, 15, 32, 42, 44, 48, 53, 56, 59, 69, 113, 114, 118, 134, 135, 138, 139, 140, 141, 150, 158, 157, 165, 182, 185, 183, 188, 191, 199, 205, 259, 276, 297, 299, 304, 315, 336, 334, 380, 383, 386, 388, 398, 434, 430, 465, 470, 514, 536, 551, 560, 563, 564, 580, 585, 594, 604, 631, 676, 682, 699, 702, 709, 717, 723, 726, 737, 755, 760, 762, 780, 782, 785, 804]. procedure [465, 594, 628]. procedures [25, 121, 120, 320, 343]. Process [99, 634, 689]. processes [23, 100, 236, 259, 433, 512, 558, 629]. produced [255, 278]. product [21, 91, 89, 455]. products [250]. programming [75, 266, 293, 333, 437, 572, 735, 803]. progressing [8]. prolate [37, 38, 693]. proof [190, 678, 738]. Proofs [625]. Propagation [329, 363, 6, 83, 84, 100, 272, 328, 390, 388, 453, 577, 576, 727, 750, 798]. Properties [389, 655, 720, 73, 74, 172, 173, 215, 558, 614, 624, 625, 653, 689, 719, 742, 754]. protective [621, 622]. pulse [298].

Quadratic [624, 625, 24, 29, 73, 74, 75, 231, 293, 295, 327, 442, 617]. Quadrature [464, 214, 387, 451, 455, 463, 467, 499, 589, 596, 646, 655, 656, 657, 668]. quadratures [466, 645]. qualitative [504, 505, 83]. Quantitative [84]. quantities [108]. quantum [566]. Quasi [332, 383, 101, 225, 454, 703, 794, 670]. quasi-linear [225, 454]. Quasi-Newton [332, 383, 101, 703, 794, 670]. queueing [469]. quick [752]. quotients [515].

Radiation [783, 748, 445, 481, 556]. radiative [19, 348, 413, 609]. random [328, 329, 539, 570, 672, 702]. randomly [321]. Range [319]. Rank [91, 97, 98]. Rank-one [91]. rank-two [91]. rapid [549]. Rate [429, 50, 533]. rates [169, 310, 308]. Ratio [136, 497]. Ratio-like [136]. rational [61, 62, 240, 241, 353, 352, 364, 365, 366, 401, 412, 483, 494, 557, 636, 756, 757]. ratios [252]. ray [88, 168]. Rayleigh [336, 456, 594, 639, 781]. reacting [577, 576]. reaction [32, 448, 633, 717]. reactors [593]. Real [30, 432]. rectangle [471]. rectangular [190, 434, 731]. recurrence [136]. recurrent [714]. recursive [75]. reduced [125]. reducing [626]. Reduction [244, 780]. reference [627]. reflection [700]. refraction [88]. region [449, 788]. regions [436, 555]. Regula [304]. Regula-Falsi [304]. regular [368, 48]. regularization [2, 1]. regularized [397]. regulator [134]. Reinforced [356, 270, 355, 554, 573, 731, 736]. reinforcement [240]. related [166]. relation [135, 136, 277]. relations [112, 316, 495, 509]. relative [375]. relaxation [72, 215, 276, 801]. reliable [352, 610]. rely [621]. Remainders [153, 518]. Remarks [401, 711, 734, 439, 726]. renewal [572]. replacement [569, 622]. replacing [540]. reply [494]. Representation [181]. representations [223, 224, 547, 647, 660, 774]. requirements [528]. reservoir [27]. residuals [317]. Resonance [482, 145, 146]. response [80, 403]. Restoring [522]. restrictions [398]. result [333]. resultant [662, 749]. results [23, 289, 504, 505, 624]. Review [766, 314]. revolution [350]. Reynolds [21]. Riccati [694, 782]. Richardson [689]. Richtmyer [791]. rigid [17, 130, 507, 601]. Riseman [58]. Ritz [594, 217, 336, 373, 639]. road [18, 35]. Robin [418]. rocket [763]. rods [409]. roller [449, 601]. root [210, 406]. root-finding [406]. Roots [326]. rotating [207, 209, 359, 620, 632, 772]. Rotation [374, 175]. rotational [405]. Round [309]. Round-off [309]. rounded [792]. Rounding [85, 712]. routes [544, 735]. Routh [54]. routines [338]. routing [118]. rule [452]. rules [499]. rumours [204].

Runge [121, 120, 123, 124, 198, 199, 259, 382, 422, 423, 629, 648].

S.E.R [215]. S.O.R. [584]. safety [621, 622]. sampled [321]. sampling [76, 626]. satisfies [372]. saturated [286]. scalar [691]. Scaling [202, 703]. Scattering [9, 60, 19, 37, 38, 39, 413, 426, 443, 456, 497, 691]. scheduling [260, 266]. scheme [44, 46, 525, 569, 616, 772]. schemes [122, 223, 369, 371, 420, 490, 522, 550, 621, 622, 661, 791]. Schrödinger [663]. screen [8, 17, 748]. sea [5]. search [569]. searches [230, 231, 503]. second [65, 86, 104, 158, 164, 208, 263, 264, 320, 335, 423, 506, 630, 649, 719, 741, 767, 7]. second-order [158, 335, 423, 630, 7]. sector [380]. segments [384]. selected [539]. selection [393]. Self [94, 193]. Self-induced [94]. self-similar [193]. selfadjoint [235]. semi [222, 415, 720]. semi-discrete [720]. semi-infinite [222, 415]. semiconductors [766]. semidiscrete [719]. semigroups [753]. sensitivity [571]. separated [498, 694]. separation [94]. Sequence [243]. sequences [61]. sequencing [602]. Series [591, 49, 136, 182, 195, 196, 226, 227, 401, 412, 483, 516]. set [90, 112, 127]. Sets [246, 217, 219, 250, 258, 385, 469]. settings [18]. several [307]. shapes [546]. shaping [291]. sharing [570]. shear [289]. sheet [443]. shells [396, 609, 707]. shock [77, 269, 405, 592, 798]. shop [602]. Short [213, 322, 481, 693]. Short-time [322]. sided [3, 769]. sides [757]. sidewalls [207]. signals [18, 321, 607]. significant [107]. Signorini [740]. similar [193]. Similarity [244]. simple [143, 173, 191, 489]. simplex [96, 540, 628]. Simplified [273]. simply [485]. simulation [5]. Simultaneous [162, 432, 163, 587, 628]. single [18, 19, 148, 537, 667]. Singular [759, 2, 1, 132, 164, 218, 263, 271, 361, 371, 430, 463, 464, 509, 527, 593, 614, 678, 679, 708, 787, 784]. Singularities [728, 56, 69, 183, 303, 604, 709]. sixth [141]. sixth-order [141]. size [735]. sizes [375]. skeletal [393]. slab [507]. slender [449, 682]. slightly [219]. slow [175, 773]. Slowly [462, 195, 409]. small [278, 568, 598]. smooth [307, 541]. Smoothing [803, 290]. solar [718]. solid [60, 77, 80, 130, 156, 571]. solids [778]. Solution [201, 246, 297, 386, 394, 395, 430, 500, 611, 10, 12, 31, 45, 46, 66, 81, 82, 86, 95, 103, 114, 116, 125, 150, 158, 157, 168, 164, 163, 165, 171, 184, 188, 199, 214, 216, 217, 219, 226, 227, 228, 225, 245, 254, 256, 261, 263, 273, 275, 280, 277, 276, 282, 284, 334, 344, 347, 349, 365, 366, 369, 380, 397, 417, 418, 419, 421, 460, 461, 464, 465, 479, 480, 486, 501, 506, 520, 551, 563, 603, 604, 609, 612, 628, 644, 657, 676, 707, 708, 716, 728, 729, 745, 744, 760, 771, 776]. solutions [15, 52, 53, 58, 64, 65, 322, 335, 381, 454, 468, 552, 614, 633, 663, 667, 671, 695, 706, 705, 738, 743, 772, 777, 806]. Solving [370, 469, 113, 148, 150, 185, 248, 274, 489, 527, 549, 575, 638]. Some [11, 23, 55, 99, 172, 173, 233, 289, 313, 439, 455, 495, 504, 558, 612, 636, 653, 654, 682, 689, 726, 741, 764, 768, 795, 36, 42, 71, 75, 150, 164, 163, 245, 254, 314, 324, 353, 389, 479, 480, 490, 511, 513, 514, 560, 668, 744, 747, 766]. SOR [605]. sound [83, 84, 441, 443, 577, 576, 751]. source [187, 441]. space [227, 228, 255, 413, 419, 449, 521, 522, 523, 553, 601, 631, 657, 739, 791]. spaced [14, 307, 645, 647]. spaces [455, 675, 788]. spanning [115]. Sparse [201, 244, 246, 142, 200, 245, 258]. Sparsity [243, 735]. spatially [322]. special [367, 627, 650]. specialization [396]. Spectral [321]. Spectrum [429]. sphere [128, 175, 213, 701]. spheres [427, 498]. spherical [609]. spheroid [37, 38, 693]. spheroids [37, 38, 39]. spinning [701]. spirals [664, 665]. Spline [180, 181, 263, 339, 15, 31, 67, 111, 161, 262, 264, 385, 606, 635, 647, 765, 792]. Splines [179, 14, 16, 185, 205, 316, 318, 334, 457, 458]. splitter [444]. Square [326, 210]. Squares [326, 219, 296, 385, 404, 468, 538, 568, 795, 801]. Stability [396, 490, 555, 594, 630, 643, 661, 675, 791, 71, 177, 208, 300, 301, 309, 324, 338, 474, 535, 615, 649, 663, 737, 796]. Stabilizability [753]. stabilization [630]. Stable [25, 590, 660]. stages [128]. started [222, 415]. state [42, 229, 373, 657, 666, 754]. Statement [624]. static [704]. stationary [28, 174, 440]. statistical [171]. Steady [187, 42, 229, 373]. steady-state [42, 229]. steepest [79]. Stefan [189, 666, 737]. Stenger [668]. step [124, 203, 527, 548, 637]. step-by-step [548]. Stieltjes [517]. stiff [119, 259, 637, 653, 650, 755]. Stochastic [204, 674, 194, 512, 714, 780]. stochastic-deterministic [194]. stochastically [621]. Stokes [267, 268, 495]. storage [28, 27, 397]. storage-efficient [397]. storage-stationary [28]. strain [197, 270]. Strang [791]. strategy [387, 589]. stratified [323]. Stream [405, 556]. Stream-function [405]. streams [28, 27]. stress [43, 270, 472, 731, 778]. stress-deviator [778]. stresses [571, 788]. strip [288]. structure [77, 345, 437, 650, 758]. structures [393]. studies [359, 358]. study [315, 361, 593]. Sturm [54, 683]. sub [414]. sub-critical [414]. subject [295]. subjected [396]. submerged [376]. subsonic [556, 586, 639]. successive [72, 215, 772, 801]. suddenly [255, 416]. Sufficiency [327]. suitable [563]. sum [568]. Summability [608]. superlinear [101]. supersonic [758]. surd [4]. surface [10, 156, 160, 213, 239, 255, 281, 289, 357, 363, 426, 449, 481, 561, 562, 571, 627, 654, 700]. surfaces [385, 449]. swirling [358]. switchings [337]. Symmetric [246, 257, 470, 24, 46, 89, 107, 155, 183, 237, 258, 274, 377, 395, 455, 499, 554, 600, 610, 670]. symmetrically [697]. system [232, 337, 397, 448, 523, 638, 654, 771, 789]. Systems [85, 201, 651, 7, 46, 51, 64, 119, 145, 146, 147, 155, 158, 157, 194, 245, 264, 308, 324, 325, 327, 347, 454, 469, 482, 501, 511, 521, 522, 547, 563, 564, 565, 588, 617, 618, 630, 649, 650, 652, 684, 694, 767, 804, 310].

table [515, 793]. tactics [400]. tardiness [602]. tau [170]. tau-method [170]. Tchebycheff [256]. technique [2, 1, 108, 163, 165, 292, 311, 349, 465, 500, 508, 575, 606, 626, 635, 673]. techniques [521]. temperature [305, 427]. temporal [135]. tension [354]. tensors [377]. term [719]. termination [231, 624, 625]. terms [375, 517, 681, 777]. test [196, 676]. Tests [746, 136]. textures [672]. their [198, 316, 742, 791]. theorem [324, 537, 749]. theorems [54, 474, 478, 625]. Theoretical [799]. theories [41]. theory [11, 23, 42, 83, 84, 88, 131, 145, 146, 290, 291, 306, 400, 428, 495, 504, 505, 509, 592, 607, 627, 704, 744, 750]. Thermal [87, 306, 737]. thermodynamic [236]. thermoelastic [229]. thermoelasticity [305]. thin [396]. third [417, 419]. Three [601, 228, 260, 266, 413, 462, 578, 687, 798]. Three-dimensional [601, 413]. three-machine [260, 266]. Time [211, 424, 18, 23, 51, 69, 70, 86, 134, 282, 297, 322, 473, 617, 675, 760]. time-dependent [69, 282, 675]. times [327]. timetabling [520]. timing [328]. tissue [186]. tissues [282]. tolerances [803]. topology [393]. tori [682]. Torsion [354, 154]. Torsional [130]. torus [113, 497]. total [602]. touching [427]. tournament [510]. tracking [618]. traffic [18, 35]. trajectory [763]. transfer [19, 209, 348, 413, 448, 609]. transform [4, 126, 233, 364, 446, 493, 566, 613, 657, 730, 764, 394]. transformation [159, 297, 410, 529, 540]. Transformations [115, 244, 326, 144, 604, 694, 782]. transforms [253, 439, 440, 517, 567, 612, 656, 721]. transient [571]. transition [390, 389]. transmission [388, 598]. transonic [716]. transport [46, 413]. transverse [197]. trapezoidal [710, 711]. traps [597]. travelling [633]. traversing [601]. treating [69, 182]. treatment [183, 348, 709]. trees [115]. triangle [499]. triangular [142, 242]. tridiagonal [107, 138, 139, 140, 141, 610]. tridiagonalizing [600]. trigonometric [681]. truncation [147]. tube [129, 348, 354, 573, 615]. tubes [355, 697]. turbulent [527]. TWK [602]. Two [271, 328, 457, 574, 676, 15, 28, 27, 60, 91, 114, 133, 138, 139, 140, 141, 205, 227, 228, 282, 299, 353, 364, 365, 366, 380, 402, 412, 426, 427, 459, 498, 510, 515, 516, 521, 523, 553, 673, 700, 708, 728, 737, 745, 752, 405]. two-body [752]. two-dimensional [60, 114, 282, 459, 405]. two-level [459]. two-parameter [380, 459]. two-phase [737]. two-point [15, 138, 139, 140, 141, 205, 364, 365, 366, 708]. Two-timing [328]. Two-variable [271]. type [4, 57, 104, 310, 358, 366, 398, 464, 477, 483, 602, 637, 791, 411]. types [747].

ultimate [778]. ultrasound [519]. ultraspherical [375]. un-damped [774]. unbounded [617, 618]. unconstrained [90, 332, 703, 794]. underground [117]. underwater [83, 84]. unified [670]. Uniform [381, 379, 585, 586, 725, 16, 49, 626]. uniformly [513]. unique [150]. uniqueness [453]. unloading [739]. Unsteady [583, 196, 234, 620, 632]. Unsymmetric [162, 201, 399, 308]. update [670]. Upper [21, 142, 295, 478]. upwind [46]. Use [30, 258, 317, 46, 73, 74, 233, 274, 478, 492, 531, 604, 723, 803]. useful [150]. uses [198].

Using [419, 31, 185, 190, 203, 205, 227, 253, 364, 365, 366, 370, 382, 394, 460, 509, 612, 635, 640, 765, 790].

vacuum [597]. value [15, 95, 135, 138, 139, 140, 141, 158, 176, 182, 192, 199, 205, 229, 259, 270, 275, 336, 334, 361, 417, 419, 465, 467, 470, 514, 559, 560, 563, 564, 578, 594, 604, 676, 707, 708, 726, 731, 745, 804]. Values [319]. variable [70, 95, 124, 148, 271, 407, 442, 526, 527, 623]. variables [184, 227, 228, 295, 307, 353, 367, 412, 522, 553, 791]. variance [626]. variant [63]. variants [599]. variation [496]. Variational [19, 134, 398, 450, 545, 667, 717, 22, 20, 21, 33, 57, 116, 133, 169, 234, 265, 299, 310, 308, 309, 532, 551, 560, 561, 562, 580, 585, 586, 641, 640, 724, 775]. varying [322, 409, 462]. velocity [487, 684]. vertex [237]. Vertical [502, 281, 278, 416, 426, 700]. very [369]. via [66, 500]. Vibration [376, 570]. visco [519]. visco-elastic [519]. viscoelastic [601]. viscoelasticity [42]. viscous [21, 222, 323, 773]. Volterra [1, 103, 104, 262, 263, 264, 365, 366, 489, 777]. Volterra-type [104]. volume [773]. vortex [88, 443]. vorticity [686].

wake [206]. walk [702]. wall [175, 696]. Water [687, 280, 281, 279, 598, 627]. Wave [6, 390, 727, 83, 84, 272, 279, 285, 288, 328, 381, 379, 633, 635, 654, 684, 750]. wave-energy [279]. wave-front [750]. wavelength [497]. wavelengths [693]. Waves [212, 8, 60, 112, 197, 213, 239, 269, 280, 281, 278, 289, 357, 363, 409, 426, 476, 481, 495, 561, 562, 592, 598, 627, 687, 700, 751]. Weak [137, 188, 269]. weakly [145, 146, 565, 614]. wedge [237, 642]. weight [582]. Weighted [296, 626, 628, 802]. weights [452]. Wendroff [521]. which [24, 74, 458, 516, 621]. Whittaker [743]. Whittaker-Hill [743]. Whose [246].

Wiener [394]. wind [718]. window [376]. wing [93, 722, 758]. wing-body [758]. Without [326, 90, 134, 150, 230, 231].

zero [340, 341]. Zeros [30, 346, 543]. Zhabotinskii [633].

References

Abd-Elal:1979:AER

[1] L. F. Abd-Elal. An asymptotic expansion for a regularization technique for numerical singular integrals and its application to Volterra integral equations. Journal of the Institute of Mathematics and its Applications, 23(3):291–309, 1979. CODEN JMTAA8. ISSN 0020-2932.

Abd-Elal:1976:RTC

[2] L. F. Abd-Elal and L. M. Delves. A regularization technique for a class of singular integral problems. Journal of the Institute of Mathematics and its Applications, 18(1):37–47, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Abdelmalek:1976:DLO

[3] N. N. Abdelmalek. The discrete linear one-sided Chebyshev approximation. Journal of the Institute of Mathematics and its Applications, 18(3):361–370, 1976. CODEN JMTAA8. ISSN 0020-2932.

Aboul-Atta:1978:ASI

[4] O. A. Aboul-Atta and E. Tomchuk. Approximations to surd integrals of the inverse Laplace transform type. Journal of the Institute of Mathematics and its Applications, 21(??):157–163, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Adam:1977:PCM

[5] Y. Adam. A predictor-corrector method for the numerical simulation of sea pollution. Journal of the Institute of Mathematics and its Applications, 19(3):295– 305, 1977. CODEN JMTAA8. ISSN 0020-2932.

Advani:1969:WPI

[6] S. H. Advani. Wave propagation in an infinite plate. Journal of the Institute of Mathematics and its Applications, 5(??): 271–282, 1969. CODEN JMTAA8. ISSN 0020-2932.

Aggarwal:1967:ABP

[7] J. K. Aggarwal. Amplitude bounds on periodic and aperiodic oscillations: Second-order systems. Journal of the Institute of Mathematics and its Applications, 3(??):202–215, 1967. CODEN JM-TAA8. ISSN 0020-2932.

Ahluwalia:1969:DPW

[8] Daljit S. Ahluwalia and Robert M. Lewis. Diffraction of progressing waves by a screen. Journal of the Institute of Mathematics and its Applications, 5(??): 113–139, 1969. CODEN JMTAA8. ISSN 0020-2932.

Ahner:1977:SIM

[9] J. F. Ahner. Scattering by an inhomogeneous medium. Journal of the Institute of Mathematics and its Applications, 19 (4):425–439, 1977. CODEN JMTAA8. ISSN 0020-2932.

Aitchison:1977:NSM

[10] J. M. Aitchison. The numerical solution of a minimization problem associated with a free surface flow. Journal of the Institute of Mathematics and its Applications, 20(1):33–44, 1977. CODEN JMTAA8. ISSN 0020-2932.

Albasiny:1966:SBP

[11] E. L. Albasiny and W. A. Day. Some best-possible error bounds in deformation theory. Journal of the Institute of Mathematics and its Applications, 2(??): 266–272, 1966. CODEN JMTAA8. ISSN 0020-2932.

Albasiny:1968:NSL

[12] E. L. Albasiny and W. A. Day. The numerical solution of linear onedimensional parabolic problems by the method of moments. Journal of the Institute of Mathematics and its Applications, 4(??):140–162, 1968. CODEN JM-TAA8. ISSN 0020-2932.

Albasiny:1969:FME

[13] E. L. Albasiny and W. A. Day. The forced motion of an extensible mooring cable. Journal of the Institute of Mathematics and its Applications, 5(??):55–71, 1969. CODEN JMTAA8. ISSN 0020-2932.

Albasiny:1971:NCO

[14] E. L. Albasiny and W. D. Hoskins. The numerical calculation of odd-degree polynomial splines with equi-spaced knots. Journal of the Institute of Mathematics and its Applications, 7(??):384– 397, 1971. CODEN JMTAA8. ISSN 0020-2932.

Albasiny:1972:IAC

[15] E. L. Albasiny and W. D. Hoskins. Increased accuracy cubic spline solutions to two-point boundary value problems. Journal of the Institute of Mathematics and its Applications, 9(??):47–55, 1972. CODEN JMTAA8. ISSN 0020-2932.

Albasiny:1973:EEB

[16] E. L. Albasiny and W. D. Hoskins. Explicit error bounds for periodic splines of odd order on a uniform mesh. Journal of the Institute of Mathematics and its Applications, 12(??):303–318, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Allami:1975:NIA

[17] K. Allami. A numerical investigation of acoustic diffraction by a circular aperture in a rigid screen. Journal of the Institute of Mathematics and its Applications, 16(3):371–380, 1975. CODEN JMTAA8. ISSN 0020-2932.

Allsop:1971:DMS

[18] Richard E. Allsop. Delay-minimizing settings for fixed-time traffic signals at a single road junction. Journal of the Institute of Mathematics and its Applications, 8(??):164–185, 1971. CODEN JMTAA8. ISSN 0020-2932.

Anderson:1973:VPR

[19] D. Anderson. Variational principles for the radiative transfer equation of single scattering. Journal of the Institute of Mathematics and its Applications, 12 (??):55–62, 1973. CODEN JMTAA8. ISSN 0020-2932.

Anderson:1974:CVP

[20] N. Anderson and A. M. Arthurs. Complementary variational principles for a class of non-linear diffusion equations. Journal of the Institute of Mathematics and its Applications, 13(??):153–159, 1974. CODEN JMTAA8. ISSN 0020-2932.

Anderson:1974:ULV

[21] N. Anderson and A. M. Arthurs. Upper and lower variational bounds for the friction factor Reynolds number product in laminar flow of a viscous fluid. Journal of the Institute of Mathematics and its Applications, 14(??):69–74, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Anderson:1969:PCV

[22] N. Anderson, A. M. Arthurs, and P. D. Robinson. Pairs of complementary variational principles. Journal of the Institute of Mathematics and its Applications, 5(4):422–431, 1969. CODEN JM-TAA8. ISSN 0020-2932.

Anderson:1976:SET

[23] Oliver D. Anderson. Some eigen theory results for moving average time processes. Journal of the Institute of Mathematics and its Applications, 18(3):325– 330, 1976. CODEN JMTAA8. ISSN 0020-2932.

Anderson:1973:QCA

[24] P. Anderson and G. Loizou. On the quadratic convergence of an algorithm which diagonalizes a complex symmetric matrix. Journal of the Institute of Mathematics and its Applications, 12 (??):261–271, 1973. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Anderssen:1976:SPI

[25] R. S. Anderssen. Stable procedures for the inversion of Abel’s equation. Journal of the Institute of Mathematics and its Applications, 17(3):329–342, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Andrew:1979:ICD

[26] A. L. Andrew. Iterative computation of derivatives of eigenvalues and eigenvectors. Journal of the Institute of Mathematics and its Applications, 24(2):209– 218, 1979. CODEN JMTAA8. ISSN 0020-2932.

Anis:1970:CRS

[27] A. A. Anis and S. S. Daoud. A continuous reservoir storage problem involving two streams. Journal of the Institute of Mathematics and its Applications, 6(??): 241–249, 1970. CODEN JMTAA8. ISSN 0020-2932.

Anis:1968:SSD

[28] A. A. Anis and A. S. T. El-Naggar. The storage-stationary distribution in the case of two streams. Journal of the Institute of Mathematics and its Applications, 4(??):223–231, 1968. CODEN JMTAA8. ISSN 0020-2932.

Arthur:1972:EBM

[29] D. W. Arthur. Extension of Bairstow’s method for multiple quadratic factors. Journal of the Institute of Mathematics and its Applications, 9(??):194–197, 1972. CODEN JMTAA8. ISSN 0020-2932.

Arthur:1972:UIA

[30] D. W. Arthur. The use of interval arithmetic to bound the zeros of real polynomials. Journal of the Institute of Mathematics and its Applications, 10(??):231– 237, 1972. CODEN JMTAA8. ISSN 0020-2932.

Arthur:1973:SFI

[31] D. W. Arthur. The solution of Fredholm integral equations using spline functions. Journal of the Institute of Mathematics and its Applications, 11(??):121–129, 1973. CODEN JMTAA8. ISSN 0020-2932.

Arthurs:1974:EBC

[32] A. M. Arthurs. Error bound for a class of non-linear problems in diffusion and reaction. Journal of the Institute of Mathematics and its Applications, 13(??):63– 66, 1974. CODEN JMTAA8. ISSN 0020-2932.

Arthurs:1971:EBV

[33] A. M. Arthurs and C. W. Coles. Error bounds and variational methods for non-linear differential and integral equations. Journal of the Institute of Mathematics and its Applications, 7(??):324– 330, 1971. CODEN JMTAA8. ISSN 0020-2932.

Arthurs:1974:MMP

[34] A. M. Arthurs and R. I. Reeves. Maximum and minimum principles for a class of partial differential equations. Journal of the Institute of Mathematics and its Applications, 14(??):1–7, 1974. CODEN JMTAA8. ISSN 0020-2932.

Ashton:1971:DGR

[35] W. D. Ashton. Distributions for gaps in road traffic. Journal of the Institute of Mathematics and its Applications, 7(??): 37–46, 1971. CODEN JMTAA8. ISSN 0020-2932.

Askey:1979:PCN

[36] R. Askey. Positivity of the Cotes numbers for some Jacobi abscissas. II. Journal of the Institute of Mathematics and its Applications, 24(1):95–98, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Asvestas:1970:LFSa

[37] J. S. Asvestas and R. E. Kleinman. Low frequency scattering by spheroids and disks. I: Dirichlet problem for a prolate spheroid. Journal of the Institute of Mathematics and its Applications, 6(??): 42–56, 1970. CODEN JMTAA8. ISSN 0020-2932.

Asvestas:1970:LFSb

[38] J. S. Asvestas and R. E. Kleinman. Low frequency scattering by spheroids and disks. II: Neumann problem for a prolate spheroid. Journal of the Institute of Mathematics and its Applications, 6(??): 57–75, 1970. CODEN JMTAA8. ISSN 0020-2932.

Asvestas:1970:LFSc

[39] J. S. Asvestas and R. E. Kleinman. Low frequency scattering by spheroids and disks. III: Oblate spheroids and disks. Journal of the Institute of Mathematics and its Applications, 6(??):157–163, 1970. CODEN JMTAA8. ISSN 0020-2932.

Atanackovic:1978:OCD

[40] T. M. Atanackovic. Optimal control of dynamically controllable motions: an example. Journal of the Institute of Mathematics and its Applications, 21 (??):293–299, 1978. CODEN JMTAA8. ISSN 0020-2932.

Atkin:1976:CTM

[41] R. J. Atkin and R. E. Craine. Continuum theories of mixtures: Applications. Journal of the Institute of Mathematics and its Applications, 17(2):153–207, 1976. CODEN JMTAA8. ISSN 0020-2932.

Atkinson:1977:SSS

[42] C. Atkinson and C. J. Coleman. On some steady-state moving boundary problems in the linear theory of viscoelasticity. Journal of the Institute of Mathematics and its Applications, 20 (??):85–106, 1977. CODEN JMTAA8. ISSN 0020-2932.

Atkinson:1980:ESI

[43] C. Atkinson and H. Javaherian. Evaluation of stress intensity factors associated with bimaterial cracks. Journal of the Institute of Mathematics and its Applications, 26(3):235–258, 1980. CODEN JMTAA8. ISSN 0020-2932.

Atthey:1974:FDS

[44] D. R. Atthey. A finite difference scheme for melting problems. Journal of the Institute of Mathematics and its Applications, 13(??):353–366, 1974. CODEN JMTAA8. ISSN 0020-2932.

Axelsson:1973:NNS

[45] O. Axelsson. Notes on the numerical solution of the biharmonic equation. Journal of the Institute of Mathematics and its Applications, 11(??):213–226, 1973. CODEN JMTAA8. ISSN 0020-2932.

Axelsson:1979:MUS

[46] O. Axelsson and I. Gustafsson. A modified upwind scheme for convective transport equations and the use of a conjugate gradient method for the solution of non-symmetric systems of equations. Journal of the Institute of Mathematics and its Applications, 23(??):321–337, 1979. CODEN JMTAA8. ISSN 0020-2932.

Babolian:1979:AGM

[47] E. Babolian and L. M. Delves. An augmented Galerkin method for first kind Fredholm equations. Journal of the Institute of Mathematics and its Applications, 24(2):157–174, 1979. CODEN JM-TAA8. ISSN 0020-2932.

Backhouse:1975:RAA

[48] R. C. Backhouse and B. A. Carré. Regular Algebra applied to path-finding problems. Journal of the Institute of Mathematics and its Applications, 15(??):161– 186, 1975. CODEN JMTAA8. ISSN 0020-2932.

Bain:1978:UCG

[49] M. Bain. On the uniform convergence of generalized Fourier series. Journal of the Institute of Mathematics and its Applications, 21(??):379–386, 1978. CODEN JMTAA8. ISSN 0020-2932.

Bain:1978:RCM

[50] M. Bain, L. M. Delves, and K. O. Mead. On the rate of convergence of multi-dimensional orthogonal expansions. Journal of the Institute of Mathematics and its Applications, 22(2):225– 234, 1978. CODEN JMTAA8. ISSN 0020-2932.

Ball:1972:TOC

[51] D. J. Ball and J. R. Hewit. The time optimal control of a class of distributed systems. Journal of the Institute of Mathematics and its Applications, 10(??):193– 201, 1972. CODEN JMTAA8. ISSN 0020-2932.

Banks:1977:SFI

[52] S. P. Banks. On the solutions of Fredholm integral equations of the first kind in L^{2}. Journal of the Institute of Mathematics and its Applications, 20(2):143– 150, 1977. CODEN JMTAA8. ISSN 0020-2932.

Bareiss:1972:CSM

[53] Erwin H. Bareiss. Computational solutions of matrix problems over an integral domain. Journal of the Institute of Mathematics and its Applications, 10 (??):68–104, August 1972. CODEN JM-TAA8. ISSN 0020-2932.

Barnett:1971:NFT

[54] S. Barnett. A new formulation of the theorems of Hurwitz, Routh and Sturm. Journal of the Institute of Mathematics and its Applications, 8(??):240–250, 1971. CODEN JMTAA8. ISSN 0020-2932.

Barnett:1968:SAL

[55] S. Barnett and C. Storey. Some applications of the Lyapunov matrix equation. Journal of the Institute of Mathematics and its Applications, 4(??):33–42, 1968. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Barnhill:1975:EAG

[56] R. E. Barnhill and J. R. Whiteman. Error analysis of Galerkin methods for Dirichlet problems containing boundary singularities. Journal of the Institute of Mathematics and its Applications, 15 (??):121–125, 1975. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1974:DVP

[57] M. F. Barnsley and P. D. Robinson. Dual variational principles and Padétype approximants. Journal of the Institute of Mathematics and its Applications, 14(??):229–249, 1974. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1974:PAB

[58] M. F. Barnsley and P. D. Robinson. Padé-approximant bounds and approximate solutions for Kirkwood–Riseman integral equations. Journal of the Institute of Mathematics and its Applications, 14(??):251–265, 1974. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1977:BBN

[59] M. F. Barnsley and P. D. Robinson. Bivariational bounds for non-linear problems. Journal of the Institute of Mathematics and its Applications, 20(4):485– 504, 1977. CODEN JMTAA8. ISSN 0020-2932.

Barratt:1968:SPH

[60] P. J. Barratt. Scattering of plane harmonic waves in an infinite elastic solid by an arbitrary configuration of twodimensional obstacles. Journal of the Institute of Mathematics and its Applications, 4(??):233–249, 1968. CODEN JMTAA8. ISSN 0020-2932.

Barrett:1971:CSR

[61] W. Barrett. On the convergence of sequences of rational approximations to analytic functions of a certain class. Journal of the Institute of Mathematics and its Applications, 7(??):308–323, 1971. CODEN JMTAA8. ISSN 0020-2932.

Barrett:1974:BRA

[62] W. Barrett. On best rational approximations to certain functions defined as integrals. Journal of the Institute of Mathematics and its Applications, 13(??):107– 116, 1974. CODEN JMTAA8. ISSN 0020-2932.

Barrodale:1977:NVG

[63] I. Barrodale and G. F. Stuart. A new variant of Gaussian elimination. Journal of the Institute of Mathematics and its Applications, 19(1):39–47, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Bartholomew-Biggs:1979:MMM

[64] M. C. Bartholomew-Biggs. A matrix modification method for calculating approximate solutions to systems of linear equations. Journal of the Institute of Mathematics and its Applications, 23 (2):131–137, March 1979. CODEN JM-TAA8. ISSN 0020-2932.

Barton:1974:IPE

[65] N. G. Barton. An inequality for the pointwise estimation of solutions of second order elliptic partial differential equations. Journal of the Institute of Mathematics and its Applications, 14 (??):325–333, 1974. CODEN JMTAA8. ISSN 0020-2932.

Baumeister:1980:NSP

[66] J. Baumeister, K.-H. Hoffmann, and P. Jochum. Numerical solution of a parabolic free boundary problem via Newton’s method. Journal of the Institute of Mathematics and its Applications, 25(2):99–109, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Behforooz:1979:ECC

[67] G. H. Behforooz and N. Papamichael. End conditions for cubic spline interpolation. Journal of the Institute of Mathematics and its Applications, 23(3):355– 366, 1979. CODEN JMTAA8. ISSN 0020-2932.

Behforooz:1980:CPC

[68] G. H. Behforooz, N. Papamichael, and A. J. Worsey. A class of piecewise-cubic interpolatory polynomials. Journal of the Institute of Mathematics and its Applications, 25(1):53–65, 1980. CODEN JMTAA8. ISSN 0020-2932.

Bell:1973:MTB

[69] G. E. Bell and J. Crank. A method of treating boundary singularities in timedependent problems. Journal of the Institute of Mathematics and its Applications, 12(??):37–48, 1973. CODEN JM-TAA8. ISSN 0020-2932.

Belleni-Morante:1974:DTV

[70] A. Belleni-Morante and C. Vitocolonna. Discretization of the time variable in evolution equations. Journal of the Institute of Mathematics and its Applications, 14(??):105–112, 1974. CODEN JMTAA8. ISSN 0020-2932.

BenLashiher:1972:FSS

[71] Ali M. Ben Lashiher and C. Storey. Final-stability with some applications. Journal of the Institute of Mathematics and its Applications, 9(??):397–410, 1972. CODEN JMTAA8. ISSN 0020-2932.

Benson:1972:SPB

[72] A. Benson and D. J. Evans. The successive peripheral block over-relaxation method. Journal of the Institute of Mathematics and its Applications, 9(??): 68–79, 1972. CODEN JMTAA8. ISSN 0020-2932.

Biggs:1971:MAM

[73] M. C. Biggs. Minimization algorithms making use of non-quadratic properties of the objective function. Journal of the Institute of Mathematics and its Applications, 8(??):315–327, 1971. CODEN JMTAA8. ISSN 0020-2932.

Biggs:1973:NMA

[74] M. C. Biggs. A note on minimization algorithms which make use of nonquadratic properties of the objective function. Journal of the Institute of Mathematics and its Applications, 12 (??):337–338, 1973. CODEN JMTAA8. ISSN 0020-2932. See note [75].

Biggs:1978:CSC

[75] M. C. Biggs. On the convergence of some constrained minimization algorithms based on recursive quadratic programming. Journal of the Institute of Mathematics and its Applications, 21(1): 67–81, 1978. CODEN JMTAA8. ISSN 0020-2932. See [74].

Bignami:1971:NSC

[76] Anna Bignami and A. De Matteis. A note on sampling from combinations of distributions. Journal of the Institute of Mathematics and its Applications, 8(??): 80–81, 1971. CODEN JMTAA8. ISSN 0020-2932.

Bland:1965:SSS

[77] D. R. Bland. On shock structure in a solid. Journal of the Institute of Mathematics and its Applications, 1(??):56–75, 1965. CODEN JMTAA8. ISSN 0020-2932.

Bland:1966:FE

[78] D. R. Bland. Finite elastodynamics. Journal of the Institute of Mathematics and its Applications, 2(??):327–342, 1966. CODEN JMTAA8. ISSN 0020-2932.

Bleistein:1972:GMS

[79] N. Bleistein and R. A. Handelsman. A generalization of the method of steepest descent. Journal of the Institute of Mathematics and its Applications, 10 (??):211–230, 1972. CODEN JMTAA8. ISSN 0020-2932.

Blowers:1969:RES

[80] R. M. Blowers. On the response of an elastic solid to droplet impact. Journal of the Institute of Mathematics and its Applications, 5(??):167–193, 1969. CO-DEN JMTAA8. ISSN 0020-2932.

Blum:1978:NMS

[81] E. K. Blum and A. F. Chang. A numerical method for the solution of the double eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 22(1):29–42, 1978. CODEN JMTAA8. ISSN 0020-2932. See corrigendum [82].

Blum:1979:CNM

[82] E. K. Blum and A. F. Chang. Corrigendum to: “A numerical method for the solution of the double eigenvalue problem”. Journal of the Institute of Mathematics and its Applications, 23(3):??, 1979. CODEN JMTAA8. ISSN 0020-2932. Loose corrigendum. See [81].

Blum:1971:AWPa

[83] Joseph W. Blum and Donald S. Cohen. Acoustic wave propagation in an underwater sound channel. I. Qualitative theory. Journal of the Institute of Mathematics and its Applications, 8(??):186– 198, 1971. CODEN JMTAA8. ISSN 0020-2932.

Blum:1971:AWPb

[84] Joseph W. Blum and Donald S. Cohen. Acoustic wave propagation in an underwater sound channel. II. Quantitative theory. Journal of the Institute of Mathematics and its Applications, 8(??):199– 220, 1971. CODEN JMTAA8. ISSN 0020-2932.

Bohte:1975:BRE

[85] Z. Bohte. Bounds for rounding errors in the Gaussian elimination for band systems. Journal of the Institute of Mathematics and its Applications, 16(2):133– 142, 1975. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Bojadziev:1974:ASN

[86] G. N. Bojadžiev and R. W. Lardner. Asymptotic solution of a non-linear second order hyperbolic differential equation with large time delay. Journal of the Institute of Mathematics and its Applications, 14(??):203–210, 1974. CODEN JMTAA8. ISSN 0020-2932.

Brindley:1967:TCH

[87] J. Brindley. Thermal convection in horizontal fluid layers. Journal of the Institute of Mathematics and its Applications, 3(??):313–343, 1967. CODEN JM-TAA8. ISSN 0020-2932.

Broadbent:1977:ART

[88] E. G. Broadbent. Acoustic ray theory applied to vortex refraction. Journal of the Institute of Mathematics and its Applications, 19(1):1–27, 1977. CODEN JMTAA8. ISSN 0020-2932.

Brodlie:1974:CME

[89] K. W. Brodlie. Corrections to matrices expressed in non-symmetric product form. Journal of the Institute of Mathematics and its Applications, 14(??):141– 144, 1974. CODEN JMTAA8. ISSN 0020-2932.

Brodlie:1975:NDS

[90] K. W. Brodlie. A new direction set method for unconstrained minimization without evaluating derivatives. Journal of the Institute of Mathematics and its Applications, 15(??):385–396, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Brodlie:1973:ROR

[91] K. W. Brodlie, A. R. Gourlay, and J. Greenstadt. Rank-one and rank-two corrections to positive definite matrices expressed in product form. Journal of the Institute of Mathematics and its Applications, 11(??):73–82, 1973. CODEN JMTAA8. ISSN 0020-2932.

Brodlie:1975:CCJ

[92] K. W. Brodlie and M. J. D. Powell. On the convergence of cyclic Jacobi methods. Journal of the Institute of Mathematics and its Applications, 15(3):279– 287, June 1975. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Brown:1969:FNA

[93] S. N. Brown and K. Stewartson. Flow near the apex of a plane delta wing. Journal of the Institute of Mathematics and its Applications, 5(??):206–216, 1969. CODEN JMTAA8. ISSN 0020-2932.

Brown:1975:SIS

[94] S. N. Brown and P. G. Williams. Selfinduced separation. III. Journal of the Institute of Mathematics and its Applications, 16(??):175–191, 1975. CODEN JMTAA8. ISSN 0020-2932.

Browne:1976:NSE

[95] B. T. Browne and P. J. Lawrenson. Numerical solution of an elliptic boundaryvalue problem in the complex variable. Journal of the Institute of Mathematics and its Applications, 17(3):311–327, 1976. CODEN JMTAA8. ISSN 0020-2932.

Broyden:1966:ADS

[96] C. G. Broyden. An alternative derivation of simplex method. Journal of the Institute of Mathematics and its Applications, 2(??):40–54, 1966. CODEN JM-TAA8. ISSN 0020-2932.

Broyden:1970:CCDa

[97] C. G. Broyden. The convergence of a class of double-rank minimization algorithms. I: General considerations. Journal of the Institute of Mathematics and its Applications, 6(??):76–90, 1970. CO-DEN JMTAA8. ISSN 0020-2932.

Broyden:1970:CCDb

[98] C. G. Broyden. The convergence of a class of double-rank minimization algorithms, II: The new algorithm. Journal of the Institute of Mathematics and its Applications, 6(??):222–231, 1970. CO-DEN JMTAA8. ISSN 0020-2932.

Broyden:1973:SCN

[99] C. G. Broyden. Some condition number bounds for the Gaussian elimination process. Journal of the Institute of Mathematics and its Applications, 12 (??):273–286, 1973. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Broyden:1974:EPN

[100] C. G. Broyden. Error propagation in numerical processes. Journal of the Institute of Mathematics and its Applications, 14(??):131–140, 1974. CODEN JMTAA8. ISSN 0020-2932.

Broyden:1973:LSC

[101] C. G. Broyden, J. E. Dennis, Jr., and Jorge J. More. On the local and superlinear convergence of quasi-Newton methods. Journal of the Institute of Mathematics and its Applications, 12(??):223– 245, 1973. CODEN JMTAA8. ISSN 0020-2932.

Broyden:1975:NMP

[102] C. G. Broyden and J. A. Ford. A new method of polynomial deflation. Journal of the Institute of Mathematics and its Applications, 16(3):271–281, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Brunner:1973:SVI

[103] H. Brunner. The solution of Volterra integral equations of the first kind by piecewise polynomials. Journal of the Institute of Mathematics and its Applications, 12(??):295–302, 1973. CODEN JMTAA8. ISSN 0020-2932.

Brunner:1977:PPC

[104] H. Brunner and M. D. Evans. Piecewise polynomial collocation for Volterra-type integral equations of the second kind. Journal of the Institute of Mathematics and its Applications, 20(4):415–423, 1977. CODEN JMTAA8. ISSN 0020-2932.

Budden:1979:NLE

[105] P. J. Budden and J. Norbury. A nonlinear elliptic eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 24(??):9–33, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Budden:1979:NEE

[106] P. J. Budden and J. Norbury. A nonlinear elliptic eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 24(1):9–33, 1979. CODEN JMTAA8. ISSN 0020-2932.

Bui:1974:SOS

[107] Thuy T. T. Bui and G. Hunter. The significant order of symmetric tridiagonal matrices. Journal of the Institute of Mathematics and its Applications, 14 (??):293–301, 1974. CODEN JMTAA8. ISSN 0020-2932.

Burnett:1972:ENT

[108] D. S. Burnett and W. W. Soroka. An efficient numerical technique for evaluating large quantities of highly oscillatory integrals. Journal of the Institute of Mathematics and its Applications, 10 (??):325–332, 1972. CODEN JMTAA8. ISSN 0020-2932.

Burrows:1980:NAN

[109] B. L. Burrows. A new approach to numerical integration. Journal of the Institute of Mathematics and its Applications, 26(2):151–173, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Burville:1974:HCO

[110] P. J. Burville. HEAPS: A concept in optimization. Journal of the Institute of Mathematics and its Applications, 13 (??):263–278, 1974. CODEN JMTAA8. ISSN 0020-2932.

Butterfield:1976:CAD

[111] K. R. Butterfield. The computation of all the derivatives of a B-spline basis. Journal of the Institute of Mathematics and its Applications, 17(1):15–25, 1976. CODEN JMTAA8. ISSN 0020-2932.

Byatt-Smith:1979:EBE

[112] J. G. B. Byatt-Smith. The equivalence of Bernoulli’s equation and a set of integral relations for periodic waves. Journal of the Institute of Mathematics and its Applications, 23(2):121–130, March 1979. CODEN JMTAA8. ISSN 0020-

Cade:1978:PMS

[113] R. Cade. A perturbation method for solving torus problems in electrostatics. Journal of the Institute of Mathematics and its Applications, 21(??):265–284, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cade:1980:IES

[114] R. Cade. The integral equation solution of two-dimensional electrostatic problems involving circular cylinders. Journal of the Institute of Mathematics and its Applications, 25(3):211–230, 1980. CODEN JMTAA8. ISSN 0020-2932.

Cahit:1978:TST

[115] I. Cahit and R. Cahit. Transformations on spanning trees. Journal of the Institute of Mathematics and its Applications, 21(3):353–358, 1978. CODEN JM-TAA8. ISSN 0020-2932.

Cahlon:1978:NSV

[116] B. Cahlon. The numerical solution of a variational problem in automatic control. Journal of the Institute of Mathematics and its Applications, 22(4):479– 495, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cameron:1968:DIU

[117] I. G. Cameron and G. C. Scorgie. Dynamics of intense underground explosions. Journal of the Institute of Mathematics and its Applications, 4(??):194– 222, 1968. CODEN JMTAA8. ISSN 0020-2932.

Carre:1971:ANR

[118] B. A. Carré. An algebra for network routing problems. Journal of the Institute of Mathematics and its Applications, 7(??):273–294, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Cash:1977:CIA

[119] J. R. Cash. A class of iterative algorithms for the integration of stiff systems of ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 19(3):325–335, 1977. CODEN JMTAA8. ISSN 0020-2932.

Cash:1977:NCA

[120] J. R. Cash. A note on the computational aspects of a class of implicit Runge– Kutta procedures. Journal of the Institute of Mathematics and its Applications, 20(4):425–441, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Cash:1977:CIR

[121] J. R. Cash. On a class of implicit Runge– Kutta procedures. Journal of the Institute of Mathematics and its Applications, 19(4):455–470, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Cash:1978:NCM

[122] J. R. Cash. A note on a class of modified backward differentiation schemes. Journal of the Institute of Mathematics and its Applications, 21(3):301–313, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cash:1979:DIR

[123] J. R. Cash. Diagonally implicit Runge– Kutta formulae with error estimates. Journal of the Institute of Mathematics and its Applications, 24(??):293–301, 1979. CODEN JMTAA8. ISSN 0020-

Cash:1980:DVO

[124] J. R. Cash and C. B. Liem. On the design of a variable order, variable step diagonally implicit Runge–Kutta algorithm. Journal of the Institute of Mathematics and its Applications, 26(1):87– 91, 1980. CODEN JMTAA8. ISSN 0020-2932.

Caudrey:1974:EMS

[125] P. J. Caudrey, J. C. Eilbeck, and J. D. Gibbon. Exact multisoliton solution of the reduced Maxwell–Bloch equations of non-linear optics. Journal of the Institute of Mathematics and its Applications, 14(??):375–386, 1974. CODEN JMTAA8. ISSN 0020-2932.

Cavers:1978:FFT

[126] J. K. Cavers. On the fast Fourier transform inversion of probability generating functions. Journal of the Institute of Mathematics and its Applications, 22(3): 275–282, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cecchi:1977:BSF

[127] M. M. Cecchi and A. Pica. A basis set of functions for curved elements in the finite element methods. Journal of the Institute of Mathematics and its Applications, 20(2):151–162, 1977. CODEN JMTAA8. ISSN 0020-2932.

Chadwick:1966:ISC

[128] P. Chadwick. The initial stages of cooling of a non-homogeneous sphere. Journal of the Institute of Mathematics and its Applications, 2(??):29–39, 1966. CO-DEN JMTAA8. ISSN 0020-2932.

Chadwick:1972:IEE

[129] P. Chadwick and E. W. Haddon. Inflation-extension and eversion of a tube of incompressible isotropic elastic material. Journal of the Institute of Mathematics and its Applications, 10 (??):258–278, 1972. CODEN JMTAA8. ISSN 0020-2932.

Chadwick:1969:TOR

[130] P. Chadwick and A. F. Johnson. Torsional oscillations of a rigid convex inclusion embedded in an elastic solid. Journal of the Institute of Mathematics and its Applications, 5(??):283–307, 1969. CODEN JMTAA8. ISSN 0020-2932.

Chako:1965:AED

[131] Nicholas Chako. Asymptotic expansions of double and multiple integrals occurring in diffraction theory. Journal of the Institute of Mathematics and its Applications, 1(??):372–422, 1965. CODEN JMTAA8. ISSN 0020-2932.

Chakrabarti:1980:NSI

[132] A. Chakrabarti and W. E. Williams.

A note on a singular integral equation. Journal of the Institute of Mathematics and its Applications, 26(3):321–323, 1980. CODEN JMTAA8. ISSN 0020-2932.

Chambers:1967:CTV

[133] L. G. Chambers. A comparison of two variational methods. Journal of the Institute of Mathematics and its Applications, 3(??):166–168, 1967. CODEN JM-TAA8. ISSN 0020-2932.

Chan:1976:VDL

[134] W. L. Chan. Variational dualities in the linear regulator and estimation problems with and without time delay. Journal of the Institute of Mathematics and its Applications, 18(2):237–247, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Chandra:1980:RBM

[135] J. Chandra and P. W. Davis. On the relation between monotone methods and temporal evolution in fully nonlinear boundary value problems. Journal of the Institute of Mathematics and its Applications, 25(3):231–240, 1980. CODEN JMTAA8. ISSN 0020-2932.

Chang:1980:RLR

[136] Y. F. Chang and G. Corliss. Ratio-like and recurrence relation tests for convergence of series. Journal of the Institute of Mathematics and its Applications, 25 (4):349–359, 1980. CODEN JMTAA8. ISSN 0020-2932.

Chapman:1965:WCI

[137] P. B. Chapman. Weak cylindrical implosions. Journal of the Institute of Mathematics and its Applications, 1(??):184– 197, 1965. CODEN JMTAA8. ISSN 0020-2932.

Chawla:1978:FOTa

[138] M. M. Chawla. A fourth-order tridiagonal finite difference method for general non- linear two-point boundary value problems with mixed boundary conditions. Journal of the Institute of Mathematics and its Applications, 21(??):83– 93, 1978. CODEN JMTAA8. ISSN 0020-2932.

Chawla:1978:FOTb

[139] M. M. Chawla. A fourth-order tridiagonal finite difference method for general two-point boundary value problems with non-linear boundary conditions. Journal of the Institute of Mathematics and its Applications, 22(??):89–97, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Chawla:1978:HAT

[140] M. M. Chawla. High-accuracy tridiagonal finite difference approximations for nonlinear two-point boundary value problems. Journal of the Institute of Mathematics and its Applications, 22(2): 203–209, 1978. CODEN JMTAA8. ISSN 0020-2932.

Chawla:1979:SOT

[141] M. M. Chawla. A sixth-order tridiagonal finite difference method for general nonlinear two-point boundary value problems. Journal of the Institute of Mathematics and its Applications, 24(1):35– 42, 1979. CODEN JMTAA8. ISSN 0020-2932.

Chen:1972:PIS

[142] Y. T. Chen. Permutation of irreducible sparse matrices to upper triangular forms. Journal of the Institute of Mathematics and its Applications, 10 (??):15–18, 1972. CODEN JMTAA8. ISSN 0020-2932.

Chester:1975:FOS

[143] W. Chester. The forced oscillations of a simple pendulum. Journal of the Institute of Mathematics and its Applications, 15(??):289–306, 1975. CODEN JMTAA8. ISSN 0020-2932.

Chester:1977:CTD

[144] W. Chester. Continuous transformations and differential equations. Journal of the Institute of Mathematics and its Applications, 19(3):343–376, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Chester:1978:GTR

[145] W. Chester. A general theory of resonance between weakly coupled oscillatory systems. Journal of the Institute of Mathematics and its Applications, 22 (4):383–399, 1978. CODEN JMTAA8. ISSN 0020-2932.

Chester:1980:GTR

[146] W. Chester. A general theory of resonance between weakly coupled oscillatory systems. II. Nonconservative systems. Journal of the Institute of Mathematics and its Applications, 26(2):199– 207, 1980. CODEN JMTAA8. ISSN 0020-2932.

Chew:1980:EBT

[147] K. H. Chew, P. N. Shivakumar, and J. J. Williams. Error bounds for the truncation of infinite linear differential systems. Journal of the Institute of Mathematics and its Applications, 25(1):37– 51, 1980. CODEN JMTAA8. ISSN 0020-2932.

Chien:1972:MAS

[148] Henry H. Y. Chien. A multiphase algorithm for single variable equation solving. Journal of the Institute of Mathematics and its Applications, 9(??):290– 298, 1972. CODEN JMTAA8. ISSN 0020-2932.

Christiansen:1978:IPI

[149] S. Christiansen and P. Hougaard. An investigation of a pair of integral equations for the biharmonic problem. Journal of the Institute of Mathematics and its Applications, 22(1):15–27, 1978. CODEN JMTAA8. ISSN 0020-2932.

Christiansen:1975:IEU

[150] Søren Christiansen. Integral equations without a unique solution can be made useful for solving some plane harmonic problems. Journal of the Institute of Mathematics and its Applications, 16(2):

143–159, 1975. CODEN JMTAA8. ISSN 0020-2932.

Churchhouse:1969:CFA

[151] R. F. Churchhouse and S. T. E. Muir.

Continued fractions, algebraic numbers and modular invariants. Journal of the Institute of Mathematics and its Applications, 5(??):318–328, 1969. CODEN JMTAA8. ISSN 0020-2932.

Cicenia:1976:NIF

[152] R. A. Cicenia. Numerical integration formulae involving derivatives. Journal of the Institute of Mathematics and its Applications, 18(1):79–85, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Cicenia:1979:RNI

[153] R. A. Cicenia. Remainders in numerical integration formulae involving derivatives. Journal of the Institute of Mathematics and its Applications, 24(3):347– 352, 1979. CODEN JMTAA8. ISSN 0020-2932.

Cioranescu:1979:EPT

[154] Doina Cioranescu, Jeannine Saint Jean Paulin, and Hélène Lanchon. Elasticplastic torsion of heterogeneous cylindrical bars. Journal of the Institute of Mathematics and its Applications, 24(4): 353–378, 1979. CODEN JMTAA8. ISSN 0020-2932.

Clarke:1977:BCL

[155] B. M. N. Clarke. Boundary controllability of linear symmetric hyperbolic systems. Journal of the Institute of Mathematics and its Applications, 20(3):283– 298, 1977. CODEN JMTAA8. ISSN 0020-2932.

Clements:1971:MHC

[156] D. L. Clements. The motion of a heavy cylinder over the surface of anisotropic elastic solid. Journal of the Institute of Mathematics and its Applications, 7(??): 198–206, 1971. CODEN JMTAA8. ISSN 0020-2932.

Clements:1979:MNS

[157] D. L. Clements and G. W. King. A method for the numerical solution of problems governed by elliptic systems in the cut plane. Journal of the Institute of Mathematics and its Applications, 24(1): 81–93, 1979. CODEN JMTAA8. ISSN 0020-2932.

Clements:1978:MNS

[158] D. L. Clements and F. J. Rizzo. A method for the numerical solution of boundary value problems governed by second-order elliptic systems. Journal of the Institute of Mathematics and its Applications, 22(2):197–202, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Clements:1974:ABT

[159] D. L. Clements and C. Rogers. On the application of a Bäcklund transformation to linear isotropic elasticity. Journal of the Institute of Mathematics and its Applications, 14(??):23–30, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Clenshaw:1965:CSF

[160] C. W. Clenshaw and J. G. Hayes. Curve and surface fitting. Journal of the Institute of Mathematics and its Applications, 1(??):164–183, 1965. CODEN JM-TAA8. ISSN 0020-2932.

Clenshaw:1978:CSA

[161] C. W. Clenshaw and B. Negus. The cubic X-spline and its application to interpolation. Journal of the Institute of Mathematics and its Applications, 22(1): 109–119, August 1978. CODEN JM-TAA8. ISSN 0020-2932.

Clint:1971:SIM

[162] Maurice Clint and Alan Jennings. A simultaneous iteration method for the unsymmetric eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 8(??):111–121, 1971. CO-DEN JMTAA8. ISSN 0020-2932.

Cohen:1977:DTA

[163] A. M. Cohen. A divergent technique for the approximate solution of some simultaneous linear equations by iterative methods. Journal of the Institute of Mathematics and its Applications, 19(2): 187–193, 1977. CODEN JMTAA8. ISSN 0020-2932.

Cohen:1974:NNS

[164] A. M. Cohen and D. E. Jones. A note on the numerical solution of some singular second order differential equations. Journal of the Institute of Mathematics and its Applications, 13(??):379–384, 1974. CODEN JMTAA8. ISSN 0020-2932.

Cohen:1977:TSE

[165] A. M. Cohen and D. E. Jones. A technique for the solution of eigenvalue problems. Journal of the Institute of Mathematics and its Applications, 20(1):1–7, 1977. CODEN JMTAA8. ISSN 0020-2932.

Cohen:1979:GCO

[166] E. Cohen. A general class of optimal polynomials related to electric filters. Journal of the Institute of Mathematics and its Applications, 24(2):197–208, 1979. CODEN JMTAA8. ISSN 0020-2932.

Cohen:1969:EDD

[167] J. K. Cohen and D. L. Hector. Edge ’diffraction’ for the diffusion equation. Journal of the Institute of Mathematics and its Applications, 5(??):1–18, 1969. CODEN JMTAA8. ISSN 0020-2932.

Cohen:1967:RMA

[168] Jack K. Cohen and Robert M. Lewis. A ray method for the asymptotic solution of the diffusion equation. Journal of the Institute of Mathematics and its Applications, 3(??):266–290, 1967. CODEN JMTAA8. ISSN 0020-2932.

Cole:1977:CVP

[169] R. J. Cole. Complementary variational principles for Knudsen flow rates. Journal of the Institute of Mathematics and its Applications, 20(??):107–115, 1977. CODEN JMTAA8. ISSN 0020-2932.

Coleman:1976:LTM

[170] J. P. Coleman. The Lanczos taumethod. Journal of the Institute of Mathematics and its Applications, 17(1): 85–97, 1976. CODEN JMTAA8. ISSN 0020-2932.

Collins:1969:SSD

[171] R. Collins and A. Wragg. The solution of a statistical decision problem with moving boundaries by the image method. Journal of the Institute of Mathematics and its Applications, 5(??):308–317, 1969. CODEN JMTAA8. ISSN 0020-2932.

Common:1974:SPC

[172] A. K. Common and P. R. Graves-Morris. Some properties of Chisholm approximants. Journal of the Institute of Mathematics and its Applications, 13(??): 229–232, 1974. CODEN JMTAA8. ISSN 0020-2932.

Constantine:1978:SPA

[173] A. G. Constantine and J. C. Gower.

Some properties and applications of simple orthogonal matrices. Journal of the Institute of Mathematics and its Applications, 21(4):445–454, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cooke:1978:NPS

[174] J. C. Cooke. Note on the principle of stationary phase. Journal of the Institute of Mathematics and its Applications, 22(3): 297–303, 1978. CODEN JMTAA8. ISSN 0020-2932.

Cooley:1968:SRS

[175] M. D. A. Cooley and M. E. O’Neil. On the slow rotation of a sphere about a diameter parallel to a nearby plane wall. Journal of the Institute of Mathematics and its Applications, 4(??):163–173, 1968. CODEN JMTAA8. ISSN 0020-2932.

Copson:1966:BVP

[176] E. T. Copson and P. Keast. On a boundary-value problem for the equation of heat. Journal of the Institute of Mathematics and its Applications, 2(??): 358–363, 1966. CODEN JMTAA8. ISSN 0020-2932.

Cotsaftis:1969:SM

[177] M. Cotsaftis. On stability of motion.

Journal of the Institute of Mathematics and its Applications, 5(??):19–54, 1969. CODEN JMTAA8. ISSN 0020-2932.

Cox:1971:CFP

[178] M. G. Cox. Curve fitting with piecewise polynomials. Journal of the Institute of Mathematics and its Applications, 8(??): 36–52, 1971. CODEN JMTAA8. ISSN 0020-2932.

Cox:1972:NEB

[179] M. G. Cox. The numerical evaluation of B-splines. Journal of the Institute of Mathematics and its Applications, 10 (??):134–149, 1972. CODEN JMTAA8. ISSN 0020-2932.

Cox:1975:ASI

[180] M. G. Cox. An algorithm for spline interpolation. Journal of the Institute of Mathematics and its Applications, 15 (??):95–108, 1975. CODEN JMTAA8. ISSN 0020-2932.

Cox:1978:NES

[181] M. G. Cox. The numerical evaluation of a spline from its B-spline representation. Journal of the Institute of Mathematics and its Applications, 21(??):135– 143, 1978. CODEN JMTAA8. ISSN 0020-2932.

Craggs:1978:PSM

[182] J. W. Craggs and P. W. Duck. A power series method for treating mixed boundary value problems. Journal of the Institute of Mathematics and its Applications, 21(??):1–12, 1978. CODEN JM-TAA8. ISSN 0020-2932.

Crank:1977:TBS

[183] J. Crank and R. M. Furzeland. The treatment of boundary singularities in axially symmetric problems containing discs. Journal of the Institute of Mathematics and its Applications, 20(3):355– 370, 1977. CODEN JMTAA8. ISSN 0020-2932.

Crank:1980:NSF

[184] J. Crank and T. Ozis. Numerical solution of a free boundary problem by interchanging dependent and independent variables. Journal of the Institute of Mathematics and its Applications, 26(1): 77–85, 1980. CODEN JMTAA8. ISSN 0020-2932.

Crank:1972:MSM

[185] John Crank and R. S. Gupta. A method for solving moving boundary problems in heat flow using cubic splines or polynomials. Journal of the Institute of Mathematics and its Applications, 10 (??):296–304, 1972. CODEN JMTAA8. ISSN 0020-2932.

Crank:1972:MBP

[186] John Crank and Radhey S. Gupta. A moving boundary problem arising from the diffusion of oxygen in absorbing tissue. Journal of the Institute of Mathematics and its Applications, 10(??):19– 33, 1972. CODEN JMTAA8. ISSN 0020-2932.

Crapper:1965:SMF

[187] G. D. Crapper. Steady magnetogasdynamic flow past a source. Journal of the Institute of Mathematics and its Applications, 1(??):241–257, 1965. CODEN JMTAA8. ISSN 0020-2932.

Crowley:1979:WSM

[188] A. B. Crowley. On the weak solution of moving boundary problems. Journal of the Institute of Mathematics and its Applications, 24(1):43–57, 1979. CODEN JMTAA8. ISSN 0020-2932.

Crowley:1977:SPN

[189] A. B. Crowley and J. R. Ockendon. A Stefan problem with a non-monotone boundary. Journal of the Institute of Mathematics and its Applications, 20(3): 269–281, 1977. CODEN JMTAA8. ISSN 0020-2932.

Cryer:1980:PCF

[190] C. W. Cryer. A proof of the convexity of the free boundary for porous flow through a rectangular dam using the maximum principle. Journal of the Institute of Mathematics and its Applications, 25(2):111–120, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Cullen:1973:SFE

[191] M. J. P. Cullen. A simple finite element method for meteorological problems. Journal of the Institute of Mathematics and its Applications, 11(??):15– 31, 1973. CODEN JMTAA8. ISSN 0020-2932.

Cullen:1974:FEM

[192] M. J. P. Cullen. A finite element method for a non-linear initial value problem. Journal of the Institute of Mathematics and its Applications, 13(??):233–247, 1974. CODEN JMTAA8. ISSN 0020-2932.

Cumberbatch:1966:GSS

[193] E. Cumberbatch and E. Varley. Generalized self-similar flows. Journal of the Institute of Mathematics and its Applications, 2(??):1–11, 1966. CODEN JM-TAA8. ISSN 0020-2932.

Cupini:1977:MSD

[194] E. Cupini. A mixed stochasticdeterministic method for systems of linear equations. Journal of the Institute of Mathematics and its Applications, 19 (4):447–454, 1977. CODEN JMTAA8. ISSN 0020-2932.

Curle:1979:ACS

[195] N. Curle. Analysis of certain slowly converging series. Journal of the Institute of Mathematics and its Applications, 23 (3):265–275, 1979. CODEN JMTAA8. ISSN 0020-2932.

Curle:1980:UBL

[196] S. N. Curle. An unsteady boundary layer problem—a critical test of extended series expansions. Journal of the Institute of Mathematics and its Applications, 25 (3):199–209, 1980. CODEN JMTAA8. ISSN 0020-2932.

Currie:1969:LTW

[197] P. Currie and M. Hayes. Longitudinal and transverse waves in finite elastic strain. Hadamard and Green materials. Journal of the Institute of Mathematics and its Applications, 5(??):140– 161, 1969. CODEN JMTAA8. ISSN 0020-2932.

Curtis:1975:HOE

[198] A. R. Curtis. High-order explicit Runge– Kutta formulae, their uses, and limitations. Journal of the Institute of Mathematics and its Applications, 16(1):35– 55, 1975. CODEN JMTAA8. ISSN 0020-2932.

Curtis:1979:IIR

[199] A. R. Curtis. The implementation of implicit Runge–Kutta formulae for the solution of initial-value problems. Journal of the Institute of Mathematics and its Applications, 23(3):339–353, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Curtis:1974:ESJ

[200] A. R. Curtis, M. J. D. Powel, and J. K. Reid. On the estimation of sparse Jacobian matrices. Journal of the Institute of Mathematics and its Applications, 13 (??):117–119, 1974. CODEN JMTAA8. ISSN 0020-2932.

Curtis:1971:SLS

[201] A. R. Curtis and J. K. Reid. The solution of large sparse unsymmetric systems of linear equations. Journal of the Institute of Mathematics and its Applications, 8(??):344–353, 1971. CODEN JMTAA8. ISSN 0020-2932.

Curtis:1972:ASM

[202] A. R. Curtis and J. K. Reid. On the automatic scaling of matrices for Gaussian elimination. Journal of the Institute of Mathematics and its Applications, 10 (??):118–124, 1972. CODEN JMTAA8. ISSN 0020-2932.

Curtis:1974:CSL

[203] A. R. Curtis and J. K. Reid. The choice of step lengths when using differences to approximate Jacobian matrices. Journal of the Institute of Mathematics and its Applications, 13(??):121–126, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Daley:1965:SR

[204] D. J. Daley and D. G. Kendall. Stochastic rumours. Journal of the Institute of Mathematics and its Applications, 1(??): 42–55, 1965. CODEN JMTAA8. ISSN 0020-2932.

Daniel:1975:ECT

[205] J. W. Daniel and B. K. Swartz. Extrapolated collocation for two-point boundary-value problems using cubic splines. Journal of the Institute of Mathematics and its Applications, 16(2):161– 174, 1975. CODEN JMTAA8. ISSN 0020-2932.

Daniels:1976:NAI

[206] P. G. Daniels. A numerical and asymptotic investigation of boundary-layer wake evolution. Journal of the Institute of Mathematics and its Applications, 17 (3):367–386, 1976. CODEN JMTAA8. ISSN 0020-2932.

Daniels:1980:ESO

[207] P. G. Daniels. The effect of sidewalls on overstable convection in a rotating fluid. Journal of the Institute of Mathematics and its Applications, 26(2):103–118, 1980. CODEN JMTAA8. ISSN 0020-2932.

Datko:1968:ASC

[208] R. Datko. The asymptotic stability of a class of second order differential equations. Journal of the Institute of Mathematics and its Applications, 4(??):276– 281, 1968. CODEN JMTAA8. ISSN 0020-2932.

Datta:1977:HEH

[209] N. Datta and R. N. Jana. Hall effects on hydromagnetic flow and heat transfer in a rotating channel. Journal of the Institute of Mathematics and its Applications, 19(??):217–229, 1977. CODEN JMTAA8. ISSN 0020-2932.

Davies:1976:IPS

[210] M. Davies and B. Dawson. The incrementation parameter in square root iteration. Journal of the Institute of Mathematics and its Applications, 17(2):219– 223, 1976. CODEN JMTAA8. ISSN 0020-2932.

Davies:1972:TOC

[211] M. J. Davies. Time optimal control and the Duffing oscillator. Journal of the Institute of Mathematics and its Applications, 9(??):357–369, 1972. CODEN JM-TAA8. ISSN 0020-2932.

Davis:1970:WPI

[212] A. M. J. Davis. Waves in the presence of an infinite dock with gap. Journal of the Institute of Mathematics and its Applications, 6(??):141–156, 1970. CODEN JMTAA8. ISSN 0020-2932.

Davis:1975:SSW

[213] A. M. J. Davis. Short surface waves due to a heaving sphere in a hemispherical lake. Journal of the Institute of Mathematics and its Applications, 16(2):221– 237, 1975. CODEN JMTAA8. ISSN 0020-2932.

Day:1973:GQM

[214] J. T. Day. A Gaussian quadrature method for the numerical solution of the Goursat problem. Journal of the Institute of Mathematics and its Applications, 12(??):155–158, 1973. CODEN JMTAA8. ISSN 0020-2932.

DellaTorre:1973:CPS

[215] E. Della Torre and W. Kinsner. Convergence properties of the successive extrapolated relaxation (S.E.R.) method. Journal of the Institute of Mathematics and its Applications, 12(??):175–185, 1973. CODEN JMTAA8. ISSN 0020-2932.

Delves:1977:FMS

[216] L. M. Delves. A fast method for the solution of Fredholm integral equations. Journal of the Institute of Mathematics and its Applications, 20(2):173–182, 1977. CODEN JMTAA8. ISSN 0020-2932.

Delves:1977:NSS

[217] L. M. Delves. The numerical solution of sets of linear equations arising from Ritz–Galerkin methods. Journal of the Institute of Mathematics and its Applications, 20(2):163–171, 1977. CODEN JMTAA8. ISSN 0020-2932.

Delves:1979:FGA

[218] L. M. Delves, L. F. Abd-Elal, and J. A.

Hendry. A fast Galerkin algorithm for singular integral equations. Journal of the Institute of Mathematics and its Applications, 23(2):139–166, March 1979. CODEN JMTAA8. ISSN 0020-2932.

Delves:1979:FDM

[219] L. M. Delves and I. Barrodale. A fast direct method for the least squares solution of slightly overdetermined sets of linear equations. Journal of the Institute of Mathematics and its Applications, 24 (2):149–156, 1979. CODEN JMTAA8. ISSN 0020-2932.

Delves:1979:IMP

[220] L. M. Delves and C. A. Hall. An implicit matching principle for global element calculations. Journal of the Institute of Mathematics and its Applications, 23(2):223–234, March 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Delves:1980:FIG

[221] L. M. Delves and C. Phillips. A fast implementation of the global element method. Journal of the Institute of Mathematics and its Applications, 25(2): 177–197, 1980. CODEN JMTAA8. ISSN 0020-2932.

Dennis:1972:MVF

[222] S. C. R. Dennis. The motion of a viscous fluid past an impulsively started semi-infinite flat plate. Journal of the Institute of Mathematics and its Applications, 10(??):105–117, 1972. CODEN JMTAA8. ISSN 0020-2932.

Dennis:1979:ARP

[223] S. C. R. Dennis and J. D. Hudson. Accurate representations of partial differential equations by finite-difference schemes. Journal of the Institute of Mathematics and its Applications, 23(1): 43–51, 1979. CODEN JMTAA8. ISSN 0020-2932.

Dennis:1980:FAR

[224] S. C. R. Dennis and J. D. Hudson. Further accurate representations of partial differential equations by finite-difference methods. Journal of the Institute of Mathematics and its Applications, 26(4): 369–379, 1980. CODEN JMTAA8. ISSN 0020-2932.

Dew:1978:NNS

[225] P. M. Dew. A note on the numerical solution of quasi-linear parabolic partial differential equations with error estimates. Journal of the Institute of Mathematics and its Applications, 21(??):401– 417, 1978. CODEN JMTAA8. ISSN 0020-2932.

Dew:1972:IMS

[226] P. M. Dew and R. E. Scraton. An improved method for the solution of the heat equation in Chebyshev series. Journal of the Institute of Mathematics and its Applications, 9(3):299–309, 1972. CODEN JMTAA8. ISSN 0020-2932.

Dew:1973:SHE

[227] P. M. Dew and R. E. Scraton. The solution of the heat equation in two space variables using Chebyshev series. Journal of the Institute of Mathematics and its Applications, 11(2):231–240, 1973. CODEN JMTAA8. ISSN 0020-2932.

Dew:1975:CMN

[228] P. M. Dew and R. E. Scraton. Chebyshev methods for the numerical solution of parabolic partial differential equations in two and three space variables. Journal of the Institute of Mathematics and its Applications, 16(1):121–131, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Dhaliwal:1971:SST

[229] Ranjit S. Dhaliwal. The steadystate thermoelastic mixed boundaryvalue problem for the elastic layer. Journal of the Institute of Mathematics and its Applications, 7(??):295–302, 1971. CODEN JMTAA8. ISSN 0020-2932.

Dixon:1973:CDL

[230] L. C. W. Dixon. Conjugate directions without linear searches. Journal of the Institute of Mathematics and its Applications, 11(??):317–328, 1973. CODEN JMTAA8. ISSN 0020-2932.

Dixon:1975:CGA

[231] L. C. W. Dixon. Conjugate gradient algorithms: quadratic termination without linear searches. Journal of the Institute of Mathematics and its Applications, 15(??):9–18, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Djukic:1978:MDS

[232] Dj.S. Djukic. On the motion of a dynamic system with minimal dissipation.

Journal of the Institute of Mathematics and its Applications, 21(??):41–46, 1978. CODEN JMTAA8. ISSN 0020-2932.

Dollimore:1973:SAU

[233] Jean Dollimore. Some algorithms for use with the fast Fourier transform. Journal of the Institute of Mathematics and its Applications, 12(??):115–117, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Dolph:1965:NVP

[234] Charles L. Dolph and John M. Richardson. A note on variational principles for unsteady compressible flows in one dimension. Journal of the Institute of Mathematics and its Applications, 1(??): 113–117, 1965. CODEN JMTAA8. ISSN 0020-2932.

Donnelly:1974:BEN

[235] J. D. P. Donnelly. Bounds for the eigenvalues of non-selfadjoint differential operators. Journal of the Institute of Mathematics and its Applications, 13 (??):249–261, 1974. CODEN JMTAA8. ISSN 0020-2932.

Donnelly:1979:MNT

[236] J. D. P. Donnelly. A model for nonequilibrium thermodynamic processes involving phase changes. Journal of the Institute of Mathematics and its Applications, 24(4):425–438, 1979. CODEN JMTAA8. ISSN 0020-2932.

Doran:1969:WSC

[237] H. E. Doran. The wedge with a symmetric crack at the vertex in plane elastostatics. Journal of the Institute of Mathematics and its Applications, 5(4):363– 372, 1969. CODEN JMTAA8. ISSN 0020-2932.

Doran:1969:HPE

[238] H. E. Doran and V. T. Buchwald. The half-plane with an edge crack in plane elastostatics. Journal of the Institute of Mathematics and its Applications, 5(??): 91–112, 1969. CODEN JMTAA8. ISSN 0020-2932.

Dore:1974:EES

[239] B. D. Dore and D. Porter. The edge effect of a surface film on gravity waves. Journal of the Institute of Mathematics and its Applications, 14(??):119–129, 1974. CODEN JMTAA8. ISSN 0020-2932.

Downs:1972:IRM

[240] T. Downs. Inversion of rational matrices by the reinforcement method. Journal of the Institute of Mathematics and its Applications, 10(??):34–41, 1972. CODEN JMTAA8. ISSN 0020-2932.

Drew:1977:BPF

[241] D. M. Drew and J. A. Murphy. Branch points, M fractions and rational approximants generated by linear equations. Journal of the Institute of Mathematics and its Applications, 19(2):169–185, 1977. CODEN JMTAA8. ISSN 0020-2932.

Duff:1977:PBT

[242] Iain S. Duff. On permutations to block triangular form. Journal of the Institute of Mathematics and its Applications, 19 (3):339–342, 1977. CODEN JMTAA8. ISSN 0020-2932.

Duff:1974:CSO

[243] Iain S. Duff and John K. Reid. A comparison of sparsity orderings for obtaining a pivotal sequence in Gaussian elimination. Journal of the Institute of Mathematics and its Applications, 14 (??):281–291, 1974. CODEN JMTAA8. ISSN 0020-2932.

Duff:1975:RSM

[244] Iain S. Duff and John K. Reid. On the reduction of sparse matrices to condensed forms by similarity transformations. Journal of the Institute of Mathematics and its Applications, 15(??):217– 224, 1975. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Duff:1976:CSM

[245] Iain S. Duff and John K. Reid. A comparison of some methods for the solution of sparse overdetermined systems of linear equations. Journal of the Institute of Mathematics and its Applications, 17 (3):267–280, 1976. CODEN JMTAA8. ISSN 0020-2932.

Duff:1979:DSS

[246] Iain S. Duff, John K. Reid, Neils Munskgaard, and Hans B. Nielsen. Direct solution of sets of linear equations whose matrix is sparse, symmetric and indefinite. Journal of the Institute of Mathematics and its Applications, 23(2):235– 250, March 1979. CODEN JMTAA8. ISSN 0020-2932.

Duggan:1980:DEP

[247] R. C. Duggan. On dual extremum principles and the hypercircle for a class of partial differential equations. Journal of the Institute of Mathematics and its Applications, 26(3):259–267, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Dunaway:1972:MGJ

[248] Donna K. Dunaway and B. L. Turlington. Modifications to the Garside, Jarratt P. Mack method for solving illconditioned polynomial equations. Journal of the Institute of Mathematics and its Applications, 10(??):358–368, 1972. CODEN JMTAA8. ISSN 0020-2932.

Dunham:1971:CA

[249] Charles B. Dunham. Chebyshev approximation by A + B log(1 + CX). Journal of the Institute of Mathematics and its Applications, 8(??):371–373, 1971. CO-DEN JMTAA8. ISSN 0020-2932.

Dunham:1972:AEP

[250] Charles B. Dunham. Approximation by exponential-polynomial products on finite point sets. Journal of the Institute of Mathematics and its Applications, 10 (??):125–127, 1972. CODEN JMTAA8. ISSN 0020-2932.

Dunham:1972:CAI

[251] Charles B. Dunham. Chebyshev approximation by A+B^{∗} log(1+CX). II. Journal of the Institute of Mathematics and its Applications, 10(??):369–372, 1972. CODEN JMTAA8. ISSN 0020-2932.

Dunham:1973:CAE

[252] Charles B. Dunham. Chebyshev approximation by exponential ratios. Journal of the Institute of Mathematics and its Applications, 11(??):247–249, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Durbin:1979:AEL

[253] P. Durbin. Asymptotic expansion of Laplace transforms about the origin using generalized functions. Journal of the Institute of Mathematics and its Applications, 23(2):181–192, March 1979. CODEN JMTAA8. ISSN 0020-2932.

Dyakonov:1971:SSE

[254] E. G. D’yakonov. On the solution of some elliptic difference equations. Journal of the Institute of Mathematics and its Applications, 7(??):1–20, 1971. CO-DEN JMTAA8. ISSN 0020-2932.

Eason:1966:DPE

[255] G. Eason. The displacement produced in an elastic half-space by a suddenly applied surface force. Journal of the Institute of Mathematics and its Applications, 2(??):299–326, 1966. CODEN JM-TAA8. ISSN 0020-2932.

Easton:1973:FPM

[256] M. C. Easton. A fixed point method for Tchebycheff solution of inconsistent linear equations. Journal of the Institute of Mathematics and its Applications, 12 (??):137–154, 1973. CODEN JMTAA8. ISSN 0020-2932.

Eberlein:1971:DCS

[257] P. J. Eberlein. On the diagonalization of complex symmetric matrices. Journal of the Institute of Mathematics and its Applications, 7(??):377–383, 1971. CO-DEN JMTAA8. ISSN 0020-2932. Cited in [?].

Edwards:1979:ULM

[258] J. T. Edwards, D. C. Licciardello, and D. J. Thouless. Use of the Lanczos method for finding complete sets of eigenvalues of large sparse symmetric matrices. Journal of the Institute of Mathematics and its Applications, 23(3): 277–283, 1979. CODEN JMTAA8. ISSN 0020-2932.

Ehle:1975:GRK

[259] B. L. Ehle and J. D. Lawson. Generalized Runge–Kutta processes for stiff initial-value problems. Journal of the Institute of Mathematics and its Applications, 16(1):11–21, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Eilon:1976:COS

[260] S. Eilon and I. G. Chowdhury. Conditions for optimal scheduling for the three-machine problem. Journal of the Institute of Mathematics and its Applications, 17(??):1–4, 1976. CODEN JM-TAA8. ISSN 0020-2932.

El-Gendi:1975:CSP

[261] S. E. El-Gendi. On Chebyshev solution of parabolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 16(3):283– 289, 1975. CODEN JMTAA8. ISSN 0020-2932.

ElTom:1971:ASF

[262] M. E. A. El Tom. Application of spline functions to Volterra integral equations. Journal of the Institute of Mathematics and its Applications, 8(??):354–357, 1971. CODEN JMTAA8. ISSN 0020-2932.

ElTom:1974:SFA

[263] M. E. A. El Tom. Spline function approximations to the solution of singular Volterra integral equations of the second kind. Journal of the Institute of Mathematics and its Applications, 14(??):303– 309, 1974. CODEN JMTAA8. ISSN 0020-2932.

ElTom:1976:ASF

[264] M. E. A. El Tom. Application of spline functions to systems of Volterra integral equations of the first and second kinds. Journal of the Institute of Mathematics and its Applications, 17(3):295–310, 1976. CODEN JMTAA8. ISSN 0020-2932.

Elliott:1980:VIF

[265] C. M. Elliott. On a variational inequality formulation of an electrochemical machining moving boundary problem and its approximation by the finite element method. Journal of the Institute of Mathematics and its Applications, 25 (2):121–131, 1980. CODEN JMTAA8. ISSN 0020-2932.

Ellman:1974:DPA

[266] D. Ellman and W. Townsend. A dynamic programming approach to the three-machine scheduling problem. Journal of the Institute of Mathematics and its Applications, 13(??):161–167, 1974. CODEN JMTAA8. ISSN 0020-2932.

Emamzadeh:1977:NISa

[267] A. M. Emamzadeh. Numerical investigations into the Stokes phenomenon. I. Journal of the Institute of Mathematics and its Applications, 19(1):77–86, 1977. CODEN JMTAA8. ISSN 0020-2932.

Emamzadeh:1977:NISb

[268] A. M. Emamzadeh. Numerical investigations into the Stokes phenomenon. II. Journal of the Institute of Mathematics and its Applications, 19(2):149–157, 1977. CODEN JMTAA8. ISSN 0020-2932.

Emenalo:1972:IWS

[269] S. I. Emenalo. Interaction of weak shock waves with magnetohydrodynamic boundary layers. Journal of the Institute of Mathematics and its Applications, 9(??):234–249, 1972. CODEN JMTAA8. ISSN 0020-2932.

England:1972:SBV

[270] A. H. England. The stress boundary value problem for plane strain deformations of an ideal fibre-reinforced material. Journal of the Institute of Mathematics and its Applications, 9(??):310– 322, 1972. CODEN JMTAA8. ISSN 0020-2932.

Erdelyi:1968:TVE

[271] A. Erdélyi. Two-variable expansions for singular perturbations. Journal of the Institute of Mathematics and its Applications, 4(??):113–119, 1968. CODEN JMTAA8. ISSN 0020-2932.

Estrada:1980:ADD

[272] R. Estrada and R. P. Kanwal. Applications of distributional derivatives to wave propagation. Journal of the Institute of Mathematics and its Applications, 26(1):39–63, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Evans:1966:SIM

[273] D. J. Evans. Simplified implicit methods for the finite difference solution of parabolic and hyperbolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 2 (??):76–96, 1966. CODEN JMTAA8. ISSN 0020-2932.

Evans:1968:UPC

[274] D. J. Evans. The use of pre-conditioning in iterative methods for solving linear equations with symmetric positive definite matrices. Journal of the Institute of Mathematics and its Applications, 4 (??):295–314, 1968. CODEN JMTAA8. ISSN 0020-2932.

Evans:1971:NSF

[275] D. J. Evans. The numerical solution of the fourth boundary value problem for parabolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 7(??):61–75, 1971. CODEN JMTAA8. ISSN 0020-2932.

Evans:1977:NSE

[276] D. J. Evans. Numerical solution of exterior problems by the peripheral block over- relaxation method. Journal of the Institute of Mathematics and its Applications, 19(4):399–405, 1977. CODEN JMTAA8. ISSN 0020-2932.

Evans:1975:AND

[277] D. J. Evans and S. O. Fatunla. Accurate numerical determination of the intersection point of the solution of a differential equation with a given algebraic relation. Journal of the Institute of Mathematics and its Applications, 16(3):355– 359, 1975. CODEN JMTAA8. ISSN 0020-2932.

Evans:1976:NWP

[278] D. V. Evans. A note on the waves produced by the small oscillations of a partially vertical plate. Journal of the Institute of Mathematics and its Applications, 17(??):135–140, 1976. CODEN JMTAA8. ISSN 0020-2932.

Evans:1978:OWC

[279] D. V. Evans. The oscillating water column wave-energy device. Journal of the Institute of Mathematics and its Applications, 22(??):423–433, 1978. CODEN JMTAA8. ISSN 0020-2932.

Evans:1972:CAS

[280] D. V. Evans and C. A. N. Morris. Complementary approximations to the solution of a problem in water waves. Journal of the Institute of Mathematics and its Applications, 10(??):1–9, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Evans:1972:EFV

[281] D. V. Evans and C. A. N. Morris.

The effect of a fixed vertical barrier on obliquely incident surface waves in deep water. Journal of the Institute of Mathematics and its Applications, 9(??):198– 204, 1972. CODEN JMTAA8. ISSN 0020-2932.

Evans:1977:STD

[282] N. T. S. Evans and A. R. Gourlay.

The solution of a two-dimensional timedependent diffusion problem concerned with oxygen metabolism in tissues. Journal of the Institute of Mathematics and its Applications, 19(2):239–251, 1977. CODEN JMTAA8. ISSN 0020-2932.

Exton:1980:BAH

[283] Harold Exton. A basic analogue of Hermite’s equation. Journal of the Institute of Mathematics and its Applications, 26 (3):315–320, 1980. CODEN JMTAA8. ISSN 0020-2932.

Fairweather:1979:ASD

[284] G. Fairweather. On the approximate solution of a diffusion problem by Galerkin methods. Journal of the Institute of Mathematics and its Applications, 24(2): 121–137, 1979. CODEN JMTAA8. ISSN 0020-2932.

Fairweather:1965:HAA

[285] G. Fairweather and A. R. Mitchell.

A high accuracy alternating direction method for the wave equation. Journal of the Institute of Mathematics and its Applications, 1(??):309–316, 1965. CO-DEN JMTAA8. ISSN 0020-2932.

Fasano:1979:LFP

[286] A. Fasano and M. Primicerio. Liquid flow in partially saturated porous media. Journal of the Institute of Mathematics and its Applications, 23(4):503– 517, 1979. CODEN JMTAA8. ISSN 0020-2932.

Fasano:1980:MPH

[287] Antonio Fasano, Mario Primicerio, and L. Rubinstein. A model problem for heat conduction with a free boundary in a concentrated capacity. Journal of the Institute of Mathematics and its Applications, 26(4):327–347, 1980. CODEN JMTAA8. ISSN 0020-2932.

Faulkner:1965:DEP

[288] T. R. Faulkner. Diffraction of an electromagnetic plane-wave by a metallic strip. Journal of the Institute of Mathematics and its Applications, 1(??):149–163, 1965. CODEN JMTAA8. ISSN 0020-2932.

Fenton:1973:SRS

[289] J. D. Fenton. Some results for surface gravity waves on shear flows. Journal of the Institute of Mathematics and its Applications, 12(??):1–20, 1973. CODEN JMTAA8. ISSN 0020-2932.

Fitz-Gerald:1969:MTEa

[290] J. M. Fitz-Gerald and J. A. McGeough. Mathematical theory of electrochemical machining. I: Anodic smoothing. Journal of the Institute of Mathematics and its Applications, 5(??):387–408, 1969. CODEN JMTAA8. ISSN 0020-2932.

Fitz-Gerald:1969:MTEb

[291] J. M. Fitz-Gerald, J. A. McGeough, and L. McL. Marsh. Mathematical theory of electrochemical machining. II: Anodic shaping. Journal of the Institute of Mathematics and its Applications, 5 (??):409–421, 1969. CODEN JMTAA8. ISSN 0020-2932.

Fletcher:1969:TO

[292] R. Fletcher. A technique for orthogonalization. Journal of the Institute of Mathematics and its Applications, 5(??): 162–166, 1969. CODEN JMTAA8. ISSN 0020-2932.

Fletcher:1971:GQP

[293] R. Fletcher. A general quadratic programming algorithm. Journal of the Institute of Mathematics and its Applications, 7(??):76–91, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Fletcher:1975:IPF

[294] R. Fletcher. An ideal penalty function for constrained optimization. Journal of the Institute of Mathematics and its Applications, 15(??):319–342, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Fletcher:1974:MQF

[295] R. Fletcher and M. P. Jackson. Minimization of a quadratic function of many variables subject only to lower and upper bounds. Journal of the Institute of Mathematics and its Applications, 14 (??):159–174, 1974. CODEN JMTAA8. ISSN 0020-2932.

Ford:1975:WLS

[296] J. A. Ford. Weighted least-squares polynomial deflation. Journal of the Institute of Mathematics and its Applications, 16 (3):263–269, 1975. CODEN JMTAA8. ISSN 0020-2932.

Forsyth:1979:STD

[297] P. Forsyth, Jr. and H. Rasmussen. Solution of time dependent electrochemical machining problems by a co- ordinate transformation. Journal of the Institute of Mathematics and its Applications, 24 (??):411–424, 1979. CODEN JMTAA8. ISSN 0020-2932.

Foschini:1975:DPD

[298] G. J. Foschini and A. C. Salazar. Data pulse design — intersymbol interference aspects. Journal of the Institute of Mathematics and its Applications, 15 (??):225–237, 1975. CODEN JMTAA8. ISSN 0020-2932.

Foster:1971:DCV

[299] K. Foster and R. Anderson. Dual and complementary variational principles for two dimensional electrostatic problems. Journal of the Institute of Mathematics and its Applications, 8(??):221–224, 1971. CODEN JMTAA8. ISSN 0020-2932.

Fowler:1978:LNL

[300] A. C. Fowler. Linear and non-linear stability of heat exchangers. Journal of the Institute of Mathematics and its Applications, 22(??):361–382, 1978. CODEN JMTAA8. ISSN 0020-2932.

Fowler:1978:LNS

[301] A. C. Fowler. Linear and nonlinear stability of heat exchangers. Journal of the Institute of Mathematics and its Applications, 22(3):361–382, 1978. CODEN JMTAA8. ISSN 0020-2932.

Fox:1971:FDE

[302] L. Fox, D. F. Mayers, J. R. Ockendon, and A. B. Tayler. On a functional differential equation. Journal of the Institute of Mathematics and its Applications, 8 (??):271–307, 1971. CODEN JMTAA8. ISSN 0020-2932.

Fox:1969:BSL

[303] L. Fox and R. Sankar. Boundary singularities in linear elliptic differential equations. Journal of the Institute of Mathematics and its Applications, 5(??):340– 350, 1969. CODEN JMTAA8. ISSN 0020-2932.

Fox:1973:RFM

[304] L. Fox and R. Sankar. The Regula-Falsi method for free-boundary problems. Journal of the Institute of Mathematics and its Applications, 12(??):49– 54, 1973. CODEN JMTAA8. ISSN 0020-2932.

Fox:1969:LTE

[305] N. Fox. Low temperature effects and generalized thermoelasticity. Journal of the Institute of Mathematics and its Applications, 5(4):373–386, 1969. CODEN JMTAA8. ISSN 0020-2932.

Fradkin:1977:CEP

[306] L. Ju. Fradkin and G. C. Wake. The critical explosion parameter in the theory of thermal ignition. Journal of the Institute of Mathematics and its Applications, 20(4):471–484, 1977. CODEN JMTAA8. ISSN 0020-2932.

Franke:1977:LDS

[307] R. Franke. Locally determined smooth interpolation at irregularly spaced points in several variables. Journal of the Institute of Mathematics and its Applications, 19(4):471–482, 1977. CODEN JMTAA8. ISSN 0020-2932.

Freeman:1974:CRVb

[308] T. L. Freeman and L. M. Delves. On the convergence rates of variational methods. III. Unsymmetric systems. Journal of the Institute of Mathematics and its Applications, 14(??):311–323, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Freeman:1975:RES

[309] T. L. Freeman and L. M. Delves. Roundoff errors and stability in variational calculations. Journal of the Institute of Mathematics and its Applications, 16(3): 345–353, 1975. CODEN JMTAA8. ISSN 0020-2932.

Freeman:1974:CRVa

[310] T. L. Freeman, L. M. Delves, and J. K. Reid. On the convergence rates of variational methods. II. Systems of type B, C. Journal of the Institute of Mathematics and its Applications, 14(??):145–157, 1974. CODEN JMTAA8. ISSN 0020-2932.

Frey:1977:EBT

[311] A. Frey, C. A. Hall, and J. C. Heinrich. On an exact boundary technique. Journal of the Institute of Mathematics and its Applications, 20(2):191–198, 1977. CODEN JMTAA8. ISSN 0020-2932.

Frost:1975:PFD

[312] M. G. Frost and A. Sackfield. Polynomials of the form _{2}F_{0} as determinants. Journal of the Institute of Mathematics and its Applications, 16(3):389–392, 1975. CODEN JMTAA8. ISSN 0020-2932.

Fryer:1976:SEA

[313] M. J. Fryer. Some errors associated with the non-parametric estimation of density functions. Journal of the Institute of Mathematics and its Applications, 18 (??):371–380, 1976. CODEN JMTAA8. ISSN 0020-2932.

Fryer:1977:RSN

[314] M. J. Fryer. A review of some nonparametric methods of density estimation. Journal of the Institute of Mathematics and its Applications, 20(3):335– 354, 1977. CODEN JMTAA8. ISSN 0020-2932.

Furzeland:1980:CSN

[315] R. M. Furzeland. A comparative study of numerical methods for moving boundary problems. Journal of the Institute of Mathematics and its Applications, 26(4): 411–429, 1980. CODEN JMTAA8. ISSN 0020-2932.

Fyfe:1971:LDR

[316] D. J. Fyfe. Linear dependence relations connecting equal interval Nth degree splines and their derivatives. Journal of the Institute of Mathematics and its Applications, 7(??):398–406, 1971. CODEN JMTAA8. ISSN 0020-2932.

Fyfe:1979:URF

[317] D. J. Fyfe and K. H. Oke. Use of residuals in Forsythe’s method for polynomial curve fitting. Journal of the Institute of Mathematics and its Applications, 24(1): 99–108, 1979. CODEN JMTAA8. ISSN 0020-2932.

Gaffney:1976:CII

[318] P. W. Gaffney. The calculation of indefinite integrals of B-splines. Journal of the Institute of Mathematics and its Applications, 17(1):37–41, 1976. CODEN JMTAA8. ISSN 0020-2932.

Gaffney:1978:RPV

[319] P. W. Gaffney. The range of possible values of f(x). Journal of the Institute of Mathematics and its Applications, 21 (2):211–226, 1978. CODEN JMTAA8. ISSN 0020-2932.

Gane:1977:BHP

[320] C. R. Gane and A. R. Gourlay. Block hopscotch procedures for second order parabolic differential equations. Journal of the Institute of Mathematics and its Applications, 19(2):205–216, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Gaster:1975:SAR

[321] M. Gaster and J. B. Roberts. Spectral analysis of randomly sampled signals. Journal of the Institute of Mathematics and its Applications, 15(??):195– 216, 1975. CODEN JMTAA8. ISSN 0020-2932.

Gavalas:1980:STA

[322] G. R. Gavalas and Y. C. Yortsos. Shorttime asymptotic solutions of the heat conduction equation with spatially varying coefficients. Journal of the Institute of Mathematics and its Applications, 26 (3):209–219, 1980. CODEN JMTAA8. ISSN 0020-2932.

Gelder:1976:NCV

[323] D. Gelder. The numerical calculation of viscous stratified flows. Journal of the Institute of Mathematics and its Applications, 17(2):225–231, 1976. CODEN JMTAA8. ISSN 0020-2932.

Genin:1966:STC

[324] J. Genin and J. S. Maybee. A stability theorem for a class of damped dynamic systems and some applications. Journal of the Institute of Mathematics and its Applications, 2(??):343–357, 1966. CO-DEN JMTAA8. ISSN 0020-2932.

Genin:1971:NLS

[325] Joseph Genin and John S. Maybee. Nonconservative linear systems with constant coefficients. Journal of the Institute of Mathematics and its Applications, 8(??):358–370, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Gentleman:1973:LSC

[326] W. Morven Gentleman. Least squares computations by Givens transformations without square roots. Journal of the Institute of Mathematics and its Applications, 12(??):329–336, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Getz:1977:SCF

[327] W. M. Getz and D. H. Jacobson. Sufficiency conditions for finite escape times in systems of quadratic differential equations. Journal of the Institute of Mathematics and its Applications, 19(4):377– 383, 1977. CODEN JMTAA8. ISSN 0020-2932.

Ghandour:1976:TTO

[328] E. Ghandour. Two-timing and onedimensional random wave propagation. Journal of the Institute of Mathematics and its Applications, 17(3):355–366, 1976. CODEN JMTAA8. ISSN 0020-2932.

Ghandour:1977:PDR

[329] E. Ghandour. Propagation of disturbances in a random medium. Journal of the Institute of Mathematics and its Applications, 20(4):393–407, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Gilder:1966:CPM

[330] J. Gilder. A class of pencils of matrices. Journal of the Institute of Mathematics and its Applications, 2(??):126– 132, 1966. CODEN JMTAA8. ISSN 0020-2932. See errata [331].

Gilder:1966:ECP

[331] J. Gilder. Errata: A class of pencils of matrices. Journal of the Institute of Mathematics and its Applications, 2(??): 283, 1966. CODEN JMTAA8. ISSN 0020-2932. See [330].

Gill:1972:QNM

[332] P. E. Gill and W. Murray. Quasi-Newton methods for unconstrained optimization. Journal of the Institute of Mathematics and its Applications, 9(??): 91–108, 1972. CODEN JMTAA8. ISSN 0020-2932.

Gittins:1967:GRL

[333] J. C. Gittins. A generalization of a result in linear programming. Journal of the Institute of Mathematics and its Applications, 3(??):193–201, 1967. CODEN JMTAA8. ISSN 0020-2932.

Gladwell:1975:EBC

[334] I. Gladwell and D. J. Mullings. On the effect of boundary conditions in collocation by polynomial splines for the solution of boundary value problems in ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 16(1):93–107, 1975. CODEN JMTAA8. ISSN 0020-2932.

Gladwell:1972:PEB

[335] Ian Gladwell. A posteriori error bounds for approximate solutions of linear second-order ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 9(??):323– 349, 1972. CODEN JMTAA8. ISSN 0020-2932.

Gladwell:1973:RRM

[336] Ian Gladwell. Rayleigh–Ritz methods for nonlinear boundary value problems. with an appendix: Analytical and numerical examples, by I. Gladwell and J. L. Lloyd. Journal of the Institute of Mathematics and its Applications, 11 (??):191–211, 1973. CODEN JMTAA8. ISSN 0020-2932.

Glen:1973:COC

[337] J. A. Glen. A criterion for an optimal control system to have finite switchings. Journal of the Institute of Mathematics and its Applications, 11(??):45–55, 1973. CODEN JMTAA8. ISSN 0020-2932.

Glover:1973:SNI

[338] Keith Glover and Jan C. Willems. On the stability of numerical integration routines for ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 11(??):171– 180, 1973. CODEN JMTAA8. ISSN 0020-2932.

Godwin:1979:SCL

[339] A. N. Godwin. Spline curves with length constraints. Journal of the Institute of Mathematics and its Applications, 24(2): 219–229, 1979. CODEN JMTAA8. ISSN 0020-2932.

Goodinson:1968:EMF

[340] P. A. Goodinson and R. A. Newing. Einstein–Maxwell fields with non-zero charge-current distribution. Journal of the Institute of Mathematics and its Applications, 4(??):270–275, 1968. CODEN JMTAA8. ISSN 0020-2932.

Goodinson:1969:EMN

[341] P. A. Goodinson and R. A. Newing. Einstein–Maxwell null fields with nonzero current distribution. Journal of the Institute of Mathematics and its Applications, 5(??):72–90, 1969. CODEN JM-TAA8. ISSN 0020-2932.

Goodinson:1970:EMF

[342] P. A. Goodinson and R. A. Newing. Einstein–Maxwell fields in the presence of matter and pressure. Journal of the Institute of Mathematics and its Applications, 6(??):212–221, 1970. CODEN JMTAA8. ISSN 0020-2932.

Gourlay:1970:CAP

[343] A. R. Gourlay. On Chebychev acceleration procedures for alternating direction iterative methods. Journal of the Institute of Mathematics and its Applications, 6(??):1–11, 1970. CODEN JM-TAA8. ISSN 0020-2932.

Gourlay:1971:GHA

[344] A. R. Gourlay and G. R. McGuire. General hopscotch algorithm for the numerical solution of partial differential equations. Journal of the Institute of Mathematics and its Applications, 7(??):216– 227, 1971. CODEN JMTAA8. ISSN 0020-2932.

Gourlay:1972:SAD

[345] A. R. Gourlay and A. R. Mitchell. On the structure of alternating direction implicit (A.D.I.) and locally one dimensional (L.O.D.) difference methods. Journal of the Institute of Mathematics and its Applications, 9(??):80–90, 1972. CODEN JMTAA8. ISSN 0020-2932.

Grad:1973:MEZ

[346] J. Grad and E. Zakrajšek. Method for evaluation of zeros of Bessel functions. Journal of the Institute of Mathematics and its Applications, 11(??):57–72, 1973. CODEN JMTAA8. ISSN 0020-2932.

Granoff:1966:ASH

[347] B. Granoff and R. M. Lewis. Asymptotic solution of hyperbolic systems with constant coefficients. Journal of the Institute of Mathematics and its Applications, 2(??):144–156, 1966. CODEN JM-TAA8. ISSN 0020-2932.

Grant:1974:NTO

[348] I. P. Grant and D. Ellison. A numerical treatment of one-dimensional models of radiative heat transfer in co-current tube furnaces. Journal of the Institute of Mathematics and its Applications, 13 (??):169–185, 1974. CODEN JMTAA8. ISSN 0020-2932.

Grant:1971:ACM

[349] J. A. Grant and G. D. Hitchins. An always convergent minimization technique for the solution of polynomial equations. Journal of the Institute of Mathematics and its Applications, 8(??):122–129, 1971. CODEN JMTAA8. ISSN 0020-2932.

Grattan-Guinness:1969:JFR

[350] I. Grattan-Guinness. Joseph Fourier and the revolution in mathematical physics. Journal of the Institute of Mathematics and its Applications, 5(??):230–253, 1969. CODEN JMTAA8. ISSN 0020-2932.

Graves-Morris:1978:PAL

[351] P. R. Graves-Morris. Padé approximants for linear integral equations? Journal of the Institute of Mathematics and its Applications, 21(??):375–378, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Graves-Morris:1980:PRR

[352] P. R. Graves-Morris. Practical, reliable, rational interpolation. Journal of the Institute of Mathematics and its Applications, 25(3):267–286, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Graves-Morris:1974:CSR

[353] P. R. Graves-Morris, R. Hughes Jones, and G. J. Makinson. The calculation of some rational approximants in two variables. Journal of the Institute of Mathematics and its Applications, 13(??):311– 320, 1974. CODEN JMTAA8. ISSN 0020-2932.

Green:1977:TET

[354] A. E. Green and J. Z. Mkrtichian. Torsion and extension of a tube with different moduli in tension and compression. Journal of the Institute of Mathematics and its Applications, 20(??):221–226, 1977. CODEN JMTAA8. ISSN 0020-2932.

Green:1978:NRC

[355] A. E. Green and D. Nicol. A note on reinforced cylindrical tubes. Journal of the Institute of Mathematics and its Applications, 21(??):227–235, 1978. CODEN JMTAA8. ISSN 0020-2932.

Green:1974:RH

[356] A. E. Green and M. Troth. Reinforced hosepipes. Journal of the Institute of Mathematics and its Applications, 13 (??):39–46, 1974. CODEN JMTAA8. ISSN 0020-2932.

Green:1971:PCO

[357] M. W. Green. A problem connected with the oblique incidence of surface waves on an immersed cylinder. Journal of the Institute of Mathematics and its Applications, 8(??):82–98, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Greenspan:1978:CSS

[358] D. Greenspan and J. Collier. Computer studies of swirling particle fluids and the evolution of planetary-type bodies. Journal of the Institute of Mathematics and its Applications, 22(??):235–253, 1978. CODEN JMTAA8. ISSN 0020-2932.

Greenspan:1972:NSF

[359] Donald Greenspan. Numerical studies of flow between rotating coaxial disks. Journal of the Institute of Mathematics and its Applications, 9(??):370–377, 1972. CODEN JMTAA8. ISSN 0020-2932.

Griffel:1976:ADP

[360] D. H. Griffel. Age-dependent population growth. Journal of the Institute of Mathematics and its Applications, 17(2): 141–152, 1976. CODEN JMTAA8. ISSN 0020-2932.

Griffiths:1977:NSS

[361] D. F. Griffiths. A numerical study of a singular elliptic boundary value problem. Journal of the Institute of Mathematics and its Applications, 19(1):59– 69, 1977. CODEN JMTAA8. ISSN 0020-

Griffiths:1976:MMP

[362] J. D. Griffiths and C. Cresswell. A mathematical model of a pelican crossing. Journal of the Institute of Mathematics and its Applications, 18(3):381–394, 1976. CODEN JMTAA8. ISSN 0020-2932.

Grimshaw:1968:PSW

[363] R. Grimshaw. Propagation of surface waves at high frequencies. Journal of the Institute of Mathematics and its Applications, 4(??):174–193, 1968. CODEN JMTAA8. ISSN 0020-2932.

Grundy:1977:LTI

[364] R. E. Grundy. Laplace transform inversion using two-point rational approximants. Journal of the Institute of Mathematics and its Applications, 20(3):299– 306, 1977. CODEN JMTAA8. ISSN 0020-2932.

Grundy:1978:SNL

[365] R. E. Grundy. On the solution of nonlinear Volterra integral equations using two-point rational approximants. Journal of the Institute of Mathematics and its Applications, 22(3):317–320, 1978. CODEN JMTAA8. ISSN 0020-2932.

Grundy:1978:SVI

[366] R. E. Grundy. The solution of Volterra integral equations of the convolution type using two-point rational approximants. Journal of the Institute of Mathematics and its Applications, 22(2):147– 158, 1978. CODEN JMTAA8. ISSN 0020-2932.

Gunson:1978:ASF

[367] J. Gunson and P. H. Ng. Approximation of special functions of complex variables.

Journal of the Institute of Mathematics and its Applications, 22(1):43–52, 1978. CODEN JMTAA8. ISSN 0020-2932.

Guttmann:1975:DFR

[368] A. J. Guttmann. Derivation of 1/2 mimic functions1/2 from regular perturbation expansions in fluid mechanics. Journal of the Institute of Mathematics and its Applications, 15(??):307–317, 1975. CODEN JMTAA8. ISSN 0020-2932.

Hackbusch:1980:FNS

[369] Wolfgang Hackbusch. The fast numerical solution of very large elliptic difference schemes. Journal of the Institute of Mathematics and its Applications, 26 (2):119–132, 1980. CODEN JMTAA8. ISSN 0020-2932.

Hadjidimos:1970:SFD

[370] A. Hadjidimos. Solving the fourdimensional Laplace’s equation by using extrapolated alternating direction implicit methods. Journal of the Institute of Mathematics and its Applications, 6 (??):232–240, 1970. CODEN JMTAA8. ISSN 0020-2932.

Hadjidimos:1971:OEA

[371] Apostolos Hadjidimos. Optimum extrapolated alternating direction implicit schemes in the presence of singular matrices. Journal of the Institute of Mathematics and its Applications, 7(??):361– 366, 1971. CODEN JMTAA8. ISSN 0020-2932.

Hall:1978:FES

[372] C. A. Hall and J. Heinrich. A finite element that satisfies natural boundary conditions exactly. Journal of the Institute of Mathematics and its Applications, 21(??):237–250, 1978. CODEN JMTAA8. ISSN 0020-2932.

Hall:1973:CRA

[373] C. A. Hall and J. W. Kennedy. Convergence of a Ritz approximation for the steady state heat flow problem. Journal of the Institute of Mathematics and its Applications, 12(??):187–196, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Hammarling:1974:NMG

[374] Sven Hammarling. A note on modifications to the Givens plane rotation. Journal of the Institute of Mathematics and its Applications, 13(??):215–218, 1974. CODEN JMTAA8. ISSN 0020-2932.

Handscomb:1973:RST

[375] D. C. Handscomb. The relative sizes of the terms in Chebyshev and other ultraspherical expansions. Journal of the Institute of Mathematics and its Applications, 11(??):241–246, 1973. CODEN JMTAA8. ISSN 0020-2932.

Handscomb:1977:VSW

[376] D. C. Handscomb. Vibration of a submerged window. Journal of the Institute of Mathematics and its Applications, 20 (??):183–189, 1977. CODEN JMTAA8. ISSN 0020-2932.

Hannabuss:1974:ICH

[377] K. C. Hannabuss. The irreducible components of homogeneous functions and symmetric tensors. Journal of the Institute of Mathematics and its Applications, 14(??):83–88, 1974. CODEN JM-TAA8. ISSN 0020-2932.

Hansen:1974:MBP

[378] E. Hansen and P. Hougaard. On a moving boundary problem from biomechanics. Journal of the Institute of Mathematics and its Applications, 13(??):385– 398, 1974. CODEN JMTAA8. ISSN 0020-2932.

Hargrave:1974:UAEb

[379] B. A. Hargrave. Uniform asymptotic expansions for ellipsoidal wave functions. II: Ellipsoidal wave functions. Journal of the Institute of Mathematics and its Applications, 14(??):189–196, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Hargrave:1974:NST

[380] B. A. Hargrave and B. D. Sleeman. The numerical solution of two-parameter eigenvalue problems in ordinary differential equations with an application to the problem of diffraction by a plane angular sector. Journal of the Institute of Mathematics and its Applications, 14(??):9–22, 1974. CODEN JMTAA8. ISSN 0020-2932.

Hargrave:1974:UAEa

[381] B. A. Hargrave and B. D. Sleeman. Uniform asymptotic expansions for ellipsoidal wave functions. I. High frequency solutions of the ellipsoidal wave equations. Journal of the Institute of Mathematics and its Applications, 14(??):31– 40, 1974. CODEN JMTAA8. ISSN 0020-2932.

Harley:1976:ECP

[382] P. J. Harley and A. R. Mitchell. Exact control of a parabolic differential equation using an implicit Runge–Kutta method. Journal of the Institute of Mathematics and its Applications, 18(1): 9–14, 1976. CODEN JMTAA8. ISSN 0020-2932.

Hart:1973:QNM

[383] W. E. Hart and S. O. W. Soul. Quasi-Newton methods for discretized nonlinear boundary problems. Journal of the Institute of Mathematics and its Applications, 11(??):351–359, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Hawkins:1972:CSP

[384] Douglas M. Hawkins. On the choice of segments in piecewise approximation. Journal of the Institute of Mathematics and its Applications, 9(??):250–256, 1972. CODEN JMTAA8. ISSN 0020-2932.

Hayes:1974:LSF

[385] J. G. Hayes and J. Halliday. The leastsquares fitting of cubic spline surfaces to general data sets. Journal of the Institute of Mathematics and its Applications, 14(??):89–103, 1974. CODEN JM-TAA8. ISSN 0020-2932.

Hayes:1976:SNL

[386] L. Hayes and E. Wasserstrom. Solution of non-linear eigenvalue problems by the continuation method. Journal of the Institute of Mathematics and its Applications, 17(1):5–14, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Hazlewood:1977:ASC

[387] L. J. Hazlewood. An alternative strategy for cautious adaptive quadrature. Journal of the Institute of Mathematics and its Applications, 20(4):505–518, 1977. CODEN JMTAA8. ISSN 0020-2932.

Heading:1978:CTC

[388] J. Heading. Calculation of transmission coefficients for generalized propagation problems. Journal of the Institute of Mathematics and its Applications, 21 (??):387–400, 1978. CODEN JMTAA8. ISSN 0020-2932.

Heading:1980:PSG

[389] J. Heading. Properties of some generalized hypergeometric functions governing coupling at transition points. Journal of the Institute of Mathematics and its Applications, 25(3):311–324, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Heading:1968:WPG

[390] John Heading. Wave propagation governed by an infinite number of transition points. Journal of the Institute of Mathematics and its Applications, 4(??): 123–139, 1968. CODEN JMTAA8. ISSN 0020-2932.

Heading:1972:FEA

[391] John Heading. Further exact and approximate considerations of the barrier problem. Journal of the Institute of Mathematics and its Applications, 10 (??):312–324, 1972. CODEN JMTAA8. ISSN 0020-2932.

Hegedus:1979:DMC

[392] C. J. Hegedus, G. Kadar, and E. Della Torre. Demagnetization matrices for cylindrical bodies. Journal of the Institute of Mathematics and its Applications, 24(??):279–291, 1979. CODEN JMTAA8. ISSN 0020-2932.

Henderson:1969:PAT

[393] J. C. Henderson and E. A. W. Maunder.

A problem in applied topology: On the selection of cycles for the flexibility analysis of skeletal structures. Journal of the Institute of Mathematics and its Applications, 5(??):254–269, 1969. CODEN JMTAA8. ISSN 0020-2932.

Henery:1974:SWH

[394] R. J. Henery. Solution of Wiener–Hopf integral equations using the Fast Fourier Transform. Journal of the Institute of Mathematics and its Applications, 13 (??):89–96, 1974. CODEN JMTAA8. ISSN 0020-2932.

Henery:1977:SFE

[395] R. J. Henery. Solution of Fredholm equations with symmetric difference kernels. Journal of the Institute of Mathematics and its Applications, 19(1):29–37, 1977. CODEN JMTAA8. ISSN 0020-2932.

Herbert:1966:SEB

[396] R. E. Herbert. Stability of elastic bodies subjected to body and pressure forces with specialization to thin shells. Journal of the Institute of Mathematics and its Applications, 2(??):248–265, 1966. CODEN JMTAA8. ISSN 0020-2932.

Herman:1980:SEA

[397] G. T. Herman, A. Lent, and H. Hurwitz. A storage-efficient algorithm for finding the regularized solution of a large, inconsistent system of equations. Journal of the Institute of Mathematics and its Applications, 25(4):361–366, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Herrera:1980:VPP

[398] I. Herrera. Variational principles for problems with linear constraints: prescribed jumps and continuation type restrictions. Journal of the Institute of Mathematics and its Applications, 25(1): 67–96, 1980. CODEN JMTAA8. ISSN 0020-2932.

Herrera:1978:DEP

[399] I. Herrera and M. J. Sewell. Dual extremum principles for non-negative unsymmetric operators. Journal of the Institute of Mathematics and its Applications, 21(??):95–115, 1978. CODEN JM-TAA8. ISSN 0020-2932.

Hill:1970:ATG

[400] W. Hill and P. R. Wallis. The application of the theory of games to mine countermeasures tactics. Journal of the Institute of Mathematics and its Applications, 6(??):27–41, 1970. CODEN JM-TAA8. ISSN 0020-2932.

Hillion:1977:RRA

[401] P. Hillion. Remarks on rational approximations of multiple power series. Journal of the Institute of Mathematics and its Applications, 19(3):281–293, 1977. CODEN JMTAA8. ISSN 0020-2932.

Hocking:1971:DDPa

[402] L. M. Hocking. Diffusion from disjoint plane domains. I. The interaction of two domains. Journal of the Institute of Mathematics and its Applications, 8(??): 139–152, 1971. CODEN JMTAA8. ISSN 0020-2932.

Hocking:1971:DDPb

[403] L. M. Hocking. Diffusion from disjoint plane domains. II. Arrays of domains, with application to the response of a muscle to a drug. Journal of the Institute of Mathematics and its Applications, 8 (??):153–163, 1971. CODEN JMTAA8. ISSN 0020-2932.

Holt:1979:ACN

[404] J. N. Holt and R. Fletcher. An algorithm for constrained non-linear least-squares. Journal of the Institute of Mathematics and its Applications, 23(4):449–463, 1979. CODEN JMTAA8. ISSN 0020-

Honda:1965:SFC

[405] M. Honda. Stream-function co-ordinates in rotational flow and an analysis of the flow in shock layer. I. Two-dimensional flow. Journal of the Institute of Mathematics and its Applications, 1(??):127– 148, 1965. CODEN JMTAA8. ISSN 0020-2932.

Hopgood:1974:IRF

[406] D. J. Hopgood and S. McKee. Improved root-finding methods derived from inverse interpolation. Journal of the Institute of Mathematics and its Applications, 14(??):217–228, 1974. CODEN JMTAA8. ISSN 0020-2932.

Hoshino:1972:FVM

[407] S. Hoshino. A formulation of variable metric methods. Journal of the Institute of Mathematics and its Applications, 10 (??):394–403, 1972. CODEN JMTAA8. ISSN 0020-2932.

Hoshino:1973:MCA

[408] Satoshi Hoshino. A method of centres algorithm with arbitrary parameter and its application. Journal of the Institute of Mathematics and its Applications, 12 (??):319–328, 1973. CODEN JMTAA8. ISSN 0020-2932.

Howard:1971:ASV

[409] I. C. Howard. An approximation for slowly varying waves in elastic rods. Journal of the Institute of Mathematics and its Applications, 7(??):47–60, 1971. CODEN JMTAA8. ISSN 0020-2932.

Howe:1973:ANM

[410] D. Howe. The application of numerical methods to the conformal transformation of polygonal boundaries. Journal of the Institute of Mathematics and its Applications, 12(??):125–136, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Huang:1979:IJT

[411] C. P. Huang. The implementation of a Jacobi-type method. Journal of the Institute of Mathematics and its Applications, 23(4):497–501, 1979. CODEN JM-TAA8. ISSN 0020-2932.

HughesJones:1974:GCR

[412] R. Hughes Jones and G. J. Makinson. The generation of Chisholm rational polynomial approximants to power series in two variables. Journal of the Institute of Mathematics and its Applications, 13(??):299–310, 1974. CODEN JMTAA8. ISSN 0020-2932.

Hunt:1967:TER

[413] G. E. Hunt. The transport equation of radiative transfer in a three-dimensional space with anisotropic scattering. Journal of the Institute of Mathematics and its Applications, 3(??):181–192, 1967. CODEN JMTAA8. ISSN 0020-2932.

Ingham:1975:PSC

[414] D. B. Ingham. Plane sub-critical flow past a lifting aerofoil. Journal of the Institute of Mathematics and its Applications, 16(??):73–80, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Ingham:1977:FPI

[415] D. B. Ingham. Flow past an impulsively started semi-infinite flat plate in the presence of a magnetic field. Journal of the Institute of Mathematics and its Applications, 20(4):459–469, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Ingham:1978:FPS

[416] D. B. Ingham. Flow past a suddenly cooled vertical plate. Journal of the Institute of Mathematics and its Applications, 22(2):189–196, 1978. CODEN JM-TAA8. ISSN 0020-2932.

Iordanidis:1971:NST

[417] K. I. Iordanidis. Numerical solution of the third boundary value problem for Laplace’s equation. Journal of the Institute of Mathematics and its Applications, 8(??):308–314, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Iordanidis:1973:NSR

[418] K. I. Iordanidis. Numerical solution of the Robin problem by extrapolated A. D. I. methods. Journal of the Institute of Mathematics and its Applications, 12 (??):91–96, 1973. CODEN JMTAA8. ISSN 0020-2932.

Iordanidis:1975:UCE

[419] K. I. Iordanidis. Using crude eigenvalue bounds in the numerical solution of the third boundary value problem in 2- and 3-dimensional space. Journal of the Institute of Mathematics and its Applications, 15(??):249–256, 1975. CODEN JMTAA8. ISSN 0020-2932.

Iyengar:1978:HAD

[420] S. R. K. Iyengar and R. C. Mittal. High accuracy difference schemes for the cylindrical heat conduction equation. Journal of the Institute of Mathematics and its Applications, 22(3):321– 330, 1978. CODEN JMTAA8. ISSN 0020-2932.

Jain:1974:HOD

[421] M. K. Jain and Settaluri R. K. Iyengar. Higher order difference formulae for the numerical solution of the heat conduction equation. Journal of the Institute of Mathematics and its Applications, 13 (??):147–151, 1974. CODEN JMTAA8. ISSN 0020-2932.

Jain:1971:ORK

[422] R. K. Jain and M. K. Jain. Optimum Runge–Kutta Fehlberg methods for first order differential equations. Journal of the Institute of Mathematics and its Applications, 8(??):386–396, 1971. CODEN JMTAA8. ISSN 0020-2932.

Jain:1972:ORK

[423] R. K. Jain and M. K. Jain. Optimum Runge–Kutta–Fehlberg methods for second-order differential equations. Journal of the Institute of Mathematics and its Applications, 10(??):202–210, 1972. CODEN JMTAA8. ISSN 0020-2932.

James:1974:TOC

[424] Eleanor M. James. Time optimal control and the van der Pol oscillator. Journal of the Institute of Mathematics and its Applications, 13(??):67–81, 1974. CODEN JMTAA8. ISSN 0020-2932.

Jarratt:1974:ABC

[425] P. Jarratt and N. Parkin. On the asymptotic behaviour of a certain non-linear difference equation. Journal of the Institute of Mathematics and its Applications, 13(??):83–88, 1974. CODEN JM-TAA8. ISSN 0020-2932.

Jarvis:1971:SSW

[426] R. J. Jarvis. The scattering of surface waves by two vertical plane barriers. Journal of the Institute of Mathematics and its Applications, 7(??):207– 215, 1971. CODEN JMTAA8. ISSN 0020-2932.

Jeffrey:1978:TFE

[427] D. J. Jeffrey. The temperature field or electric potential around two almost touching spheres. With an appendix by M. Van Dyke. Journal of the Institute of Mathematics and its Applications, 22 (3):337–351, 1978. CODEN JMTAA8. ISSN 0020-2932.

Jeffreys:1977:PTG

[428] Harold Jeffreys. Probability theory in geophysics. Journal of the Institute of Mathematics and its Applications, 19(1): 87–96, 1977. CODEN JMTAA8. ISSN 0020-2932.

Jennings:1977:IES

[429] A. Jennings. Influence of the eigenvalue spectrum on the convergence rate of the conjugate gradient method. Journal of the Institute of Mathematics and its Applications, 20(1):61–72, 1977. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Jennings:1978:SLG

[430] A. Jennings, J. Halliday, and M. J. Cole. Solution of linear generalized eigenvalue problems containing singular matrices. Journal of the Institute of Mathematics and its Applications, 22(4):401–410, 1978. CODEN JMTAA8. ISSN 0020-2932.

Jennings:1977:PE

[431] A. Jennings and G. M. Malik. Partial elimination. Journal of the Institute of Mathematics and its Applications, 20(3): 307–316, 1977. CODEN JMTAA8. ISSN 0020-2932.

Jennings:1975:SIP

[432] A. Jennings and W. J. Stewart. Simultaneous iteration for partial eigensolution of real matrices. Journal of the Institute of Mathematics and its Applications, 15 (??):351–361, 1975. CODEN JMTAA8. ISSN 0020-2932.

Jennings:1971:ACM

[433] Alan Jennings. Accelerating the convergence of matrix iterative processes. Journal of the Institute of Mathematics and its Applications, 8(??):99–110, 1971. CODEN JMTAA8. ISSN 0020-2932.

Jennings:1977:GEP

[434] L. S. Jennings and M. R. Osborne. Generalized eigenvalue problems for rectangular matrices. Journal of the Institute of Mathematics and its Applications, 20 (4):443–458, 1977. CODEN JMTAA8. ISSN 0020-2932.

Johnson:1973:EAN

[435] L. W. Johnson and R. D. Riess. An error analysis for numerical differentiation. Journal of the Institute of Mathematics and its Applications, 11(??):115–120, 1973. CODEN JMTAA8. ISSN 0020-2932.

Johnson:1979:GHI

[436] L. W. Johnson and R. D. Riess. Gauss harmonic interpolation formulas for circular regions. Journal of the Institute of Mathematics and its Applications, 23 (2):215–222, March 1979. CODEN JM-TAA8. ISSN 0020-2932.

Johnson:1979:BOP

[437] M. A. Johnson. Bounds for an optimization problem with a linear programming structure. Journal of the Institute of Mathematics and its Applications, 24(4): 395–410, 1979. CODEN JMTAA8. ISSN 0020-2932.

Jones:1976:EHO

[438] B. Jones. Extrapolation for higher orders of convergence. Journal of the Institute of Mathematics and its Applications, 17(1):27–36, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Jones:1965:SRH

[439] D. S. Jones. Some remarks on Hilbert transforms. Journal of the Institute of Mathematics and its Applications, 1(??): 226–240, 1965. CODEN JMTAA8. ISSN 0020-2932.

Jones:1966:FTM

[440] D. S. Jones. Fourier transforms and the method of stationary phase. Journal of the Institute of Mathematics and its Applications, 2(??):197–222, 1966. CODEN JMTAA8. ISSN 0020-2932.

Jones:1972:ASD

[441] D. S. Jones. Aerodynamic sound due to a source near a half-plane. Journal of the Institute of Mathematics and its Applications, 9(??):114–122, 1972. CODEN JMTAA8. ISSN 0020-2932.

Jones:1973:VMA

[442] D. S. Jones. The variable metric algorithm for non-definite quadratic functions. Journal of the Institute of Mathematics and its Applications, 12(??):63– 71, 1973. CODEN JMTAA8. ISSN 0020-2932.

Jones:1975:SSV

[443] D. S. Jones. The scattering of sound by a vortex sheet. Journal of the Institute of Mathematics and its Applications, 15(??):33–57, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Jones:1978:ASP

[444] D. S. Jones. Acoustics of a splitter plate.

Journal of the Institute of Mathematics and its Applications, 21(??):197–209, 1978. CODEN JMTAA8. ISSN 0020-2932.

Jones:1979:LFE

[445] D. S. Jones. Low frequency electromagnetic radiation. Journal of the Institute of Mathematics and its Applications, 23 (4):421–447, 1979. CODEN JMTAA8. ISSN 0020-2932.

Jones:1980:KLT

[446] D. S. Jones. The Kontorovich–Lebedev transform. Journal of the Institute of Mathematics and its Applications, 26(2): 133–141, 1980. CODEN JMTAA8. ISSN 0020-2932.

Jones:1979:NID

[447] D. S. Jones and G. G. Swinerd. A note on the intersection of dispersion curves of a fluid-loaded elastic plate. Journal of the Institute of Mathematics and its Applications, 23(??):311–319, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Joslin:1975:MNL

[448] T. Joslin and S. McKee. A mildly non-linear system of parabolic differential equations arising from mass transfer with chemical reaction in electrochemistry. Journal of the Institute of Mathematics and its Applications, 15(??):373– 383, 1975. CODEN JMTAA8. ISSN 0020-2932.

Kalker:1977:SDE

[449] J. J. Kalker. The surfaces displacement of an elastic half-space loaded in a slender, bounded, curved surface region with application to the calculation of the contact pressure under a roller. Journal of the Institute of Mathematics and its Applications, 19(??):127–144, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Kalker:1977:VPC

[450] J. J. Kalker. Variational principles of contact elastostatics. Journal of the Institute of Mathematics and its Applications, 20(2):199–219, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Kambo:1971:ECG

[451] N. S. Kambo. Error of certain Gauss quadrature formulas. Journal of the Institute of Mathematics and its Applications, 7(??):303–307, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Kambo:1973:NAM

[452] N. S. Kambo. A note on asymptotic minimal weights for an integration rule. Journal of the Institute of Mathematics and its Applications, 12(??):159–164, 1973. CODEN JMTAA8. ISSN 0020-2932.

Karlsson:1971:UPE

[453] Per W. Karlsson. On the uniqueness and propagation of the electromagnetic field. Journal of the Institute of Mathematics and its Applications, 8(??):1–7, 1971. CODEN JMTAA8. ISSN 0020-2932.

Kartsatos:1975:BEP

[454] A. G. Kartsatos and J. R. Ward. Boundedness and existence of periodic solutions of quasi-linear systems. Journal of the Institute of Mathematics and its Applications, 15(??):187–194, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Keast:1979:SFS

[455] P. Keast. Some fully symmetric quadrature formulae for product spaces. Journal of the Institute of Mathematics and its Applications, 23(3):251–264, 1979. CODEN JMTAA8. ISSN 0020-2932.

Keller:1972:DMR

[456] J. B. Keller, R. E. Kleinman, and T. B. A. Senior. Dipole moments in Rayleigh scattering. Journal of the Institute of Mathematics and its Applications, 9(??):14–22, 1972. CODEN JM-TAA8. ISSN 0020-2932.

Kershaw:1973:TIC

[457] D. Kershaw. Two interpolatory cubic splines. Journal of the Institute of Mathematics and its Applications, 11 (??):329–333, 1973. CODEN JMTAA8. ISSN 0020-2932.

Kershaw:1977:BIB

[458] D. Kershaw. A bound on the inverse of a band matrix which occurs in interpolation by periodic odd order splines. Journal of the Institute of Mathematics and its Applications, 20(2):227–228, 1977. CODEN JMTAA8. ISSN 0020-2932.

Khalil:1975:TPF

[459] H. M. Khalil and J. H. Giese. A two-parameter family of two-level higher order difference methods for the twodimensional heat equation. Journal of the Institute of Mathematics and its Applications, 16(2):193–205, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Knibb:1973:NSP

[460] D. Knibb. The numerical solution of parabolic partial differential equations using the method of Lanczos. Journal of the Institute of Mathematics and its Applications, 11(??):181–190, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Knibb:1978:CMN

[461] D. Knibb and R. E. Scraton. A collocation method for the numerical solution of non-linear parabolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 22 (3):305–315, 1978. CODEN JMTAA8. ISSN 0020-2932.

Kotorynski:1979:SVC

[462] W. P. Kotorynski. Slowly varying channel flows in three dimensions. Journal of the Institute of Mathematics and its Applications, 24(1):71–80, 1979. CODEN JMTAA8. ISSN 0020-2932.

Krenk:1978:NQP

[463] S. Krenk. Numerical quadrature of periodic singular integral equations. Journal of the Institute of Mathematics and its Applications, 21(??):181–187, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Krenk:1978:QFC

[464] S. Krenk. Quadrature formulae of closed type for solution of singular integral equations. Journal of the Institute of Mathematics and its Applications, 22(1): 99–107, 1978. CODEN JMTAA8. ISSN 0020-2932.

Kubicek:1973:GPM

[465] M. Kubíček and V. Hlaváček. General parameter mapping technique — a procedure of solution of non-linear boundary value problems depending on an actual parameter. Journal of the Institute of Mathematics and its Applications, 12 (??):287–293, 1973. CODEN JMTAA8. ISSN 0020-2932.

Kumar:1974:CGQ

[466] R. Kumar. Certain Gaussian quadratures. Journal of the Institute of Mathematics and its Applications, 14(??):175– 182, 1974. CODEN JMTAA8. ISSN 0020-2932.

Kumar:1980:NQF

[467] S. Kumar. A note on quadrature formulae for Cauchy principal value integrals. Journal of the Institute of Mathematics and its Applications, 26(4):447– 451, 1980. CODEN JMTAA8. ISSN 0020-2932.

Kydes:1977:HMD

[468] A. S. Kydes and R. P. Tewarson. Householder modification for damped least squares solutions. Journal of the Institute of Mathematics and its Applications, 19(4):407–423, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Lam:1978:SLS

[469] C. F. Lam and J. Pedersen. Solving large sets of differential equations describing queueing systems with the aid of a dynamic allocation algorithm. Journal of the Institute of Mathematics and its Applications, 22(4):435–441, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Lambert:1976:SMM

[470] J. D. Lambert and I. A. Watson. Symmetric multistep methods for periodic initial value problems. Journal of the Institute of Mathematics and its Applications, 18(2):189–202, 1976. CODEN JMTAA8. ISSN 0020-2932.

Lancaster:1976:IRF

[471] P. Lancaster. Interpolation in a rectangle and finite elements of high degree.

Journal of the Institute of Mathematics and its Applications, 18(1):67–77, 1976. CODEN JMTAA8. ISSN 0020-2932.

Lardner:1971:DMC

[472] R. W. Lardner. Dislocations in materials with couple stress. Journal of the Institute of Mathematics and its Applications, 7(??):126–137, 1971. CODEN JMTAA8. ISSN 0020-2932.

Lardner:1976:NLD

[473] R. W. Lardner and G. N. Bojadziev. Non-linear differential equations with damping and large time delay. Journal of the Institute of Mathematics and its Applications, 18(1):25–35, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Lashiher:1978:CTF

[474] A. M. Ben Lashiher. Converse theorems for final-stability. Journal of the Institute of Mathematics and its Applications, 21(??):127–134, 1978. CODEN JMTAA8. ISSN 0020-2932.

Laurie:1977:OPC

[475] D. P. Laurie. Optimal parameter choice in the θ-method for parabolic equations. Journal of the Institute of Mathematics and its Applications, 19(1):119–125, 1977. CODEN JMTAA8. ISSN 0020-2932.

Lavenda:1977:NLD

[476] B. H. Lavenda, G. Peluso, and A. di Chiara. Non-linear dissipative and dispersive waves. Journal of the Institute of Mathematics and its Applications, 20 (1):117–131, 1977. CODEN JMTAA8. ISSN 0020-2932.

Lavery:1979:PEB

[477] J. E. Lavery. A posteriori error bounds for linear elliptic equations of potential type. Journal of the Institute of Mathematics and its Applications, 23(4):465– 472, 1979. CODEN JMTAA8. ISSN 0020-2932.

Laws:1975:UET

[478] N. Laws. The use of energy theorems to obtain upper and lower bounds. Journal of the Institute of Mathematics and its Applications, 15(??):109–119, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Lee:1971:SSD

[479] C. F. Lee. On the solution of some diffusion equations with concentrationdependent diffusion coefficients. Journal of the Institute of Mathematics and its Applications, 8(??):251–259, 1971. CO-DEN JMTAA8. ISSN 0020-2932.

Lee:1972:SSD

[480] C. F. Lee. On the solution of some diffusion equations with concentrationdependent diffusion coefficients. II. Journal of the Institute of Mathematics and its Applications, 10(??):129–133, 1972. CODEN JMTAA8. ISSN 0020-2932.

Leppington:1970:RSS

[481] F. G. Leppington. On the radiation of short surface waves by a finite dock. Journal of the Institute of Mathematics and its Applications, 6(??):319–340, 1970. CODEN JMTAA8. ISSN 0020-2932.

Levey:1968:RAL

[482] H. C. Levey and J. J. Mahony. Resonance in almost linear systems. Journal of the Institute of Mathematics and its Applications, 4(??):282–294, 1968. CO-DEN JMTAA8. ISSN 0020-2932.

Levin:1976:GOP

[483] D. Levin. General order Padé-type rational approximants defined from double power series. Journal of the Institute of Mathematics and its Applications, 18 (1):1–8, 1976. CODEN JMTAA8. ISSN 0020-2932.

Levin:1980:CCA

[484] D. Levin. Corrected collocation approximations for the harmonic Dirichlet problem. Journal of the Institute of Mathematics and its Applications, 26(1):65– 75, 1980. CODEN JMTAA8. ISSN 0020-2932.

Levin:1978:BKM

[485] D. Levin, N. Papamichael, and A. Sideridis. The Bergman kernel method for the numerical conformal mapping of simply connected domains. Journal of the Institute of Mathematics and its Applications, 22(2):171–187, 1978. CODEN JMTAA8. ISSN 0020-2932.

Lewis:1975:NSF

[486] B. A. Lewis. On the numerical solution of Fredholm integral equations of the first kind. Journal of the Institute of Mathematics and its Applications, 16 (2):207–220, 1975. CODEN JMTAA8. ISSN 0020-2932.

Lighthill:1965:GV

[487] M. J. Lighthill. Group velocity. Journal of the Institute of Mathematics and its Applications, 1(??):1–28, 1965. CODEN JMTAA8. ISSN 0020-2932.

Lighthill:1968:MC

[488] M. J. Lighthill. Mathematics in control. Journal of the Institute of Mathematics and its Applications, 4(??):1–19, 1968. CODEN JMTAA8. ISSN 0020-2932.

Linz:1974:SAM

[489] P. Linz. A simple approximation method for solving Volterra integro-differential equations of the first kind. Journal of the Institute of Mathematics and its Applications, 14(??):211–215, 1974. CODEN JMTAA8. ISSN 0020-2932.

Linz:1977:SCD

[490] P. Linz. Stability and convergence of difference schemes for some functional differential equations. Journal of the Institute of Mathematics and its Applications, 19(1):71–76, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Littlewood:1976:NEF

[491] R. K. Littlewood and V. Zakian. Numerical evaluation of Fourier integrals. Journal of the Institute of Mathematics and its Applications, 18(3):331–339, 1976. CODEN JMTAA8. ISSN 0020-2932.

Long:1977:UDD

[492] M. J. Long and K. W. Morton. The use of divided differences in finite element calculations. Journal of the Institute of Mathematics and its Applications, 19(3): 307–323, 1977. CODEN JMTAA8. ISSN 0020-2932.

Longman:1968:NIL

[493] I. M. Longman. On the numerical inversion of the Laplace transform of a discontinuous original. Journal of the Institute of Mathematics and its Applications, 4(??):320–328, 1968. CODEN JM-TAA8. ISSN 0020-2932.

Longman:1972:CPJ

[494] I. M. Longman. Comments on a paper by J. A. Murphy: “Certain rational function approximations to (1 + x^{2})^{−1/2}”. With a reply by J. A. Murphy. Journal of the Institute of Mathematics and its Applications, 9(??):411– 412, 1972. CODEN JMTAA8. ISSN 0020-2932.

Longuet-Higgins:1978:SNR

[495] M. S. Longuet-Higgins. Some new relations between Stokes’s coefficients in the theory of gravity waves. Journal of the Institute of Mathematics and its Applications, 22(3):261–273, 1978. CODEN JMTAA8. ISSN 0020-2932.

Lord:1979:MNV

[496] M. E. Lord. The method of nonlinear variation of constants for difference equations. Journal of the Institute of Mathematics and its Applications, 23(3): 285–290, 1979. CODEN JMTAA8. ISSN 0020-2932.

Love:1974:LWA

[497] J. D. Love. Long wavelength acoustic scattering by a torus of arbitrary aspect ratio. Journal of the Institute of Mathematics and its Applications, 13(??):321– 344, 1974. CODEN JMTAA8. ISSN 0020-2932.

Love:1979:NCT

[498] J. D. Love. A note on the capacitance of two closely separated spheres. Journal of the Institute of Mathematics and its Applications, 24(3):255–257, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Lyness:1975:MDS

[499] J. N. Lyness and D. Jespersen. Moderate degree symmetric quadrature rules for the triangle. Journal of the Institute of Mathematics and its Applications, 15 (??):19–32, 1975. CODEN JMTAA8. ISSN 0020-2932.

MacDonald:1970:SIB

[500] D. A. MacDonald. Solution of the incompressible boundary layer equations via the Galerkin-Kantorovich technique. Journal of the Institute of Mathematics and its Applications, 6(??):115–130, 1970. CODEN JMTAA8. ISSN 0020-2932.

Madsen:1975:AMS

[501] K. Madsen. An algorithm for minimax solution of overdetermined systems of non-linear equations. Journal of the Institute of Mathematics and its Applications, 16(3):321–328, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Majumdar:1976:VBE

[502] S. R. Majumdar and M. E. O’Neill. Vertical boundary effects on the electrohydrostatic instability of a fluid interface. Journal of the Institute of Mathematics and its Applications, 17(3):343–354, 1976. CODEN JMTAA8. ISSN 0020-2932.

Marimont:1979:NNS

[503] R. B. Marimont and M. B. Shapiro.

Nearest neighbour searches and the curse of dimensionality. Journal of the Institute of Mathematics and its Applications, 24(??):59–70, 1979. CODEN JMTAA8. ISSN 0020-2932.

Maroulas:1978:SNR

[504] J. Maroulas and S. Barnett. Some new results on the qualitative theory of generalized polynomials. Journal of the Institute of Mathematics and its Applications, 22(??):53–70, 1978. CODEN JM-TAA8. ISSN 0020-2932.

Maroulas:1979:FRQ

[505] J. Maroulas and S. Barnett. Further results on the qualitative theory of generalized polynomials. Journal of the Institute of Mathematics and its Applications, 23(1):33–42, 1979. CODEN JM-TAA8. ISSN 0020-2932.

Marsh:1976:IMS

[506] T. Marsh and M. Wadsworth. An iterative method for the solution of Fredholm integral equations of the second kind. Journal of the Institute of Mathematics and its Applications, 18(1):57–65, 1976. CODEN JMTAA8. ISSN 0020-2932.

Marshall:1970:DPD

[507] E. A. Marshall. The dynamic plane deformation of a rigid perfectly-plastic slab. Journal of the Institute of Mathematics and its Applications, 6(??):201– 211, 1970. CODEN JMTAA8. ISSN 0020-2932.

Marshall:1973:EBT

[508] J. A. Marshall and A. R. Mitchell. An exact boundary technique for improved accuracy in the finite element method. Journal of the Institute of Mathematics and its Applications, 12(??):355–362, 1973. CODEN JMTAA8. ISSN 0020-2932.

Mason:1975:DOR

[509] D. P. Mason. Derivation of orthogonality relations in plasma theory using singular integral equations. Journal of the Institute of Mathematics and its Applications, 15(??):257–265, 1975. CODEN JMTAA8. ISSN 0020-2932.

Mauldon:1973:EAT

[510] J. G. Mauldon. The elimination of all but two from a tournament. Journal of the Institute of Mathematics and its Applications, 11(??):151–156, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Mavrommatis:1978:ESI

[511] E. Mavrommatis. Evaluation of some integrals encountered in calculations in many-fermion systems. Journal of the Institute of Mathematics and its Applications, 22(3):255–260, 1978. CODEN JMTAA8. ISSN 0020-2932.

Mayne:1967:RAC

[512] D. Q. Mayne. Recent approaches to the control of stochastic processes. Journal of the Institute of Mathematics and its Applications, 3(??):46–59, 1967. CO-DEN JMTAA8. ISSN 0020-2932.

McAllister:1971:NPS

[513] G. T. McAllister. On the Neumann problem for some linear and non-linear uniformly elliptic difference equations. Journal of the Institute of Mathematics and its Applications, 7(??):151–167, 1971. CODEN JMTAA8. ISSN 0020-2932.

McAllister:1972:MBV

[514] G. T. McAllister. Mixed boundary value problems for some linear and non-linear elliptic difference equations. Journal of the Institute of Mathematics and its Applications, 9(??):205–217, 1972. CODEN JMTAA8. ISSN 0020-2932.

McCabe:1975:FEP

[515] J. H. McCabe. A formal extension of the Padé table to include two point Padé quotients. Journal of the Institute of Mathematics and its Applications, 15 (??):363–372, 1975. CODEN JMTAA8. ISSN 0020-2932.

McCabe:1976:CFW

[516] J. H. McCabe and J. A. Murphy. Continued fractions which correspond to power series expansions at two points. Journal of the Institute of Mathematics and its Applications, 17(2):233–247, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

McClure:1978:EET

[517] J. P. McClure and R. Wong. Explicit error terms for asymptotic expansions of Stieltjes transforms. Journal of the Institute of Mathematics and its Applications, 22(2):129–145, 1978. CODEN JM-TAA8. ISSN 0020-2932.

McClure:1979:ERA

[518] J. P. McClure and R. Wong. Exact remainders for asymptotic expansions of fractional integrals. Journal of the Institute of Mathematics and its Applications, 24(2):139–147, 1979. CODEN JM-TAA8. ISSN 0020-2932.

McCormack:1975:VEA

[519] P. D. McCormack and D. Fitzgerald. The visco-elastic arterial flow model and the ultrasound Doppler probe. Journal of the Institute of Mathematics and its Applications, 16(3):361–370, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

McDiarmid:1972:STP

[520] C. J. H. McDiarmid. The solution of a timetabling problem. Journal of the Institute of Mathematics and its Applications, 9(??):23–34, 1972. CODEN JM-TAA8. ISSN 0020-2932.

McGuire:1972:BTM

[521] G. R. McGuire and J. Ll. Morris. Boundary techniques for the multistep formulation of the optimized Lax–Wendroff method for non-linear hyperbolic systems in two space dimensions. Journal of the Institute of Mathematics and its Applications, 10(??):150–165, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

McGuire:1976:ROA

[522] G. R. McGuire and J. Ll. Morris. Restoring orders of accuracy for multilevel schemes for non-linear hyperbolic systems in many space variables. Journal of the Institute of Mathematics and its Applications, 17(1):53–67, 1976. CODEN JMTAA8. ISSN 0020-2932.

McKee:1971:ADM

[523] S. McKee. Alternating direction methods for a system of parabolic equations in two space dimensions with a mixed derivative. Journal of the Institute of Mathematics and its Applications, 8(??): 376–385, 1971. CODEN JMTAA8. ISSN 0020-2932.

McKee:1972:ADM

[524] S. McKee. Alternating direction methods for hyperbolic equations with mixed derivatives. Journal of the Institute of Mathematics and its Applications, 9(??): 350–356, 1972. CODEN JMTAA8. ISSN 0020-2932.

McKee:1972:GDF

[525] S. McKee. A generalization of the Du Fort-Frankel scheme. Journal of the Institute of Mathematics and its Applications, 10(??):42–48, 1972. CODEN JM-TAA8. ISSN 0020-2932.

McKee:1973:HAD

[526] S. McKee. High accuracy A.D.I. methods for hyperbolic equations with variable coefficients. Journal of the Institute of Mathematics and its Applications, 11 (??):105–109, 1973. CODEN JMTAA8. ISSN 0020-2932.

McKee:1979:AVS

[527] S. McKee. The analysis of a variable step, variable coefficient linear multistep method for solving a singular integrodifferential equation arising from the diffusion of discrete particles in a turbulent fluid. Journal of the Institute of Mathematics and its Applications, 23(3):373– 388, 1979. CODEN JMTAA8. ISSN 0020-2932.

McLeod:1976:NRH

[528] R. McLeod. Node requirements for highorder approximation over curved finite elements. Journal of the Institute of Mathematics and its Applications, 17(2): 249–254, 1976. CODEN JMTAA8. ISSN 0020-2932.

McLeod:1978:HOT

[529] R. McLeod. High order transformation methods for curved finite elements. Journal of the Institute of Mathematics and its Applications, 21(??):419–428, 1978. CODEN JMTAA8. ISSN 0020-2932.

McLeod:1972:CBF

[530] R. McLeod and A. R. Mitchell. The construction of basis functions for curved elements in the finite element method. Journal of the Institute of Mathematics and its Applications, 10(??):382–393, 1972. CODEN JMTAA8. ISSN 0020-2932.

McLeod:1975:UPA

[531] R. J. Y. McLeod and A. R. Mitchell. The use of parabolic arcs in matching curved boundaries in the finite element method. Journal of the Institute of Mathematics and its Applications, 16(2):239–246, 1975. CODEN JMTAA8. ISSN 0020-2932.

Mead:1972:PEC

[532] K. O. Mead and L. M. Delves. On the prediction of the expansion coefficients in a variational calculation. Journal of the Institute of Mathematics and its Applications, 10(??):166–175, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Mead:1973:CRG

[533] K. O. Mead and L. M. Delves. On the convergence rate of generalized Fourier expansions. Journal of the Institute of Mathematics and its Applications, 12 (??):247–259, 1973. CODEN JMTAA8. ISSN 0020-2932.

Mees:1972:DFM

[534] A. I. Mees. The describing function matrix. Journal of the Institute of Mathematics and its Applications, 10(??):49– 67, 1972. CODEN JMTAA8. ISSN 0020-2932.

Mees:1973:LCS

[535] A. I. Mees. Limit cycle stability. Journal of the Institute of Mathematics and its Applications, 11(??):281–295, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Meyer:1977:AML

[536] G. H. Meyer. An application of the method of lines to multidimensional free boundary problems. Journal of the Institute of Mathematics and its Applications, 20(3):317–329, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Meyer:1967:ADL

[537] R. E. Meyer. On the approximation of double limits by single limits and the Kaplun extension theorem. Journal of the Institute of Mathematics and its Applications, 3(??):245–249, 1967. CODEN JMTAA8. ISSN 0020-2932.

Meyer:1972:MDL

[538] Robert R. Meyer and Philip M. Roth. Modified damped least squares: An algorithm for non-linear estimation. Journal of the Institute of Mathematics and its Applications, 9(??):218–233, 1972. CODEN JMTAA8. ISSN 0020-2932.

Mezei:1973:EFA

[539] M. Mezei. The error of a function approximation based on random selected points. Journal of the Institute of Mathematics and its Applications, 12(??):97– 102, 1973. CODEN JMTAA8. ISSN 0020-2932.

Mezei:1973:PRB

[540] M. Mezei and T. Szondy. On the possibility of replacing the backward transformation in the modified simplex method by a forward transformation. Journal of the Institute of Mathematics and its Applications, 11(??):339–342, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Micchelli:1979:OES

[541] C. A. Micchelli. Optimal estimation of smooth functions from inaccurate data. Journal of the Institute of Mathematics and its Applications, 23(4):473–495, 1979. CODEN JMTAA8. ISSN 0020-2932.

Miller:1978:FEM

[542] J. V. Miller, K. W. Morton, and M. J.

Baines. A finite element moving boundary computation with an adaptive mesh. Journal of the Institute of Mathematics and its Applications, 22(4):467–477, 1978. CODEN JMTAA8. ISSN 0020-2932.

Miller:1971:LZC

[543] John J. H. Miller. On the location of zeros of certain classes of polynomials with applications to numerical analysis. Journal of the Institute of Mathematics and its Applications, 8(??):397–406, 1971. CODEN JMTAA8. ISSN 0020-2932.

Minieka:1973:NAB

[544] E. Minieka and D. R. Shier. A note on an algebra for the k best routes in a network. Journal of the Institute of Mathematics and its Applications, 11(??):145– 149, 1973. CODEN JMTAA8. ISSN 0020-2932.

Mitchell:1972:VPF

[545] A. R. Mitchell. Variational principles and the finite element method. Journal of the Institute of Mathematics and its Applications, 9(??):378–389, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Mitchell:1971:FSF

[546] A. R. Mitchell, G. Phillips, and E. Wachspress. Forbidden shapes in the finite element method. Journal of the Institute of Mathematics and its Applications, 8 (??):260–269, 1971. CODEN JMTAA8. ISSN 0020-2932.

Mitra:1968:IRP

[547] Debasis Mitra. The irreducible representations of 2p-pole linear Lagrangian systems. Journal of the Institute of Mathematics and its Applications, 4(??):329– 338, 1968. CODEN JMTAA8. ISSN 0020-2932.

Mocarsky:1971:CSS

[548] William L. Mocarsky. Convergence of step-by-step methods for non-linear integro-differential equations. Journal of the Institute of Mathematics and its Applications, 8(??):235–239, 1971. CODEN JMTAA8. ISSN 0020-2932.

Moore:1976:CRA

[549] J. B. Moore. A consistently rapid algorithm for solving polynomial equations. Journal of the Institute of Mathematics and its Applications, 17(1):99–110, 1976. CODEN JMTAA8. ISSN 0020-2932.

Moore:1978:FEM

[550] P. Moore. Finite element multistep multiderivative schemes for parabolic equations. Journal of the Institute of Mathematics and its Applications, 21(??):331– 344, 1978. CODEN JMTAA8. ISSN 0020-2932.

Morley:1965:VMS

[551] L. S. D. Morley. A variational method of solution for problems in plane elasticity. Journal of the Institute of Mathematics and its Applications, 1(??):76–100, 1965. CODEN JMTAA8. ISSN 0020-2932.

Morris:1976:PAP

[552] H. C. Morris. A perturbative approach to periodic solutions of delay-differential equations. Journal of the Institute of Mathematics and its Applications, 18(1): 15–24, 1976. CODEN JMTAA8. ISSN 0020-2932.

Morris:1973:MLO

[553] J. Ll. Morris and A. R. Gourlay. Modified locally one dimensional methods for parabolic partial differential equations in two space variables. Journal of the Institute of Mathematics and its Applications, 12(??):349–353, 1973. CODEN JMTAA8. ISSN 0020-2932.

Moss:1980:ASD

[554] R. L. Moss, T. G. Rogers, and A. J. M.

Spencer. Axially symmetric deformations of fibre-reinforced cylinders. Journal of the Institute of Mathematics and its Applications, 26(1):21–38, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Mostaghel:1977:SRH

[555] N. Mostaghel. Stability regions of Hill’s equation. Journal of the Institute of Mathematics and its Applications, 19(3): 253–259, 1977. CODEN JMTAA8. ISSN 0020-2932.

Munt:1975:ARC

[556] R. M. Munt. Acoustic radiation from a circular cylinder in a subsonic stream. Journal of the Institute of Mathematics and its Applications, 16(1):1–10, 1975. CODEN JMTAA8. ISSN 0020-2932.

Murphy:1971:CRF

[557] J. A. Murphy. Certain rational function approximations to (1 + x^{2})^{−1/2}. Journal of the Institute of Mathematics and its Applications, 7(??):138–150, 1971. CO-DEN JMTAA8. ISSN 0020-2932.

Murphy:1975:SPC

[558] J. A. Murphy and M. R. O’Donohoe.

Some properties of continued fractions with applications in Markov processes. Journal of the Institute of Mathematics and its Applications, 16(1):57–71, 1975. CODEN JMTAA8. ISSN 0020-2932.

Murray:1972:NGL

[559] J. C. Murray. A note on a general linear initial-boundary value problem. Journal of the Institute of Mathematics and its Applications, 10(??):305–311, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Murray:1973:NSV

[560] J. C. Murray. A note on some variational principles for a class of linear initialboundary value problems. Journal of the Institute of Mathematics and its Applications, 12(??):119–123, 1973. CODEN JMTAA8. ISSN 0020-2932.

Murray:1975:AVP

[561] J. C. Murray. Addendum to: “A variational principle associated with surface waves” (J. Inst. Math. Appl. 15 (1975), 127–134). Journal of the Institute of Mathematics and its Applications, 16(3): 387–388, 1975. CODEN JMTAA8. ISSN 0020-2932. See [562].

Murray:1975:VPA

[562] J. C. Murray. A variational principle associated with surface waves. Journal of the Institute of Mathematics and its Applications, 15(??):127–134, 1975. CO-DEN JMTAA8. ISSN 0020-2932. See addendum [561].

Murray:1978:CCS

[563] J. C. Murray. Curvilinear coordinate systems suitable for the numerical solution of boundary-value problems. Journal of the Institute of Mathematics and its Applications, 22(4):411–421, 1978. CODEN JMTAA8. ISSN 0020-2932.

Murray:1980:CCS

[564] J. C. Murray. Curvilinear coordinate systems associated with a class of boundary value problems. Journal of the Institute of Mathematics and its Applications, 25(4):397–411, 1980. CODEN JMTAA8. ISSN 0020-2932.

Musa:1979:WAD

[565] F. A. Musa and L. M. Delves. On weakly asymptotically diagonal systems. Journal of the Institute of Mathematics and its Applications, 23(4):389–403, 1979. CODEN JMTAA8. ISSN 0020-2932.

Namias:1980:FOF

[566] Victor Namias. The fractional order Fourier transform and its application to quantum mechanics. Journal of the Institute of Mathematics and its Applications, 25(3):241–265, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Namias:1980:FHT

[567] Victor Namias. Fractionalization of Hankel transforms. Journal of the Institute of Mathematics and its Applications, 26(2):187–197, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Nash:1977:MNL

[568] J. C. Nash. Minimizing a non-linear sum of squares function on a small computer. Journal of the Institute of Mathematics and its Applications, 19(2):231–237, 1977. CODEN JMTAA8. ISSN 0020-2932.

Nelson:1979:PIS

[569] P. R. Nelson. A prediction interval search scheme for the move-to-the-front replacement algorithm. Journal of the Institute of Mathematics and its Applications, 24(2):231–236, 1979. CODEN JMTAA8. ISSN 0020-2932.

Newland:1965:ESR

[570] D. E. Newland. Energy sharing in the random vibration of non-linearly coupled modes. Journal of the Institute of Mathematics and its Applications, 1(??): 199–207, 1965. CODEN JMTAA8. ISSN 0020-2932.

Newlands:1977:STE

[571] A. G. Newlands. On the sensitivity of transient elastic stresses in a solid to the form of surface loading. Journal of the Institute of Mathematics and its Applications, 19(??):195–203, 1977. CODEN JMTAA8. ISSN 0020-2932.

Nicholson:1971:EDP

[572] T. A. J. Nicholson. Extension of dynamic programming to multi-plant renewal. Journal of the Institute of Mathematics and its Applications, 8(??):225– 234, 1971. CODEN JMTAA8. ISSN 0020-2932.

Nicol:1978:BFR

[573] D. A. C. Nicol. Bending and flexure of a reinforced tube. Journal of the Institute of Mathematics and its Applications, 21 (??):117–126, 1978. CODEN JMTAA8. ISSN 0020-2932.

Nieminen:1974:TBD

[574] J. Nieminen. Two bounds for the domination number of a graph. Journal of the Institute of Mathematics and its Applications, 14(??):183–187, 1974. CODEN JMTAA8. ISSN 0020-2932.

Niki:1978:NPT

[575] Hiroshi Niki, Hideo Sawami, and Tetsuya Ondoh. A note on a practical technique for solving elliptic equations. Journal of the Institute of Mathematics and its Applications, 22(3):353–359, 1978. CODEN JMTAA8. ISSN 0020-2932.

Nilsson:1980:PSCb

[576] Börje Nilsson and Olle Brander. The propagation of sound in cylindrical ducts with mean flow and bulk-reacting lining. I. Modes in an infinite duct. Journal of the Institute of Mathematics and its Applications, 26(3):269–298, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Nilsson:1980:PSCa

[577] Börje Nilsson and Olle Brander. The propagation of sound in cylindrical ducts with mean flow and bulk-reacting lining. II. Bifurcated ducts. Journal of the Institute of Mathematics and its Applications, 26(4):381–410, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Noble:1980:EET

[578] B. Noble and S. Beighton. Error estimates for three methods of evaluating Cauchy principal value integrals. Journal of the Institute of Mathematics and its Applications, 26(4):431–446, 1980. CODEN JMTAA8. ISSN 0020-2932.

Noble:1972:DEP

[579] B. Noble and M. J. Sewell. On dual extremum principles in applied mathematics. Journal of the Institute of Mathematics and its Applications, 9(??):123– 193, 1972. CODEN JMTAA8. ISSN 0020-2932.

Noble:1969:VMI

[580] Ben Noble and M. Hussain. A variational method for inclusion and indentation problems. Journal of the Institute of Mathematics and its Applications, 5 (??):194–205, 1969. CODEN JMTAA8. ISSN 0020-2932.

Nonweiler:1977:IAE

[581] T. Nonweiler. Integrals and areas expressed as bilinear forms of discrete defining data. Journal of the Institute of Mathematics and its Applications, 19 (2):159–168, 1977. CODEN JMTAA8. ISSN 0020-2932.

Nuttall:1978:GIC

[582] J. Nuttall and C. J. Wherry. Gaussian integration for complex weight functions. Journal of the Institute of Mathematics and its Applications, 21(??):165– 170, 1978. CODEN JMTAA8. ISSN 0020-2932.

Nye:1977:UMF

[583] V. A. Nye. Unsteady magnetohydrodynamic flows. Journal of the Institute of Mathematics and its Applications, 20 (??):229–243, 1977. CODEN JMTAA8. ISSN 0020-2932.

OCarroll:1973:IRC

[584] M. J. O’Carroll. Inconsistencies and S.O.R. convergence for the discrete Neumann problem. Journal of the Institute of Mathematics and its Applications, 11 (??):343–350, 1973. CODEN JMTAA8. ISSN 0020-2932.

OCarroll:1974:UCVa

[585] M. J. O’Carroll. Uniform convergence of variational methods for elliptic problems in N dimensions. Journal of the Institute of Mathematics and its Applications, 13 (??):1–5, 1974. CODEN JMTAA8. ISSN 0020-2932.

OCarroll:1974:UCVb

[586] M. J. O’Carroll. Uniform convergence of variational methods for plane subsonic flow. Journal of the Institute of Mathematics and its Applications, 13(??):7–12, 1974. CODEN JMTAA8. ISSN 0020-2932.

Ohiwa:1978:NAN

[587] H. Ohiwa. A new algorithm for nonlinear simultaneous equations. Journal of the Institute of Mathematics and its Applications, 21(??):189–196, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Oliver:1968:IIS

[588] J. Oliver. Inherent instability in systems of first order linear differential equations. Journal of the Institute of Mathematics and its Applications, 4(??):399– 409, 1968. CODEN JMTAA8. ISSN 0020-2932.

Oliver:1971:PSC

[589] J. Oliver. A practical strategy for the Clenshaw–Curtis quadrature method. Journal of the Institute of Mathematics and its Applications, 8(??):53–56, 1971. CODEN JMTAA8. ISSN 0020-2932.

Oliver:1975:SME

[590] J. Oliver. Stable methods for evaluating the points cos(iπ/n). Journal of the Institute of Mathematics and its Applications, 16(2):247–257, 1975. CODEN JMTAA8. ISSN 0020-2932.

Oliver:1977:EAM

[591] J. Oliver. An error analysis of the modified Clenshaw method for evaluating Chebyshev and Fourier series. Journal of the Institute of Mathematics and its Applications, 20(3):379–391, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

OLoghlen:1973:EPA

[592] J. W. OLoghlen. An eigenvalue problem arising in the theory of shock waves. Journal of the Institute of Mathematics and its Applications, ??(??):??, 1973. CODEN JMTAA8. ISSN 0020-2932.

OMalley:1970:NLS

[593] Robert E. O’Malley, Jr. A non-linear singular perturbation problem arising in the study of chemical flow reactors. Journal of the Institute of Mathematics and its Applications, 6(??):12–20, 1970. CODEN JMTAA8. ISSN 0020-2932.

Omodei:1977:SRR

[594] B. J. Omodei. Stability of the Rayleigh– Ritz–Galerkin procedure for linear elliptic boundary value problems. Journal of the Institute of Mathematics and its Applications, 20(2):245–255, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Orley:1973:NHA

[595] D. G. Orley and S. McKee. A note on hopscotch as applied to the bending beam equation. Journal of the Institute of Mathematics and its Applications, 11 (??):335–338, 1973. CODEN JMTAA8. ISSN 0020-2932.

Osborne:1974:AEF

[596] M. R. Osborne. Asymptotic error formulae for numerical quadrature. Journal of the Institute of Mathematics and its Applications, 13(??):219–227, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Pack:1974:CFV

[597] D. C. Pack and J. M. Sharp. Collisionless flow in vacuum traps. Journal of the Institute of Mathematics and its Applications, 13(??):23–38, 1974. CODEN JMTAA8. ISSN 0020-2932.

Packham:1972:NTW

[598] B. A. Packham and W. E. Williams. A note on the transmission of water waves through small apertures. Journal of the Institute of Mathematics and its Applications, 10(??):176–184, 1972. CODEN JMTAA8. ISSN 0020-2932.

Paige:1972:CVL

[599] C. C. Paige. Computational variants of the Lanczos method for the eigenproblem. Journal of the Institute of Mathematics and its Applications, 10(3):373– 381, December 1972. CODEN JMTAA8. ISSN 0020-2932.

Paige:1976:EAL

[600] C. C. Paige. Error analysis of the Lanczos algorithm for tridiagonalizing a symmetric matrix. Journal of the Institute of Mathematics and its Applications, 18 (3):341–349, 1976. CODEN JMTAA8. ISSN 0020-2932.

Panek:1980:TDC

[601] Carl Panek and J. J. Kalker. Threedimensional contact of a rigid roller traversing a viscoelastic half space. Journal of the Institute of Mathematics and its Applications, 26(3):299–313, 1980. CODEN JMTAA8. ISSN 0020-2932.

Panwalkar:1979:OFS

[602] S. S. Panwalkar and A. W. Khan. An ordered flow shop sequencing problem with TWK type due dates and total tardiness criterion. Journal of the Institute of Mathematics and its Applications, 23 (3):367–372, 1979. CODEN JMTAA8. ISSN 0020-2932.

Papaconstantinou:1976:ISG

[603] C. Papaconstantinou. Iterative solution of the generalized eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 17(2):255–260, 1976. CODEN JMTAA8. ISSN 0020-2932.

Papamichael:1979:UCT

[604] N. Papamichael and A. Sideridis. The use of conformal transformations for the numerical solution of elliptic boundary value problems with boundary singularities. Journal of the Institute of Mathematics and its Applications, 23(??):73– 88, 1979. CODEN JMTAA8. ISSN 0020-2932.

Papamichael:1975:DCO

[605] N. Papamichael and G. D. Smith. The determination of consistent orderings for the SOR iterative method. Journal of the Institute of Mathematics and its Applications, 15(??):239–248, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Papamichael:1973:CST

[606] N. Papamichael and J. R. Whiteman. A cubic-spline technique for the one dimensional heat conduction equation. Journal of the Institute of Mathematics and its Applications, 11(??):111–113, 1973. CODEN JMTAA8. ISSN 0020-2932.

Parker:1971:ATO

[607] D. F. Parker. An asymptotic theory for oscillatory non-linear signals. Journal of the Institute of Mathematics and its Applications, 7(??):92–110, 1971. CODEN JMTAA8. ISSN 0020-2932.

Pathak:1978:SLP

[608] R. S. Pathak. Summability of the Laguerre polynomial expansion of generalized functions. Journal of the Institute of Mathematics and its Applications, 21 (??):171–180, 1978. CODEN JMTAA8. ISSN 0020-2932.

Peraiah:1973:NSR

[609] A. Peraiah and I. P. Grant. Numerical solution of the radiative transfer equation in spherical shells. Journal of the Institute of Mathematics and its Applications, 12(??):75–90, 1973. CODEN JMTAA8. ISSN 0020-2932.

Pereyra:1973:EST

[610] V. Pereyra and G. Scherer. Eigenvalues of symmetric tridiagonal matrices: a fast, accurate and reliable algorithm. Journal of the Institute of Mathematics and its Applications, 12(??):209–222, 1973. CODEN JMTAA8. ISSN 0020-2932.

Peters:1971:PPA

[611] G. Peters and J. H. Wilkinson. Practical problems arising in the solution of polynomial equations. Journal of the Institute of Mathematics and its Applications, 8(??):16–35, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Pickering:1977:SCS

[612] W. M. Pickering. Some comments on the solution of Poisson’s equation using Bickley’s formula and fast Fourier transforms. Journal of the Institute of Mathematics and its Applications, 19(3):337– 338, 1977. CODEN JMTAA8. ISSN 0020-2932.

Piessens:1972:NNM

[613] R. Piessens. A new numerical method for the inversion of the Laplace transform. Journal of the Institute of Mathematics and its Applications, 10(??):185– 192, 1972. CODEN JMTAA8. ISSN 0020-2932.

Pitkaranta:1979:DPS

[614] J. Pitkäranta. On the differential properties of solutions to Fredholm equations with weakly singular kernels. Journal of the Institute of Mathematics and its Applications, 24(2):109–119, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Pitts:1976:SDH

[615] E. Pitts. The stability of a drop hanging from a tube. Journal of the Institute of Mathematics and its Applications, 17(3): 387–397, 1976. CODEN JMTAA8. ISSN 0020-2932.

Polak:1974:IAS

[616] S. J. Polak. An increased accuracy scheme for diffusion equations in cylindrical coordinates. Journal of the Institute of Mathematics and its Applications, 14(??):197–201, 1974. CODEN JMTAA8. ISSN 0020-2932.

Pollock:1980:ITQ

[617] J. Pollock and A. J. Pritchard. The infinite time quadratic cost control problem for distributed systems with unbounded control action. Journal of the Institute of Mathematics and its Applications, 25 (3):287–309, 1980. CODEN JMTAA8. ISSN 0020-2932.

Pollock:1980:TPD

[618] J. Pollock and A. J. Pritchard. The tracking problem for distributed systems with unbounded control action. Journal of the Institute of Mathematics and its Applications, 25(4):387–396, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Pomentale:1974:HIM

[619] T. Pomentale. Homotopy iterative methods for polynomial equations. Journal of the Institute of Mathematics and its Applications, 13(??):201–213, 1974. CODEN JMTAA8. ISSN 0020-2932.

Pop:1975:UBL

[620] Ioan Pop and V. M. Soundalgekar. On unsteady boundary layers in a rotating flow. Journal of the Institute of Mathematics and its Applications, 15(??):343– 349, 1975. CODEN JMTAA8. ISSN 0020-2932.

Powell:1974:DSS

[621] J. H. Powell. Deficiencies in safety schemes which rely on stochastically failing protective equipment. Journal of the Institute of Mathematics and its Applications, 14(??):41–56, 1974. CODEN JMTAA8. ISSN 0020-2932.

Powell:1975:AME

[622] J. H. Powell. Assessment of the maintenance of the effectiveness of safety schemes by inspection and agereplacement of protective devices. Journal of the Institute of Mathematics and its Applications, 16(1):81–92, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Powell:1971:CVM

[623] M. J. D. Powell. On the convergence of the variable metric algorithm. Journal of the Institute of Mathematics and its Applications, 7(1):21–36, February 1971. CODEN JMTAA8. ISSN 0020-2932.

Powell:1972:QTPa

[624] M. J. D. Powell. Quadratic termination properties of minimization algorithms. I. Statement and discussion of results. Journal of the Institute of Mathematics and its Applications, 10(3):333– 342, December 1972. CODEN JM-TAA8. ISSN 0020-2932. URL https:/

/academic.oup.com/imamat/article-
abstract/10/3/333/824259.

Powell:1972:QTPb

[625] M. J. D. Powell. Quadratic termination properties of minimization algorithms. II. Proofs of theorems. Journal of the Institute of Mathematics and its Applications, 10(3):343–357, December 1972. CODEN JMTAA8. ISSN 0020-

2932. URL https://academic.oup.
com/imamat/article-abstract/10/3/
343/824267.

Powell:1966:WUS

[626] M. J. D. Powell and J. Swann. Weighted uniform sampling — a Monte Carlo technique for reducing variance. Journal of the Institute of Mathematics and its Applications, 2(3):228–236, September 1966. CODEN JMTAA8. ISSN 0020-2932.

Prasad:1977:TNL

[627] Phoolan Prasad and Renuka Ravindran. A theory of non-linear waves in multidimensions: With special reference to surface water waves. Journal of the Institute of Mathematics and its Applications, 20(1):9–20, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Price:1979:WSP

[628] W. L. Price. A weighted simplex procedure for the solution of simultaneous non-linear equations. Journal of the Institute of Mathematics and its Applications, 24(??):1–8, 1979. CODEN JM-TAA8. ISSN 0020-2932.

Prince:1978:RKP

[629] P. J. Prince and K. Wright. Runge– Kutta processes with exact principal error equations. Journal of the Institute of Mathematics and its Applications, 21 (??):363–373, 1978. CODEN JMTAA8. ISSN 0020-2932.

Pritchard:1971:SSS

[630] A. J. Pritchard. Stability and stabilization of second-order systems. Journal of the Institute of Mathematics and its Applications, 7(??):348–360, 1971. CODEN JMTAA8. ISSN 0020-2932.

Pritchard:1974:LBC

[631] A. J. Pritchard and K. T. Parker. A lower bound for the cost functional for control problems in Hilbert space. Journal of the Institute of Mathematics and its Applications, 13(??):97–106, 1974. CODEN JMTAA8. ISSN 0020-2932.

Purushothaman:1977:NUB

[632] R. Purushothaman. A note on unsteady boundary layers in a rotating flow. Journal of the Institute of Mathematics and its Applications, 20(??):257–260, 1977. CODEN JMTAA8. ISSN 0020-2932.

Quinney:1979:CTW

[633] D. A. Quinney. On computing travelling wave solutions in a model for the Belousov–Zhabotinskii reaction. Journal of the Institute of Mathematics and its Applications, 23(2):193–201, March 1979. CODEN JMTAA8. ISSN 0020-

Racicot:1971:FFP

[634] R. L. Racicot. Fitting a filtered Poisson process. Journal of the Institute of Mathematics and its Applications, 7(??): 260–272, 1971. CODEN JMTAA8. ISSN 0020-2932.

Raggett:1974:FIF

[635] G. F. Raggett and P. D. Wilson. A fully implicit finite difference approximation to the one-dimensional wave equation using a cubic spline technique. Journal of the Institute of Mathematics and its Applications, 14(??):75–77, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Ralston:1973:SAD

[636] Anthony Ralston. Some aspects of degeneracy in rational approximations. Journal of the Institute of Mathematics and its Applications, 11(??):157–170, 1973. CODEN JMTAA8. ISSN 0020-2932.

Rao:1976:HON

[637] C. Prabhakara Rao and Satteluri R. K. Iyengar. High order numerical differentiation type formulas with an off-step point for stiff ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 17(3):281– 293, 1976. CODEN JMTAA8. ISSN 0020-2932.

Rao:1980:OMS

[638] Ganti Prasada Rao and T. Srinivasan. An optimal method of solving differential equations characterizing the dynamics of a current collection system for an electric locomotive. Journal of the Institute of Mathematics and its Applications, 25(4):329–342, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Rasmussen:1973:CRR

[639] H. Rasmussen. On the convergence of the Rayleigh–Ritz method in plane subsonic flow. Journal of the Institute of Mathematics and its Applications, 11 (??):1–8, 1973. CODEN JMTAA8. ISSN 0020-2932.

Rasmussen:1976:CFC

[640] H. Rasmussen. Compressible flow calculations by variational method using natural boundary conditions. Journal of the Institute of Mathematics and its Applications, 17(??):111–118, 1976. CODEN JMTAA8. ISSN 0020-2932.

Rasmussen:1974:AVM

[641] H. Rasmussen and N. Heys. Application of a variational method in plane compressible flow calculation. Journal of the Institute of Mathematics and its Applications, 13(??):47–61, 1974. CODEN JMTAA8. ISSN 0020-2932.

Rawlins:1977:DDW

[642] A. D. Rawlins. Diffraction by a dielectric wedge. Journal of the Institute of Mathematics and its Applications, 19(3): 261–279, 1977. CODEN JMTAA8. ISSN 0020-2932.

Reid:1971:NSG

[643] J. K. Reid. A note on the stability of Gaussian elimination. Journal of the Institute of Mathematics and its Applications, 8(??):374–375, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Reid:1972:CCF

[644] J. K. Reid. On the construction and convergence of a finite-element solution of Laplace’s equation. Journal of the Institute of Mathematics and its Applications, 9(??):1–13, 1972. CODEN JM-TAA8. ISSN 0020-2932.

Riess:1976:GAE

[645] R. D. Riess. A Gaussian approach to equally spaced quadratures. Journal of the Institute of Mathematics and its Applications, 17(2):261–265, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Riess:1974:DQF

[646] R. D. Riess and L. W. Johnson. On the determination of quadrature formulae of highest degree of precision for approximating Fourier coefficients. Journal of the Institute of Mathematics and its Applications, 13(??):345–351, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Riess:1978:CFR

[647] R. D. Riess and L. W. Johnson. Closedform representations for errors of cubic spline approximations on equally spaced knots. Journal of the Institute of Mathematics and its Applications, 21(??):345– 351, 1978. CODEN JMTAA8. ISSN 0020-2932.

Rive:1975:RKM

[648] L. Dalle Rive and F. Pasciutti. Runge– Kutta methods with global error estimates. Journal of the Institute of Mathematics and its Applications, 16(3):381– 386, 1975. CODEN JMTAA8. ISSN 0020-2932.

Robe:1975:NME

[649] T. R. Robe and S. E. Jones. A numerical method for establishing Liapunov stability of linear second order nonautonomous dynamical systems. Journal of the Institute of Mathematics and its Applications, 16(1):109–120, 1975. CO-DEN JMTAA8. ISSN 0020-2932.

Robertson:1976:NIS

[650] H. H. Robertson. Numerical integration of systems of stiff ordinary differential equations with special structure. Journal of the Institute of Mathematics and its Applications, 18(2):249–263, 1976. CODEN JMTAA8. ISSN 0020-2932.

Robertson:1977:AEE

[651] H. H. Robertson. The accuracy of error estimates for systems of linear algebraic equations. Journal of the Institute of Mathematics and its Applications, 20 (4):409–414, 1977. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Robertson:1979:CCM

[652] H. H. Robertson. Copolymer composition in n-monomer systems. Journal of the Institute of Mathematics and its Applications, 23(??):405–411, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Robertson:1975:SPA

[653] H. H. Robertson and J. Williams. Some properties of algorithms for stiff differential equations. Journal of the Institute of Mathematics and its Applications, 16(1): 23–34, 1975. CODEN JMTAA8. ISSN 0020-2932.

Robinson:1976:SCW

[654] I. S. Robinson. Some common wave phenomena applied to an M.H.D. anisotropic surface wave system. Journal of the Institute of Mathematics and its Applications, 18(??):159–175, 1976. CODEN JMTAA8. ISSN 0020-2932.

Rodrigues:1976:PCQ

[655] A. J. Rodrigues. Properties of constants for a quadrature formula to evaluate Bromwich’s integral. Journal of the Institute of Mathematics and its Applications, 18(1):49–56, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Rodrigues:1977:EAC

[656] A. J. Rodrigues. An error analysis and convergence of a quadrature formula to invert Laplace transforms. Journal of the Institute of Mathematics and its Applications, 20(1):21–32, 1977. CODEN JMTAA8. ISSN 0020-2932.

Rodrigues:1978:AQL

[657] A. J. Rodrigues. On the accuracy of quadrature Laplace transform inversion and the solution of state space equations. Journal of the Institute of Mathematics and its Applications, 22(3):283– 296, 1978. CODEN JMTAA8. ISSN 0020-2932.

Rodrigues:1980:CPK

[658] A. J. Rodrigues. Convergence of Padé kernel approximants to the delayed Dirac function. Journal of the Institute of Mathematics and its Applications, 25(1):17–27, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Rogers:1977:FBP

[659] J. C. W. Rogers. A free boundary problem as diffusion with non-linear absorption. Journal of the Institute of Mathematics and its Applications, 20(2):261– 268, 1977. CODEN JMTAA8. ISSN 0020-2932.

Roscoe:1975:NMD

[660] D. F. Roscoe. New methods for the derivation of stable difference representations for differential equations. Journal of the Institute of Mathematics and its Applications, 16(3):291–301, 1975. CODEN JMTAA8. ISSN 0020-2932.

Rosinger:1980:SCN

[661] E. E. Rosinger. Stability and convergence for nonlinear difference schemes are equivalent. Journal of the Institute of Mathematics and its Applications, 26 (2):143–149, 1980. CODEN JMTAA8. ISSN 0020-2932.

Rowe:1972:GRM

[662] A. Rowe. The generalized resultant matrix. Journal of the Institute of Mathematics and its Applications, 9(??):390– 396, 1972. CODEN JMTAA8. ISSN 0020-2932.

Rowlands:1974:SSN

[663] G. Rowlands. On the stability of solutions of the non-linear Schrödinger equation. Journal of the Institute of Mathematics and its Applications, 13(??):367– 377, 1974. CODEN JMTAA8. ISSN 0020-2932.

Rubenfeld:1967:DAD

[664] L. Rubenfeld and J. Bazer. Diffraction by an anisotropic disk, perfectly conducting along equiangular spirals. Journal of the Institute of Mathematics and its Applications, 3(??):216–244, 1967. CODEN JMTAA8. ISSN 0020-2932. See errata and addenda [665].

Rubenfeld:1969:EAD

[665] L. Rubenfeld and J. Bazer. Errata and addenda: “Diffraction by an anisotropic disk, perfectly conducting along equiangular spirals”. Journal of the Institute of Mathematics and its Applications, 5 (1):??, 1969. CODEN JMTAA8. ISSN 0020-2932. Loose insert. See [664].

Rubinstein:1979:SPC

[666] L. Rubinstein. The Stefan problem: comments on its present state. Journal of the Institute of Mathematics and its Applications, 24(3):259–277, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Sack:1972:VSE

[667] R. A. Sack. Variational solutions for eigenvalues of single and coupled Lamé equations. Journal of the Institute of Mathematics and its Applications, 10 (??):279–288, 1972. CODEN JMTAA8. ISSN 0020-2932.

Sack:1978:CSQ

[668] R. A. Sack. Comments on some quadrature formulas by F. Stenger. Journal of the Institute of Mathematics and its Applications, 21(??):359–361, 1978. CO-DEN JMTAA8. ISSN 0020-2932. See [710, 711].

Salamon:1979:LLH

[669] N. J. Salamon and G. G. Walter. Limits of Lipschitz-Hankel integrals. Journal of the Institute of Mathematics and its Applications, 24(3):237–254, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Salane:1980:UDS

[670] D. E. Salane and R. P. Tewarson. A unified derivation of symmetric quasi-Newton update formulas. Journal of the Institute of Mathematics and its Applications, 25(1):29–36, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Savage:1968:MDL

[671] M. D. Savage. A method for determining linear solutions in magnetohydrodynamics. Journal of the Institute of Mathematics and its Applications, 4(??):20–32, 1968. CODEN JMTAA8. ISSN 0020-2932.

Schachter:1980:GRT

[672] B. Schachter. Generation of random textures. Journal of the Institute of Mathematics and its Applications, 25(??):343– 347, 1980. CODEN JMTAA8. ISSN 0020-2932.

Schagen:1979:ITD

[673] I. P. Schagen. Interpolation in two dimensions—a new technique. Journal of the Institute of Mathematics and its Applications, 23(1):53–59, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Schagen:1980:SIF

[674] I. P. Schagen. Stochastic interpolating functions—applications in optimization. Journal of the Institute of Mathematics and its Applications, 26(1):93–101, 1980. CODEN JMTAA8. ISSN 0020-2932.

Schinas:1974:SCS

[675] J. Schinas. Stability and conditional stability of time-dependent difference equations in Banach spaces. Journal of the Institute of Mathematics and its Applications, 14(??):335–346, 1974. CODEN JMTAA8. ISSN 0020-2932.

Schultz:1975:TTC

[676] D. H. Schultz. Two test cases for the numerical solution of harmonic mixed boundary value problems. Journal of the Institute of Mathematics and its Applications, 15(??):1–8, 1975. CODEN JM-TAA8. ISSN 0020-2932.

Scorer:1967:LIC

[677] R. S. Scorer. Local instability in curved flow. Journal of the Institute of Mathematics and its Applications, 3(??):250– 265, 1967. CODEN JMTAA8. ISSN 0020-2932.

Scott:1980:EPP

[678] J. F. Scott. An existence proof for a problem in singular parabolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 26(3):221–234, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Searl:1971:ESP

[679] J. W. Searl. Expansions for singular perturbations. Journal of the Institute of Mathematics and its Applications, 8(??): 131–138, 1971. CODEN JMTAA8. ISSN 0020-2932.

Serbin:1973:OCH

[680] Hyman Serbin. The operational calculus of Heaviside. Journal of the Institute of Mathematics and its Applications, 11 (??):131–143, 1973. CODEN JMTAA8. ISSN 0020-2932.

Shafai:1973:ECB

[681] L. Shafai. Expansion of cylindrical Bessel functions in terms of trigonometric functions. Journal of the Institute of Mathematics and its Applications, 11 (??):227–230, 1973. CODEN JMTAA8. ISSN 0020-2932.

Shail:1979:SPP

[682] R. Shail. Some potential problems for slender tori. Journal of the Institute of Mathematics and its Applications, 24(3): 303–325, 1979. CODEN JMTAA8. ISSN 0020-2932.

Shampine:1979:EPF

[683] L. F. Shampine. Efficiency of phase function methods for Sturm–Liouville eigenvalues. Journal of the Institute of Mathematics and its Applications, 23(4):413– 420, 1979. CODEN JMTAA8. ISSN 0020-2932.

Shercliff:1970:NEC

[684] J. A. Shercliff. A note on the envelope construction for group velocity in dispersive, anisotropic wave systems. Journal of the Institute of Mathematics and its Applications, 6(??):91–95, 1970. CO-DEN JMTAA8. ISSN 0020-2932.

Shercliff:1970:FPB

[685] J. A. Shercliff. On flow past bodies in aligned magnetic fields. Journal of the Institute of Mathematics and its Applications, 6(??):96–101, 1970. CODEN JMTAA8. ISSN 0020-2932.

Shercliff:1975:NV

[686] J. A. Shercliff. A note on vorticity.

Journal of the Institute of Mathematics and its Applications, 16(2):259–262, 1975. CODEN JMTAA8. ISSN 0020-2932.

Shinbrot:1980:WWP

[687] M. Shinbrot. Water waves over periodic bottoms in three dimensions. Journal of the Institute of Mathematics and its Applications, 25(4):367–385, 1980. CO-DEN JMTAA8. ISSN 0020-2932.

Shukla:1977:OEP

[688] J. B. Shukla and M. Isa. Optimum externally pressurized magnetohydrodynamic bearing. Journal of the Institute of Mathematics and its Applications, 19 (??):49–58, 1977. CODEN JMTAA8. ISSN 0020-2932.

Sidi:1979:SPG

[689] A. Sidi. Some properties of a generalization of the Richardson extrapolation process. Journal of the Institute of Mathematics and its Applications, 24(3): 327–346, 1979. CODEN JMTAA8. ISSN 0020-2932.

Sidi:1980:EMO

[690] A. Sidi. Extrapolation methods for oscillatory infinite integrals. Journal of the Institute of Mathematics and its Applications, 26(1):1–20, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Sleeman:1967:LFS

[691] B. D. Sleeman. The low-frequency scalar Dirichlet scattering by a general ellipsoid. Journal of the Institute of Mathematics and its Applications, 3(??):291– 312, 1967. CODEN JMTAA8. ISSN 0020-2932.

Sleeman:1968:PCF

[692] B. D. Sleeman. On parabolic cylinder functions. Journal of the Institute of Mathematics and its Applications, 4(??): 106–112, 1968. CODEN JMTAA8. ISSN 0020-2932.

Sleeman:1969:DSW

[693] B. D. Sleeman. On diffraction at short wavelengths by a prolate spheroid. Journal of the Institute of Mathematics and its Applications, 5(4):432–442, 1969. CODEN JMTAA8. ISSN 0020-2932.

Sloan:1976:RTE

[694] D. M. Sloan and G. Wilks. Riccati transformations for eigenvalues of systems of linear ordinary differential equations with separated boundary conditions. Journal of the Institute of Mathematics and its Applications, 18(1):117– 127, 1976. CODEN JMTAA8. ISSN 0020-2932.

Smarzewski:1978:NSC

[695] R. Smarzewski and H. Malinowski. Numerical solutions of a class of Abel integral equations. Journal of the Institute of Mathematics and its Applications, 22 (2):159–170, 1978. CODEN JMTAA8. ISSN 0020-2932.

Smith:1974:BLF

[696] F. T. Smith. Boundary layer flow near a discontinuity in wall conditions. Journal of the Institute of Mathematics and its Applications, 13(??):127–145, 1974. CO-DEN JMTAA8. ISSN 0020-2932.

Smith:1978:FTS

[697] F. T. Smith. Flow through symmetrically constricted tubes. Journal of the Institute of Mathematics and its Applications, 21(??):145–156, 1978. CODEN JMTAA8. ISSN 0020-2932.

Smith:1976:NDE

[698] P. Smith. A note on dual extremum principles in magnetohydrodynamic pipe flow. Journal of the Institute of Mathematics and its Applications, 18 (1):129–133, 1976. CODEN JMTAA8. ISSN 0020-2932.

Smith:1978:AOD

[699] Peter Smith. Approximate operators and dual extremum principles for linear problems. Journal of the Institute of Mathematics and its Applications, 22 (4):457–465, 1978. CODEN JMTAA8. ISSN 0020-2932.

Smith:1969:RSW

[700] S. H. Smith. The reflection of surface waves by two parallel vertical barriers. Journal of the Institute of Mathematics and its Applications, 5(??):217–229, 1969. CODEN JMTAA8. ISSN 0020-2932.

Sozou:1971:BLG

[701] C. Sozou. Boundary layer growth on a spinning sphere. Journal of the Institute of Mathematics and its Applications, 7 (??):251–259, 1971. CODEN JMTAA8. ISSN 0020-2932.

Spanier:1979:NFE

[702] J. Spanier. A new family of estimators for random walk problems. Journal of the Institute of Mathematics and its Applications, 23(??):1–31, 1979. CODEN JMTAA8. ISSN 0020-2932.

Spedicato:1978:NDS

[703] E. Spedicato. A note on the determination of the scaling parameters in a class of quasi-Newton methods for unconstrained minimization. Journal of the Institute of Mathematics and its Applications, 21(??):285–291, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Spencer:1970:STF

[704] A. J. M. Spencer. The static theory of finite elasticity. Journal of the Institute of Mathematics and its Applications, 6 (??):164–200, 1970. CODEN JMTAA8. ISSN 0020-2932.

Stassinopoulos:1980:CDC

[705] G. I. Stassinopoulos. Continuous dependence of constrained solutions of linear differential inclusions. Journal of the Institute of Mathematics and its Applications, 25(2):161–175, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Stassinopoulos:1978:CCS

[706] G. I. Stassinopoulos and R. B. Vinter. Conditions for convergence of solutions in the computation of optimal controls. Journal of the Institute of Mathematics and its Applications, 22(1):1–14, 1978. CODEN JMTAA8. ISSN 0020-2932.

Steinberg:1972:NSB

[707] J. Steinberg. Numerical solution of a boundary value problem arising in the deflection of beams and shells. Journal of the Institute of Mathematics and its Applications, 10(??):10–14, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Steinberg:1977:NSS

[708] J. Steinberg. Numerical solution of a singular two-point boundary value problem. Journal of the Institute of Mathematics and its Applications, 20(1):45– 52, 1977. CODEN JMTAA8. ISSN 0020-2932.

Steinberg:1980:NTS

[709] J. Steinberg. Numerical treatment of singularities in elliptic problems. Journal of the Institute of Mathematics and its Applications, 26(4):349–368, 1980. CODEN JMTAA8. ISSN 0020-2932.

Stenger:1973:IFB

[710] Frank Stenger. Integration formulae based on the trapezoidal formula. Journal of the Institute of Mathematics and its Applications, 12(1):103–114, 1973. CODEN JMTAA8. ISSN 0020-2932. See remarks [711, 668].

Stenger:1977:RIF

[711] Frank Stenger. Remarks on “Integration formulae based on the trapezoidal formula” (J. Inst. Math. Appl. 12 (1973), 103–114). Journal of the Institute of Mathematics and its Applications, 19(2): 145–147, 1977. CODEN JMTAA8. ISSN 0020-2932. See [710, 668].

Stewart:1979:ERE

[712] G. W. Stewart. The effects of rounding error on an algorithm for downdating a Cholesky factorization. Journal of the Institute of Mathematics and its Applications, 23(2):203–213, March 1979. CODEN JMTAA8. ISSN 0020-2932.

Stewartson:1965:MBL

[713] K. Stewartson. On magnetic boundary layers. Journal of the Institute of Mathematics and its Applications, 1(??):29–41, 1965. CODEN JMTAA8. ISSN 0020-2932.

Stirzaker:1975:PMS

[714] D. R. Stirzaker. A perturbation method for the stochastic recurrent epidemic. Journal of the Institute of Mathematics and its Applications, 15(??):135–160, 1975. CODEN JMTAA8. ISSN 0020-2932.

Stirzaker:1976:ME

[715] D. R. Stirzaker. On models for an epidemic. Journal of the Institute of Mathematics and its Applications, 18(2):135– 148, 1976. CODEN JMTAA8. ISSN 0020-2932.

Stow:1972:SIT

[716] P. Stow. The solution of isentropic transonic flow. Journal of the Institute of Mathematics and its Applications, 9(??): 35–46, 1972. CODEN JMTAA8. ISSN 0020-2932.

Strieder:1971:VBP

[717] W. Strieder and R. Aris. Variational bounds for problems in diffusion and reaction. Journal of the Institute of Mathematics and its Applications, 8(??):328– 334, 1971. CODEN JMTAA8. ISSN 0020-2932.

Summers:1978:FMS

[718] D. Summers. Fluid models of the solar wind. Journal of the Institute of Mathematics and its Applications, 22(??):71– 87, 1978. CODEN JMTAA8. ISSN 0020-2932.

Sweet:1971:OPS

[719] R. A. Sweet. Oscillation properties of a semidiscrete approximation to the beam equation with a second order term. Journal of the Institute of Mathematics and its Applications, 7(??):119–125, 1971. CODEN JMTAA8. ISSN 0020-2932.

Sweet:1969:PSD

[720] Roland A. Sweet. Properties of a semidiscrete approximation to the beam equation. Journal of the Institute of Mathematics and its Applications, 5(??): 329–339, 1969. CODEN JMTAA8. ISSN 0020-2932.

Talbot:1979:ANI

[721] A. Talbot. The accurate numerical inversion of Laplace transforms. Journal of the Institute of Mathematics and its Applications, 23(1):97–120, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Taylor:1971:NAD

[722] R. S. Taylor. A new approach to the delta wing problem. Journal of the Institute of Mathematics and its Applications, 7(??):337–347, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Taylor:1971:LEU

[723] R. S. Taylor and F. M. Arscott. Lamé’s equation and its use in elliptic cone problems. Journal of the Institute of Mathematics and its Applications, 7(??):331– 336, 1971. CODEN JMTAA8. ISSN 0020-2932.

Telega:1979:VFN

[724] J. J. Telega. On variational formulations for nonlinear, nonpotential operators. Journal of the Institute of Mathematics and its Applications, 24(2):175– 195, 1979. CODEN JMTAA8. ISSN 0020-2932.

Temme:1978:UAE

[725] N. M. Temme. Uniform asymptotic expansions of confluent hypergeometric functions. Journal of the Institute of Mathematics and its Applications, 22(2): 215–223, 1978. CODEN JMTAA8. ISSN 0020-2932.

Temperley:1976:SRC

[726] D. J. Temperley and L. Todd. Some remarks on a class of plane, linear boundary value problems. Journal of the Institute of Mathematics and its Applications, 18(3):309–324, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Teymur:1978:WPD

[727] M. Teymur and E. Suhubi. Wave propagation in dissipative or dispersive nonlinear media. Journal of the Institute of Mathematics and its Applications, 21 (??):25–40, 1978. CODEN JMTAA8. ISSN 0020-2932.

Thatcher:1975:SSL

[728] R. W. Thatcher. Singularities in the solution of Laplace’s equation in two dimensions. Journal of the Institute of Mathematics and its Applications, 16(3): 303–319, 1975. CODEN JMTAA8. ISSN 0020-2932.

Thatcher:1978:CEN

[729] R. W. Thatcher. Controlling the error in the numerical solution of elliptic equations. Journal of the Institute of Mathematics and its Applications, 22(4):443– 455, 1978. CODEN JMTAA8. ISSN 0020-2932.

Theocaris:1977:NIM

[730] P. S. Theocaris and A. C. Chrysakis.

Numerical inversion of the Mellin transform. Journal of the Institute of Mathematics and its Applications, 20(1):73– 83, 1977. CODEN JMTAA8. ISSN 0020-2932.

Thomas:1974:SBV

[731] J. N. Thomas and A. H. England. The stress boundary value problem for an ideal fibre-reinforced rectangular plate with a hole. Journal of the Institute of Mathematics and its Applications, 14 (??):347–374, 1974. CODEN JMTAA8. ISSN 0020-2932.

Thomee:1973:PDA

[732] Vidar Thomée. Polygonal domain approximation in Dirichlet’s problem. Journal of the Institute of Mathematics and its Applications, 11(??):33–44, 1973. CODEN JMTAA8. ISSN 0020-2932.

Titterington:1973:MEA

[733] D. M. Titterington. A method of extremum adaptation. Journal of the Institute of Mathematics and its Applications, 11(??):297–315, 1973. CODEN JMTAA8. ISSN 0020-2932.

Tolstoy:1977:RLD

[734] I. Tolstoy. Remarks on the linearization of differential equations. Journal of the Institute of Mathematics and its Applications, 20(1):53–60, 1977. CODEN JM-TAA8. ISSN 0020-2932.

Tomlin:1972:PSS

[735] J. A. Tomlin. Pivoting for size and sparsity in linear programming inversion routes. Journal of the Institute of Mathematics and its Applications, 10 (??):289–295, 1972. CODEN JMTAA8. ISSN 0020-2932.

Troth:1976:MRA

[736] M. R. Troth. Materials reinforced by almost inextensible fibres. Journal of the Institute of Mathematics and its Applications, 18(??):265–278, 1976. CODEN JMTAA8. ISSN 0020-2932.

Turland:1980:SPM

[737] B. D. Turland and R. S. Peckover. The stability of planar melting fronts in twophase thermal Stefan problems. Journal of the Institute of Mathematics and its Applications, 25(1):1–15, 1980. CODEN JMTAA8. ISSN 0020-2932.

Turner:1978:CPC

[738] J. R. Turner. A convergence proof for constrained finite element solutions. Journal of the Institute of Mathematics and its Applications, 22(??):121–127, 1978. CODEN JMTAA8. ISSN 0020-2932.

Turner:1979:FUP

[739] J. R. Turner. The frictional unloading problem on a linear elastic half-space. Journal of the Institute of Mathematics and its Applications, 24(4):439–469, 1979. CODEN JMTAA8. ISSN 0020-2932.

Turner:1980:SPC

[740] J. R. Turner and D. A. C. Nicol. The Signorini problem for a Cosserat plate. Journal of the Institute of Mathematics and its Applications, 25(2):133–145, 1980. CODEN JMTAA8. ISSN 0020-2932.

Twizell:1976:SAC

[741] E. H. Twizell. Some approximations of the characteristics of second order hyperbolic partial differential equations. Journal of the Institute of Mathematics and its Applications, 17(2):209–218, 1976. CODEN JMTAA8. ISSN 0020-2932.

Underhill:1973:CPP

[742] C. Underhill and A. Wragg. Convergence properties of Padé approximants to exp(z) and their derivatives. Journal of the Institute of Mathematics and its Applications, 11(??):361–367, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Urwin:1967:PSW

[743] Kathleen M. Urwin. Perturbation solutions of the Whittaker-Hill equation. Journal of the Institute of Mathematics and its Applications, 3(??):169–180, 1967. CODEN JMTAA8. ISSN 0020-2932.

Usmani:1977:FDA

[744] R. A. Usmani. An O(h^{6}) finite difference analogue for the solution of some differential equations occurring in plate deflection theory. Journal of the Institute of Mathematics and its Applications, 20 (3):331–333, 1977. CODEN JMTAA8. ISSN 0020-2932.

Usmani:1971:FDA

[745] Riaz A. Usmani. An O(h^{6}) finite difference analogue for the numerical solution of a two point boundary value problem. Journal of the Institute of Mathematics and its Applications, 8(??):335–343, 1971. CODEN JMTAA8. ISSN 0020-2932.

Vajda:1974:TOC

[746] S. Vajda. Tests of optimality in constrained optimization. Journal of the Institute of Mathematics and its Applications, 13(??):187–200, 1974. CODEN JMTAA8. ISSN 0020-2932.

Valvi:1977:EPI

[747] F. Valvi. Explicit presentation of the inverses of some types of matrices. Journal of the Institute of Mathematics and its Applications, 19(1):107–117, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

vandenBerg:1976:RPC

[748] P. M. van den Berg. Radiation from a point charge moving through a circular aperture in a screen. Journal of the Institute of Mathematics and its Applications, 17(1):119–133, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Vardulakis:1978:GRT

[749] A. I. G. Vardulakis and P. N. R. Stoyle. Generalized resultant theorem. Journal of the Institute of Mathematics and its Applications, 22(3):331–335, 1978. CO-DEN JMTAA8. ISSN 0020-2932.

Varley:1965:NLT

[750] E. Varley and E. Cumberbatch. Nonlinear theory of wave-front propagation. Journal of the Institute of Mathematics and its Applications, 1(??):101–112, 1965. CODEN JMTAA8. ISSN 0020-2932.

Varley:1966:NLH

[751] E. Varley and E. Cumberbatch. Nonlinear, high frequency sound waves.

Journal of the Institute of Mathematics and its Applications, 2(??):133–143, 1966. CODEN JMTAA8. ISSN 0020-2932.

Verhulst:1976:MAE

[752] F. Verhulst. Matched asymptotic expansions in the two-body problem with quick loss of mass. Journal of the Institute of Mathematics and its Applications, 18(1):87–98, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Vinter:1977:SSD

[753] R. B. Vinter. Stabilizability and semigroups with discrete generators. Journal of the Institute of Mathematics and its Applications, 20(3):371–378, 1977. CO-DEN JMTAA8. ISSN 0020-2932.

Vinter:1978:ESL

[754] R. B. Vinter. On the evolution of the state of linear differential delay equations in M^{2}: properties of the generator. Journal of the Institute of Mathematics and its Applications, 21(??):13–23, 1978. CODEN JMTAA8. ISSN 0020-2932.

Visnak:1978:CNM

[755] K. Visnak and M. Kubicek. A class of numerical methods of high order for stiff problems in ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 21(??):251– 264, 1978. CODEN JMTAA8. ISSN 0020-2932.

Wachspress:1971:RBF

[756] Eugene L. Wachspress. A rational basis for function approximation. Journal of the Institute of Mathematics and its Applications, 8(??):57–68, 1971. CODEN JMTAA8. ISSN 0020-2932.

Wachspress:1973:RBF

[757] Eugene L. Wachspress. A rational basis for function approximation. II: Curved sides. Journal of the Institute of Mathematics and its Applications, 11(??):83– 104, 1973. CODEN JMTAA8. ISSN 0020-2932.

Waechter:1970:ISW

[758] R. T. Waechter. The initial structure of wing-body interaction in supersonic flow. Journal of the Institute of Mathematics and its Applications, 6(??):310– 318, 1970. CODEN JMTAA8. ISSN 0020-2932.

Wait:1977:SIF

[759] R. Wait. Singular isoparametric finite elements. Journal of the Institute of Mathematics and its Applications, 20(1):133– 141, 1977. CODEN JMTAA8. ISSN 0020-2932.

Wait:1971:STD

[760] R. Wait and A. R. Mitchell. The solution of time dependent problems by Galerkin methods. Journal of the Institute of Mathematics and its Applications, 7(??):241–250, 1971. CODEN JM-TAA8. ISSN 0020-2932.

Waldron:1968:IBF

[761] R. A. Waldron. The integration of Bessel functions. Journal of the Institute of Mathematics and its Applications, 4(??): 315–319, 1968. CODEN JMTAA8. ISSN 0020-2932.

Walpole:1974:NEP

[762] L. J. Walpole. New extremum principles for linear problems. Journal of the Institute of Mathematics and its Applications, 14(??):113–118, 1974. CODEN JMTAA8. ISSN 0020-2932.

Walsh:1966:FNC

[763] G. R. Walsh. The fourth necessary condition in an optimal rocket trajectory. Journal of the Institute of Mathematics and its Applications, 2(??):237–247, 1966. CODEN JMTAA8. ISSN 0020-2932.

Walter:1976:SEM

[764] G. Walter and D. Schultz. Some eigenfunction methods for computing a numerical Fourier transform. Journal of the Institute of Mathematics and its Applications, 18(3):279–293, 1976. CO-DEN JMTAA8. ISSN 0020-2932.

Wang:1976:NCE

[765] J. Y. Wang. On the numerical computation of eigenvalues and eigenfunctions of compact integral operators using spline functions. Journal of the Institute of Mathematics and its Applications, 18(2): 177–188, 1976. CODEN JMTAA8. ISSN 0020-2932.

Wang:1973:RSM

[766] Yen-Chu Wang. Review of some mathematical models of non-linear domain dynamics in bulk- effect semiconductors. Journal of the Institute of Mathematics and its Applications, 11(??):251–280, 1973. CODEN JMTAA8. ISSN 0020-2932.

Warren:1970:MFB

[767] F. W. G. Warren. A method for finding bounds for complex eigenvalues of second order systems. Journal of the Institute of Mathematics and its Applications, 6(??):21–26, 1970. CODEN JM-TAA8. ISSN 0020-2932.

Watkins:1977:SFF

[768] D. S. Watkins and P. Lancaster. Some families of finite elements. Journal of the Institute of Mathematics and its Applications, 19(4):385–397, 1977. CODEN JMTAA8. ISSN 0020-2932.

Watson:1973:OSA

[769] G. A. Watson. One-sided approximation and operator equations. Journal of the Institute of Mathematics and its Applications, 12(??):197–208, 1973. CODEN JMTAA8. ISSN 0020-2932.

Watson:1976:MCB

[770] G. A. Watson. A method for calculating best non-linear Chebyshev approximations. Journal of the Institute of Mathematics and its Applications, 18(3):351– 360, 1976. CODEN JMTAA8. ISSN 0020-2932.

Watson:1979:MSO

[771] G. A. Watson. The minimax solution of an overdetermined system of non-linear equations. Journal of the Institute of Mathematics and its Applications, 23(2): 167–180, March 1979. CODEN JM-TAA8. ISSN 0020-2932.

Watson:1965:ESC

[772] J. Watson. On the existence of solutions for a class of rotating disc flows and the convergence of a successive approximation scheme. Journal of the Institute of Mathematics and its Applications, 1(??): 348–371, 1965. CODEN JMTAA8. ISSN 0020-2932.

Watson:1971:TMD

[773] S. R. Watson. Towards the minimum drag on a body of given volume in slow viscous flow. Journal of the Institute of Mathematics and its Applications, 7(??): 367–376, 1971. CODEN JMTAA8. ISSN 0020-2932.

Weare:1979:DDF

[774] T. J. Weare. Damping in “un-damped”

finite difference representations. Journal of the Institute of Mathematics and its Applications, 23(??):86–96, 1979. CO-DEN JMTAA8. ISSN 0020-2932.

Westbrook:1974:VPA

[775] D. R. Westbrook. A variational principle with applications in finite elements. Journal of the Institute of Mathematics and its Applications, 14(??):79–82, 1974. CODEN JMTAA8. ISSN 0020-2932.

Westreich:1980:NSE

[776] D. Westreich. Numerical solution of the eigenvalue problem for discontinuous linear ordinary differential equations. Journal of the Institute of Mathematics and its Applications, 25(2):147–160, 1980. CODEN JMTAA8. ISSN 0020-2932.

Westreich:1980:NSV

[777] David Westreich and Baruch Cahlon.

Numerical solutions of Volterra integral equations with continuous or discontinuous terms. Journal of the Institute of Mathematics and its Applications, 26(2): 175–186, 1980. CODEN JMTAA8. ISSN 0020-2932.

White:1977:USD

[778] P. S. White. On the ultimate stressdeviator in elastic-perfectly plastic solids. Journal of the Institute of Mathematics and its Applications, 19(1):97– 105, 1977. CODEN JMTAA8. ISSN 0020-2932.

Whiten:1979:MLC

[779] W. J. Whiten. Multiple linear curve fitting. Journal of the Institute of Mathematics and its Applications, 23(1):61– 71, 1979. CODEN JMTAA8. ISSN 0020-2932.

Whittle:1970:RCS

[780] P. Whittle and P. A. Gait. Reduction of a class of stochastic control problems. Journal of the Institute of Mathematics and its Applications, 6(??):131–140, 1970. CODEN JMTAA8. ISSN 0020-2932.

Wilks:1975:RPM

[781] G. Wilks. On Rayleigh’s problem in magnetohydrodynamics. Journal of the Institute of Mathematics and its Applications, 15(??):73–84, 1975. CODEN JMTAA8. ISSN 0020-2932.

Wilks:1976:IIR

[782] G. Wilks and D. M. Sloan. Invariant imbedding, Riccati transformations and eigenvalue problems. Journal of the Institute of Mathematics and its Applications, 18(1):99–116, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Williams:1966:RDD

[783] J. E. F. Williams. Radiation from a dipole distribution in inhomogeneous convective motion. Journal of the Institute of Mathematics and its Applications, 2(??):223–227, 1966. CODEN JM-TAA8. ISSN 0020-2932.

Williams:1980:ISI

[784] M. H. Williams. The inversion of singular integral equations by expansion in Jacobi polynomials. Journal of the Institute of Mathematics and its Applications, 25(4):413–426, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Williams:1965:IEF

[785] W. E. Williams. Integral equation formulation of Oseen flow problems. Journal of the Institute of Mathematics and its Applications, 1(??):339–347, 1965. CODEN JMTAA8. ISSN 0020-2932.

Williams:1966:EDA

[786] W. E. Williams. Electromagnetic diffraction in an anisotropic medium. Journal of the Institute of Mathematics and its Applications, 2(??):186–196, 1966. CODEN JMTAA8. ISSN 0020-2932.

Williams:1978:NSI

[787] W. E. Williams. Note on a singular integral equation. Journal of the Institute of Mathematics and its Applications, 22 (2):211–214, 1978. CODEN JMTAA8. ISSN 0020-2932.

Willis:1971:ISI

[788] J. R. Willis. Interfacial stresses induced by arbitrary loading of dissimilar elastic half-spaces joined over a circular region. Journal of the Institute of Mathematics and its Applications, 7(??):179– 197, 1971. CODEN JMTAA8. ISSN 0020-2932.

Wilson:1976:CPS

[789] D. A. Wilson and J. E. Rubio. A controllability problem for a system governed by the heat equation. Journal of the Institute of Mathematics and its Applications, 17(1):43–51, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Wilson:1973:NMA

[790] G. A. Wilson and A. Wragg. Numerical methods for approximating continuous probability density functions, over [0, ∞), using moments. Journal of the Institute of Mathematics and its Applications, 12(??):165–173, 1973. CODEN JMTAA8. ISSN 0020-2932.

Wilson:1972:SRT

[791] J. C. Wilson. Stability of Richtmyer type difference schemes in any finite number of space variables and their comparison with multistep Strang schemes. Journal of the Institute of Mathematics and its Applications, 10(??):238–257, 1972. CO-DEN JMTAA8. ISSN 0020-2932.

Wind:1977:RCS

[792] H. Wind. The rounded cubic spline. Journal of the Institute of Mathematics and its Applications, 19(4):441–445, 1977. CODEN JMTAA8. ISSN 0020-2932.

Winfield:1973:FMI

[793] D. Winfield. Function minimization by interpolation in a data table. Journal of the Institute of Mathematics and its Applications, 12(??):339–347, 1973. CO-DEN JMTAA8. ISSN 0020-2932.

Wolfe:1975:QNM

[794] M. A. Wolfe. A quasi-Newton method with memory for unconstrained function minimization. Journal of the Institute of Mathematics and its Applications, 15 (??):85–94, 1975. CODEN JMTAA8. ISSN 0020-2932.

Wolfe:1976:SML

[795] M. A. Wolfe. Some methods for least squares estimation. Journal of the Institute of Mathematics and its Applications, 18(2):219–236, 1976. CODEN JM-TAA8. ISSN 0020-2932.

Wong:1967:GAS

[796] James S. W. Wong. On the global asymptotic stability of (p(t)x′)′ + q(t)x^{2n−1} = 0. Journal of the Institute of Mathematics and its Applications, 3 (??):403–405, 1967. CODEN JMTAA8. ISSN 0020-2932.

Wood:1974:EED

[797] A. D. Wood. Eigenfunction expansions by the delta function and contour integral method. Journal of the Institute of Mathematics and its Applications, 14 (??):57–67, 1974. CODEN JMTAA8. ISSN 0020-2932.

Wood:1973:GES

[798] D. H. Wood. Generating examples of shock propagation by three principles. Journal of the Institute of Mathematics and its Applications, 11(??):9–14, 1973. CODEN JMTAA8. ISSN 0020-2932.

Wragg:1973:CEM

[799] A. Wragg and C. Davies. Computation of the exponential of a matrix I: Theoretical considerations. Journal of the Institute of Mathematics and its Applications, 11(??):369–375, 1973. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Wragg:1975:CEM

[800] A. Wragg and C. Davies. Computation of the exponential of a matrix II: Practical considerations. Journal of the Institute of Mathematics and its Applications, 15(??):273–278, 1975. CODEN JMTAA8. ISSN 0020-2932. Cited in [?].

Yolles:1975:LSS

[801] M. I. Yolles. Least squares successive relaxation. Journal of the Institute of Mathematics and its Applications, 16(3):

329–343, 1975. CODEN JMTAA8. ISSN 0020-2932.

Young:1979:BAA

[802] A. Young and E. A. Kiountouzis. Best approximation in an asymmetrically weighted L_{1} measure. Journal of the Institute of Mathematics and its Applications, 24(4):379–394, 1979. CODEN JM-TAA8. ISSN 0020-2932.

Young:1971:SDT

[803] Jonathan D. Young. Smoothing data with tolerances by use of linear programming. Journal of the Institute of Mathematics and its Applications, 8(??):69–79, 1971. CODEN JMTAA8. ISSN 0020-2932.

Zakian:1975:AAN

[804] V. Zakian. Application of J_{MN} approximants to numerical initial-value problems in linear differential-algebraic systems. Journal of the Institute of Mathematics and its Applications, 15(??):267– 272, 1975. CODEN JMTAA8. ISSN 0020-2932.

Zakrajsek:1976:GEP

[805] E. Zakrajšek. On the generalized eigenvalue problem. Journal of the Institute of Mathematics and its Applications, 18 (2):155–157, 1976. CODEN JMTAA8. ISSN 0020-2932.

Zauderer:1971:MHM

[806] Erich Zauderer. On a modification of Hadamard’s method for obtaining fundamental solutions for hyperbolic and parabolic equations. Journal of the Institute of Mathematics and its Applications, 8(??):8–15, 1971. CODEN JM-TAA8. ISSN 0020-2932.